### v3

# Learn With Imagination: Safe Set Guided State-wise Constrained Policy Optimization

Abstract
Deep reinforcement learning (RL) excels in various control tasks, yet the absence of safety guarantees hampers its real-world applicability. In particular, explorations during learning usually results in safety violations, while the RL agent learns from those mistakes. On the other hand, safe control techniques ensure persistent safety satisfaction but demand strong priors on system dynamics, which is usually hard to obtain in practice. To address these problems, we present Safe Set Guided State-wise Constrained Policy Optimization (S-3PO), a pioneering algorithm generating state-wise safe optimal policies with zero training violations, i.e., learning without mistakes. S-3PO first employs a safety-oriented monitor with black-box dynamics to ensure safe exploration. It then enforces an "imaginary" cost for the RL agent to converge to optimal behaviors within safety constraints. S-3PO outperforms existing methods in high-dimensional robotics tasks, managing state-wise constraints with zero training violation. This innovation marks a significant stride towards real-world safe RL deployment.
1 Introduction
Reinforcement Learning (RL) has showcased remarkable advancements in domains like control and games. However, its focus on reward maximization sometimes neglects safety, potentially leading to catastrophic outcomes (Gu et al., 2022). To rectify this, the concept of safe RL emerged, aiming to ensure safety throughout or after training. Initial endeavors centered on Constrained Markov Decision Processes, often emphasizing cumulative or chance constraints (Ray et al., 2019; Achiam et al., 2017; Liu et al., 2021). While effective, these approaches lack instantaneous safety guarantees, which is crucial for managing emergencies like collision avoidance in autonomous vehicles (Zhao et al., 2023c; He et al., 2023a). Recent strides Zhao et al. (2023b) leveraged the Maximum Markov Decision Process to ensure instantaneous safety by bounding violations while integrating trust region techniques for policy enhancement, resulting in simultaneous improvement of worst-case performance and adherence to cost constraints. However, these approaches still cannot ensure safety during learning due to potentially unsafe explorative behaviors.
Meanwhile, safe control methods to continuously meet stringent safety requirements in predictable environments are widely examined, with energy function-based methods being the most prevalent Khatib (1986); Ames et al. (2014); Liu & Tomizuka (2014); Gracia et al. (2013); Wei & Liu (2019). These methods establish energy functions assigning lower energy to safe states and project nominal control into energy dissipating control, hence persistently maintaining system safety. However, these approahces heavily rely on accessible white-box analytical models of system dynamics. The Implicit Safe Set Algorithm (ISSA) Zhao et al. (2021) addresses this problem using black-box optimization over black-box dynamics models. Nevertheless, the reliance on continuous monitoring to ensure safety imposes significant computational overhead, and the RL policy itself does not inherently guarantee safe behavior. This combination presents a major obstacle to the successful deployment of the algorithm in real-time, safety-critical applications.
Figure 1: Overview of the principles of the S-3PO algorithm.
In summary, a pressing need arises for a method leveraging both the zero-training-time-violation ability of safe control and optimal convergence of safe RL. To bridge this gap, we introduce a novel Safe Set Guided State-wise Constrained Policy Optimization (S-3PO) algorithm. S-3PO safeguards the exploration of immature policies with a black-box safe control method and derives a novel formulation where RL learns an optimal safe policy by constraining the state-wise “imaginary" safety violations. An overview of the prescribed S-3PO principles is presented in Figure 1. Empirical validation underscores S-3PO’s efficacy in training neural network policies encompassing thousands of parameters for high-dimensional simulated robot locomotion tasks. Our contribution marks a substantial advancement in the realm of practical safe RL algorithms, poised to find applications in a multitude of real-world challenges.
2 Problem Formulation
2.1 Preliminaries
Dynamics
We consider a robot system described by its state $s_{t}\in\mathcal{S}\subset\mathbb{R}^{n_{s}}$ at time step $t$, with $n_{s}$ denoting the dimension of the state space $\mathcal{S}$, and its action input $a_{t}\in\mathcal{A}\subset\mathbb{R}^{n_{a}}$ at time step $t$, where $n_{a}$ represents the dimension of the control space $\mathcal{A}$. The system dynamics are defined as follows:
$\begin{split}&s_{t+1}=f(s_{t},a_{t}),\\ \end{split}$ (1)
where $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is a deterministic function that maps the current robot state and control to the robot’s state in the next time step.
To maintain simplicity, our approach focuses on deterministic dynamics, although it is worth noting that the proposed method can be readily extended to accommodate stochastic dynamics Zhao et al. (2021); Noren et al. (2021). Additionally, we assume the access to the dynamics model $f$ is only in the training phase, and is restricted to an implicit black-box form, as exemplified by an implicit digital twin simulator or a deep neural network model Zhao et al. (2021). We also assume there is no model mismatch, while model mismatch can be addressed by robust safe control Wei et al. (2022) and is left for future work. Post training, the knowledge of the dynamics model is concealed, aligning with practical scenarios where digital twins of real-world environments are too costly to access during model deployment.
Markov Decision Process
In this research, our primary focus lies in ensuring safety for episodic tasks, which falls within the purview of finite-horizon Markov Decision Processes (MDP). An MDP is defined by a tuple $(\mathcal{S},\mathcal{A},\gamma,R,P,\mu)$. The reward function is denoted by $R:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$, the discount factor by $10\leq\gamma<$ 1, the initial state distribution by $\mu:\mathcal{S}\mapsto\mathbb{R}$, and the transition probability function by $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$.
The transition probability $P(s^{\prime}|s,a)$ represents the likelihood of transitioning to state $s^{\prime}$ when the previous state was $s$, and the agent executed action $a$ at state $s$. This paper assumes deterministic dynamics, implying that $P(s_{t+1}|s_{t},a_{t})=1$ when $s_{t+1}=f(s_{t},a_{t})$. We denote the set of all stationary policies as $\Pi$, and we further denote $\pi_{\theta}$ as a policy parameterized by the parameter $\theta$.
In the context of an MDP, our ultimate objective is to learn a policy $\pi$ that maximizes a performance measure $\mathcal{J}(\pi)$, computed via the discounted sum of rewards, as follows:
$\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{H}\gamma^{t}R(s_{t},a_{t},s_{t+1})\right],$ (2)
where $H\in\mathbb{N}$ denotes the horizon, $\tau=[s_{0},a_{0},s_{1},\cdots]$, and $\tau\sim\pi$ indicates that the distribution over trajectories depends on $\pi$, i.e., $s_{0}\sim\mu$, $a_{t}\sim\pi(\cdot|s_{t})$, and $s_{t+1}\sim P(\cdot|s_{t},a_{t})$.
Safety Specification
The safety specification requires that the system state remains within a closed subset in the state space, denoted as the “safe set” $\mathcal{S}_{S}$. This safe set is defined by the zero-sublevel set of a continuous and piecewise smooth function $\phi_{0}:\mathbb{R}^{n_{s}}\rightarrow\mathbb{R}$, where $\mathcal{S}_{S}=\{s\mid\phi_{0}(s)\leq 0\}$. Users directly specify both $\mathcal{S}_{S}$ and $\phi_{0}$, which is easy to specify. For instance, for collision avoidance, $\phi_{0}$ can be specified as the negative closest distance between the robot and environmental obstacles.
2.2 Problem
We are interested in the safety imperative of averting collisions for mobile robots navigating 2D planes. We aim to persistently satisfy safety specifications at every time step while solving MDP, following the intuition of State-wise Constrained Markov Decision Process (SCMDP) Zhao et al. (2023c). Formally, the set of feasible stationary policies for SCMDP is defined as
$\bar{\Pi}_{C}=\{\pi\in\Pi\big{|}~{}\forall s_{t}\sim\tau,s_{t}\in\mathcal{S}_{S}\},$ (3)
where $\tau\sim\pi$. Then, the objective for SCMDP is to find a feasible stationary policy from $\bar{\Pi}_{C}$ that maximizes the performance measure. Formally,
$\max_{\theta}\mathcal{J}(\pi_{\theta}),\text{ s.t. }\pi_{\theta}\in\bar{\Pi}_{C}.$ (4)
State-wise Safe Policy with Zero Violation Training
The primary focus of this paper centers on solving (4), i.e., ensuring no safety violation during the training process, while achieving convergence of the policy to the optimal solution of (4).
3 Prior Works
3.1 Safe Reinforcement Learning
Existing safe RL approaches either consider safety after convergence or safety during training Zhao et al. (2023c). End-to-end approaches are usually used to ensure safety after convergence Liang et al. (2018); Tessler et al. (2018); Bohez et al. (2019); Ma et al. (2021); He et al. (2023b). However, these approaches cannot avoid unsafe explorations.
Safety in training is achieved by hierarchical approaches which uses a safeguard to filter out unsafe explorative actions. The safeguard relies on the knowledge on the system dynamics, which can be either learned dynamics Dalal et al. (2018); Thananjeyan et al. (2021); Zhao et al. (2022), white-box dynamics Fisac et al. (2018); Shao et al. (2021), or black-box dynamics Zhao et al. (2021). Nevertheless, these methods cannot guarantee convergence of the RL policy.
The proposed approach is the first to address safety both during training and after convergence, which can potentially serve as a general framework to bridge these two types of approaches. In the following, we choose the most advanced method in each category to form the proposed approach, so that it relies on the least assumptions.
3.2 Implicit Safe Set Algorithm
Energy function-based methods  Wei & Liu (2019) achieve safe control by designing an energy function offline such that 1) the low energy states are safe and 2) there always exists a feasible control input to dissipate the energy. A typical design Liu & Tomizuka (2014) was proposed as $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$. It is shown in Liu & Tomizuka (2014) that if the control input is unbounded ($\mathcal{A}=\mathbb{R}^{n_{a}}$), then there always exist a control input that satisfies the constraint $\dot{\phi}\leq 0\text{ when }\phi=0$; and if the control input always satisfies that constraint, then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant. In practice, the actual control signal is computed through a quadratic projection of the nominal control $a^{r}$ to the control constraint
$a=$ $\mathop{\mathbf{argmin}}_{a\in\mathcal{A}}\|a-a^{r}\|^{2}\text{ s.t. }\dot{\phi}\leq-\eta(\phi),$ (5)
where $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$ according to SSA Liu & Tomizuka (2014). Here, we define the set of safe control as $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$. To match the notion of MDP, we should consider discrete-time safe control set $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$.
Based on these, the implicit safe set algorithm (ISSA) Zhao et al. (2021) was proposed to construct the black-box dynamics based safeguard ensuring persistent satisfication of safety specification. The black-box dynamics can be a digital twin or neural network. Under some mild assumptions, it first synthesizes a safety index to make sure $\mathcal{A}_{S}(s)$ is nonempty for all $s$ (details of the mild assumptions and safety index synthesis are summarized in Appendix A). Then it manages to project the reference action generated from RL policy $\pi_{\theta}$ into $\mathcal{A}_{S}(s)$ during policy training. In detail, the nominal control $a_{t}^{r}$ needs to be projected to $\mathcal{A}_{S}^{D}(s_{t})$ by solving the following optimization:
$\begin{split}&\min_{a_{t}\in\mathcal{A}}\|a_{t}-a^{r}_{t}\|^{2}\\ &\text{s.t. }\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}.\end{split}$ (6)
To solve (6), they first introduce a sample-efficient Adaptive Momentum Boundary Approximation (AdamBA) algorithm. Then, ISSA directly uses it to find the safe control with minimum deviation from the reference control along the boundaries of $\mathcal{A}_{S}^{D}(s)$. If the process fails to return a solution, grid sampling will be deployed to find a safe control; and AdamBA is deployed again to improve the solution optimality with respect to (6). We summarize details of AdamBA and ISSA in Algorithm 2 and Algorithm 3, respectively.
3.3 State-wise Constrained Policy Optimization
Safe RL algorithms under the framework of Constrained Markov Decision Process (CMDP) do not consider state-wise constraints. To address this gap, State-wise Constrained Policy Optimization (SCPO) was proposed Zhao et al. (2023b) to provide guarantees for state-wise constraint satisfaction in expectation, which is under the framework of State-wise CMDP (SCMDP). To achieve this, SCPO directly constrain the expected maximum state-wise cost along the trajectory. And they introduced Maximum MDP (MMDP). In this setup, a running maximum cost value is associated with each state, and a non-discounted finite MDP is utilized to track and accumulate non-negative increments in cost. The format of MMDP will be introduced in Section 4.
4 Safety Index Guided State-wise Constrained Policy Optimization
The core idea of S-3PO is to enforce zero safety violation during training by projecting unsafe actions to the safe set, and then constrain the "imaginary" safety violation (i.e., what if the projection is not done) to ensure convergence of the policy to an optimal safe policy.
Zero Violation Exploration
To ensure zero violation exploration, we safeguard nominal control via solving (6) at every time step during policy training. In this paper, we adopt ISSA to solve (6).
For the 2D collision avoidance problem considered in this paper, we choose the same safety index synthesis rule as [Section 4.1, (Zhao et al., 2021)], which is summarized in Appendix A. We show in Section 6 that with the safety index synthesis rule, ISSA is guaranteed to find a feasible solution of (6), making the system forward invariance in the set $\mathcal{\bar{S}}$.
Learning Safety Measures Safely
While eliminating safety violations during training is good, this also brings inevitable challenges for the policy learning as it directly eliminates the unsafe experience, making the policy unable to distinguish between safe and unsafe actions. To overcome this challenge, the main intuition we have is that instead of directly experiencing unsafe states, i.e. $s\notin\mathcal{S}_{S}$, policy can learn to act safely from “imagination", i.e., how unsafe it will be if the safeguard had not been triggered? The critical observation we rely upon is that:
Observation 1.
Define $\Delta\phi_{t}=\Delta\phi(s_{t},a_{t},s_{t+1})\doteq\phi(f(s_{t},a^{r}_{t}))-\phi(f(s_{t},a_{t}))$, i.e. the degree of required correction to safeguard $a_{t}^{r}$. Therefore, $\Delta\phi_{t}$ can be treated as an imagination on how unsafe the reference action would be, where $\Delta\phi_{t}\leq 0$ means $a_{t}^{r}\in\mathcal{A}_{S}^{D}(s_{t})$.
Following 1, Equation 4 can be translated to:
$\max_{\theta}\mathcal{J}(\pi_{\theta}),\text{ s.t. }\pi_{\theta}\in\{\pi\in\Pi\big{|}~{}\forall\Delta\phi_{t}\sim\tau,\Delta\phi_{t}\leq 0\}.$ (7)
Remark 1.
Policies satisfying (7) ensure there is no imaginary safety violation for any possible $a_{t}^{r}$, making $\pi_{\theta}$ a safe policy as required by (4).
4.1 Transfrom State-wise Constraint into Maximum Constraint
For (7), each state-action transition pair corresponds to a constraint, which is intractable to solve. Inspired by Zhao et al. (2023c), we constrain the expected maximum state-wise $\Delta\phi$ along the trajectory instead of individual state-action transition $\Delta\phi$.
Next, by treating $\Delta\phi_{t}$ as an “imaginary” cost, we define a MMDP Zhao et al. (2023c) by introducing (i) an up-to-now maximum state-wise cost $M$ within $\mathcal{M}\subset\mathbb{R}$, and (ii) a "cost increment" function $D$, where $D:(\mathcal{S},\mathcal{M})\times\mathcal{A}\times\mathcal{S}\mapsto[0,\mathbb{R}^{+}]$ maps the augmented state-action transition tuple to non-negative cost increments. We define the augmented state $\hat{s}=(s,M)\in(\mathcal{S},\mathcal{M})\doteq\hat{{\mathcal{S}}}$, where $\hat{{\mathcal{S}}}$ is the augmented state space. Formally,
$D\big{(}\hat{s}_{t},a_{t},\hat{s}_{t+1}\big{)}=\max\{\Delta\phi(s_{t},a_{t},s_{t+1})-M,0\}.$ (8)
By setting $D\big{(}\hat{s}_{0},a_{0},\hat{s}_{1}\big{)}=\Delta\phi(s_{0},a_{0},s_{1})$, we have $M=\sum_{k=0}^{t-1}D\big{(}\hat{s}_{k},a_{k},\hat{s}_{k+1}\big{)}$ for $t\geq 1$. Hence, we define expected maximum state-wise cost (or $D$-return) for $\pi$:
$\mathcal{J}_{D}(\pi)=\mathbb{E}_{\tau\thicksim\pi}\Bigg{[}\sum_{t=0}^{H}D\big{(}\hat{s}_{t},a_{t},\hat{s}_{t+1}\big{)}\Bigg{]}.$ (9)
With (9), (7) can be rewritten as:
[…]
where $\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{H}\gamma^{t}R(\hat{s}_{t},a_{t},\hat{s}_{t+1})\right]$ and $R(\hat{s},a,\hat{s}^{\prime})\doteq R(s,a,s^{\prime})$. With $R(\tau)$ being the discounted return of a trajectory, we define the on-policy value function as $V^{\pi}(\hat{s})\doteq\mathbb{E}_{\tau\sim\pi}[R(\tau)|\hat{s}_{0}=\hat{s}]$, the on-policy action-value function as $Q^{\pi}(\hat{s},a)\doteq\mathbb{E}_{\tau\sim\pi}[R(\tau)|\hat{s}_{0}=\hat{s},a_{0}=a]$, and the advantage function as $A^{\pi}(\hat{s},a)\doteq Q^{\pi}(\hat{s},a)-V^{\pi}(\hat{s})$.
Lastly, we define on-policy value functions, action-value functions, and advantage functions for the cost increments in analogy to $V^{\pi}$, $Q^{\pi}$, and $A^{\pi}$, with $D$ replacing $R$, respectively. We denote those by $V_{D}^{\pi}$, $Q_{D}^{\pi}$ and $A_{D}^{\pi}$.
Remark 2.
Equation 7 is difficult to solve since there are as many constraints as the size of trajectory $\tau$. With (10), we turn all constraints in (7) into only a single constraint on the maximal $\Delta\phi$ along the trajectory, yielding a practically solvable problem.
4.2 S-3PO
To solve (10), we propose S-3PO inspired by recent trust region optimization methods Schulman et al. (2015). S-3PO uses KL divergence distance to restrict the policy search in (10) within a trust region around the most recent policy $\pi_{k}$. Moreover, S-3PO uses surrogate functions for the objective and constraints, which can be easily estimated from sample trajectories by $\pi_{k}$. Mathematically, S-3PO updates policy via solving the following optimization:
$\pi_{k+1}=\mathop{\underset{\pi\in\Pi_{\theta}}{\mathop{\mathbf{argmax}}}}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi_{k}}\\ a\sim\pi\end{subarray}}{\mathbb{E}}[A^{\pi_{k}}(\hat{s},a)]$ (11)
$\mathop{\mathbf{s.t.}}~{}~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta,$
$~{}~{}~{}~{}\mathcal{J}_{D}(\pi_{k})+\underset{\begin{subarray}{c}\hat{s}\sim\bar{d}^{\pi_{k}}\\ a\sim{\pi}\end{subarray}}{\mathbb{E}}\Bigg{[}A^{\pi_{k}}_{D}(\hat{s},a)\Bigg{]}+2(H+1)\epsilon_{D}^{\pi}\sqrt{\frac{1}{2}\delta}\leq 0.$
where $\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$ is KL divergence between two policy $(\pi^{\prime},\pi)$ at state $\hat{s}$, the set $\{\pi\in\Pi_{\theta}~{}:~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta\}$ is called trust region, $d^{\pi_{k}}\doteq(1-\gamma)\sum_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, $\bar{d}^{\pi_{k}}\doteq\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$ and $\epsilon^{{\pi}}_{D}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim{\pi}}[A^{\pi_{k}}_{D}(\hat{s},a)]|$.
Remark 3.
Despite the complex forms, the objective and constraints in (11) can be interpreted in two steps. First, maximizing the objective (expected reward advantage) within the trust region (marked by the KL divergence constraint) theoretically guarantees the worst performance degradation. Second, constraining the current cost advantage $A_{D}^{\pi_{k}}$ based on the previous value $\mathcal{J}_{D}(\pi_{k})$ guarantees that the worst-case “imagionary” cost is non-positive at all steps as in (7). In turn, the original safety constraint (4) is satisifed.
We then show in Section 6 that S-3PO achieves (i) state-wise safety guarantee of satisfying (3), and (ii) bounded worst case performance degradation for policy update, by establishing new bounds on the difference in returns between two stochastic policies $\pi$ and $\pi^{\prime}$ for MMDPs.
5 Practical Implementation
In this section, we summarize implementation techniques that helps with S-3PO’s practical performance. The pseudocode of S-3PO is give as algorithm 1.
Weighted loss for cost value targets
A critical step in S-3PO requires fitting of the cost increment value functions, denoted as $V_{D}^{\pi}(\hat{s}_{t})$. By definition, $V_{D}^{\pi}(\hat{s}_{t})$ is equal to the maximum cost increment in any future state over the maximal state-wise cost so far. In other words, $V_{D}^{\pi}(\hat{s}_{t})$ forms a non-increasing stair shape along the trajectory. Here we visualize an example of $V_{D}^{\pi}(\hat{s}_{t})$ in Figure 2. To enhance the accuracy of fitting this stair shape function, a weighted loss strategy is adopted, capitalizing on its monotonic property. Specifically, we define a weighted loss $L_{weight}$:
$\begin{split}L_{weight}&=L(\hat{y}_{t}-y_{t})*(1+w*\text{1}[(\hat{y}_{t}-y_{t-1})>0])\end{split}$
Figure 2: $V_{[D_{i}]\pi}(\hat{s})$ target of five sampled episodes.
where $L$ denotes Mean Squared Error (MSE), $\hat{y}_{t}$ is the prediction, ${y}_{t}$ is the fitting target and ${w}$ is the penalty weight. To account for the initial step (${t=0}$), we set ${y_{t-1}}$ to sufficiently large, thereby disregarding the weighted term associated with the first step. In essence, the rationale is to penalize any prediction that violates the non-increasing characteristics of the target sequence, thereby leading to an improved fitting quality.
Line search scheduling
Note that in (11), there are two constraints: (a) the trust region and (b) the bound on expected advantages. In practice, due to approximation errors, constraints in (11) might become infeasible. In that case, we perform a recovery update that only enforces the cost advantage $A_{D}^{\pi}$ to decrease starting from early training steps (in first $k_{\mathrm{safe}}$ updates), and starts to enforce reward improvements of $A^{\pi}$ towards the end of training. This is different from Zhao et al. (2023b), where the reward improvements are enforced at all times. This is because SCPO only guarantees safety (constraint satisfaction) after convergence, while S-3PO prioritizes constraining imaginary safety violation. With our line search scheduling, S-3PO is able to first grasp a safe policy, and then improve the reward performance. In that way, S-3PO achieves zero safety violation both during training and in testing with a worst-case performance degradation guarantee.
6 Theoretical Results
In this section, We present three theorems, including (i) zero violation exploration (Theorem 1), (ii) state-wise safety guarantee of S-3PO (Theorem 2), and (iii) worst case performance guarantee of S-3PO (Theorem 3). The proofs of the three theorems are summarized in Appendix C, Appendix D and Appendix E, respectively.
Theorem 1 (Zero Training Time Violation).
If the system satisfies Assumption 1, and the safety index design follows the rule described in (12), the implicit safe set algorithm ensures the system is forward invariant in $\mathcal{\bar{S}}\subset\mathcal{S}_{S}$.
Theorem 2 (Safety Guarantee of S-3PO).
If $\pi_{k},\pi_{k+1}$ are related by applying S-3PO and $s_{t}\in\mathcal{\bar{S}}$, then $s_{t+1}=f(s_{t},\pi_{k+1}(s_{t}))\in\mathcal{\bar{S}}$ in expectation.
Theorem 3 (Worst-case Performance Degradation in S-3PO.).
After $k_{safe}$ updates, suppose $\pi_{k},\pi_{k+1}$ are related by S-3PO update rules, with $\epsilon^{\pi_{k+1}}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim\pi_{k+1}}[A^{\pi_{k}}(\hat{s},a)]|$, then performance return for $\pi_{k+1}$ satisfies
$\mathcal{J}(\pi_{k+1})-\mathcal{J}(\pi_{k})\geq-\frac{\sqrt{2\delta}\gamma\epsilon^{\pi_{k+1}}}{1-\gamma}.$
Remark 4.
Theorem 1 shows zero safety violation exploration by proving the system state will never leave $\mathcal{\bar{S}}$ under the safeguard of ISSA. Theorem 2 shows that by satisfying the constraint of (11), the new policy is guaranteed to generate safe action, i.e. $a\in\mathcal{A}_{S}^{D}$, in expectation. Intuitively, Theorem 3 shows that with enough training epochs, the reward performance of S-3PO will not deteriorate too much after each update.
7 Experiments
In our experiments, we aim to answer the following questions:
Q1: Does S-3PO achieve zero-violation during the training?
Q2: How does S-3PO without safeguard compare with other state-of-the-art safe RL methods?
Q3: Does S-3PO learn to act without safeguard?
Q4: How does weighted loss trick impact the performance of S-3PO?
Q5: Is “imaginary” cost necessary to achieve zero violation?
Q6: How does S-3PO scale to high dimensional robots?
7.1 Experiments Setup
To answer these questions, we conducted our experiments on the safe reinforcement learning benchmark GUARD Zhao et al. (2023a) which is based on Mujoco and Gym interface.
(a) Point
(b) Swimmer
(c) Ant
(d) Hazard
(e) Pillar
Figure 3: Robots and constraints for benchmark problems in our environment.
Environment Setting
[…]
Our experiments revolve around four distinct robotic entities that can be categorized into two primary types:
  1. 1.
    Wheel Robot: this category encompasses robots that use wheels for locomotion, maintaining a continuous and seamless interaction with their surrounding environment. An example is the Point in fig. 3(a) which is designated as $\mathcal{A}\subseteq\mathbb{R}^{2}$.
  2. 2.
    Link Robot: this group includes robots composed of multiple connected links, which interact intermittently with their surroundings through the extremities of these links. Furthermore, the shape of these robots can undergo dynamic changes during these interactions. Specifically, in our environment, we have 1) Swimmer shown in fig. 3(b) as a three-link robot ($\mathcal{A}\subseteq\mathbb{R}^{2}$); 2) Ant shown in fig. 3(c) as a quadrupedal robot ($\mathcal{A}\subseteq\mathbb{R}^{8}$).
Two different types of constraints are considered.
  1. 1.
    Hazard: Dangerous areas shown in fig. 3(d). Hazards are trespassable circles on the ground. The agent is penalized for entering them.
  2. 2.
    Pillar: Fixed obstacles shown in  fig. 3(e). The agent is penalized for hitting them.
    More details about the experiments are discussed in Section F.1.
(a) Point_1Hazard
(b) Point_8Hazard
(c) Point_1Pillar
(d) Swimmer_1Hazard
Figure 5: Zero violation performance from four representative test suites in low dimensional systems (evaluated with the safeguard).
Comparison Group
The methods in the comparison group include: (i) unconstrained RL algorithm TRPO (Schulman et al., 2015) and TRPO-ISSA. (ii) end-to-end constrained safe RL algorithms CPO (Achiam et al., 2017), TRPO-Lagrangian (Bohez et al., 2019), PCPO (Yang et al., 2020), SCPO (Zhao et al., 2023b). (iii) We select TRPO as our baseline method since it is state-of-the-art and already has safety-constrained derivatives that can be tested off-the-shelf. For all experiments, the policy $\pi$, the value $(V^{\pi},V_{D}^{\pi})$ are all encoded in feedforward neural networks using two hidden layers of size (64,64) with tanh activations. The full list of parameters of all methods compared can be found in Section F.2.
Evaluation Metrics
For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost and (iii) cost rate. More details are provided in Section F.3. We set the limit of cost to 0 for all the safe RL algorithms since we aim to avoid any violation of the constraints.
(a) Point_1Hazard
(b) Point_8Hazard
(c) Point_1Pillar
(d) Swimmer_1Hazard
Figure 7: Results from four representative test suites in low dimensional systems (evaluated without the safeguard).
(a) Point_1Hazard
(b) Point_8Hazard
(c) Point_1Pillar
(d) Swimmer_1Hazard
Figure 9: Triggering frequency of the safeguard from four representative test suites in low dimensional systems.
7.2 Evaluating S-3PO and Comparison Analysis
Zero Violation During Training
The training performance of four representative test suites are summarized in Figure 5, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Theorem 1. For more experiments, please check Appendix F. This superior performance is attributed to the safeguard mechanism within the S-3PO framework, which effectively corrects unsafe actions at every step, particularly during training. Furthermore, as demonstrated in Figure 7, the reward performance continues to improve throughout the learning process, remaining comparable to state-of-the-art baselines. This distinct capability of S-3PO ensures safe reinforcement learning with zero safety violations, even in real-world scenarios, addressing Q1.
State-wise Safety Without Safety Monitor
To assess the performance of the S-3PO policy without the safeguard, evaluations are conducted at the end of each epoch. During these evaluations, the S-3PO policy is tested over 10,000 steps without the safeguard. This allows us to determine whether S-3PO effectively learns a state-wise safe policy through the guidance of the safe set-guided cost. The comparative results are presented in Figure 7. When contrasted with other baseline safe RL approaches, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance. These findings are consistent with Theorem 2 and Theorem 3, highlighting that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
Learn to Act without Safeguard
As highlighted in 1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation. To illustrate this, we visualize the average number of times the ISSA-based safeguard is triggered per step in Figure 9. For comparison, TRPO-ISSA is included as a baseline, which relies continuously on the safeguard to maintain safe control. Figure 9 shows that S-3PO dramatically reduces the frequency of safeguard activations, approaching zero, indicating that a state-wise safe policy has been effectively learned, thus addressing Q3.
Ablation on Weighted Loss for Fitting Cost Increment Value Targets
As pointed in Section 5, fitting $V_{D_{i}}(\hat{s}_{t})$ is a critical step towards solving S-3PO, which is challenging due to non-increasing stair shape of the target sequence. To elucidate the necessity of weighted loss for solving this challenge, we evaluate the cost rate of S-3PO under six distinct weight settings (0.0, 0.2, 0.4, 0.6, 0.8, 1.0) on Point_4Hazard test suite. The results shown in Figure 11(a) validates that a larger weight (hence higher penalty on predictions that violate the characteristics of value targets) results in better cost rate performance. This ablation study answers Q4.
(a) Comparison of cost rate performance with 6 different weights.
(b) Comparison between “imaginary” cost and action correction cost.
Figure 11: Comparison results of S-3PO
Necessity of “Imaginary” Cost
To understand the importance of the "imaginary" cost within the S-3PO framework, we compare it to another cost based on the magnitude of action correction Chen & Liu (2021). This empirical analysis is conducted using the Point_4Hazard test suite. As shown in Figure 11(b), the "imaginary" cost yields superior cost rate performance. This suggests that the "imaginary" cost offers deeper insights into the complex dynamics between the robot and its environment, thereby addressing Q5.
Figure 13: Cost and cost rate performance of Ant_1Hazard.
Scale S-3PO to High-Dimensional Link Robots
To showcase S-3PO’s scalability and performance with complex, high-dimensional link robots, we conducted additional tests on Ant_1Hazard featuring 8 dimensional control spaces. As shown in Figure 13, S-3PO effectively drives the cost to zero and rapidly reduces the cost rate, showcasing its clear superiority in high-dimensional safety policy learning and highlighting its exceptional scalability to more complex systems.
8 Conclusion and Future Prospectus
In this study, we introduce Safe Set Guided State-wise Constrained Policy Optimization (S-3PO), a novel algorithm pioneering state-wise safe optimal policies. This distinction is underlined by the absence of training violations, signifying an error-free learning paradigm. S-3PO employs a safeguard anchored in black-box dynamics to ensure secure exploration. Subsequently, it integrates a novel “imaginary” safety cost to guide the RL agent towards optimal safe policies. S-3PO outperforms existing methods in complex high-dimensional robotics tasks.
Nevertheless, a noteworthy limitation pertains to the potential costliness of acquiring a physical engine-based simulator (black-box dynamics model). A forward-looking perspective entails replacing the black-box dynamics model with a learned surrogate model, factoring in the nuances of errors in learned dynamics. This strategic move holds the promise of obliterating the final barrier impeding the seamless integration of safe RL training into real-world applications.
Appendix A Implicit Safe Set Algorithm Details
Figure 14: Notations.
A.1 Circumstances and Assumptions
In our treatment, both the robot and obstacles assume the form of point-mass circles, confined by finite collision radii. The safety criterion adopts the form $\phi_{0}=\max_{i}{\phi_{0}}_{i}$, with ${\phi_{0}}_{i}=d_{min}-d_{i}$. Here, $d_{i}$ captures the separation between the robot’s center and the $i$-th obstacle, encompassing both static and non-static entities. In this context, we introduce $z$ and $w$ to signify the relative acceleration and relative angular velocity of the robot, respectively, within the obstacle’s frame, as depicted in Figure 14. Importantly, the synthesis of the safety index for 2D collision evasion remains independent of specific dynamic models, but under the following assumption:
Assumption 1 (2D Collision Avoidance).
1) The state space is bounded, and the relative acceleration and angular velocity are bounded and both can achieve zeros, i.e., $w\in[w_{min},w_{max}]$ for $w_{min}\leq 0\leq w_{max}$ and $z\in[z_{min},z_{max}]$ for $z_{min}\leq 0\leq z_{max}$; 2) For all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$; 3) The discrete-time system time step $dt\to 0$; 4) At any given time, there can at most be one obstacle becoming safety critical, such that $\phi-\eta\geq 0$ (Sparse Obstacle Environment).
The bounds in the first assumption will be directly used to synthesize $\phi$. The second assumption enables us to turn the question on whether these exists a feasible control in $\mathcal{A}_{S}^{D}$ to the question on whether there exists $z$ and $w$ to decrease $\phi$. The third assumption ensures that the discrete time approximation error is small. The last assumption enables safety index design rule applicable with multiple moving obstacles.
A.2 Safety Index Synthesis
Following the rules in (Liu & Tomizuka, 2014), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$. Our goal is to choose these parameters such that $\mathcal{A}_{S}^{D}(s)$ is always nonempty. By setting $\eta=0$, the parameterization rule of safety index design rule is defined as:
$\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}\leq\frac{-z_{min}}{v_{max}}$ (12)
where $v_{max}$ is the maximum relative velocity that the robot can achieve in the obstacle frame. Note that the kinematic constraints $v_{max},z_{min}$ can be obtained by sampling the environment Zhao et al. (2021).
Appendix B Algorithms
B.1 S-3PO Algorithm
Algorithm 1 S-3PO
Input: Initial policy $\pi_{0}\in\Pi_{\theta}$.
for $k=0,1,2,\dots$ do
     for $t=0,1,2,\dots$ do
         Sample nominal action $a^{r}_{t}\sim\pi_{k}(s_{t})$
         Compute and execute $a_{t}=\mathrm{ISSA}(s_{t},a^{r}_{t})$
         Log $\tau\leftarrow\tau\cup\{(s_{t},a_{t},r_{t},s_{t+1},\Delta\phi_{t})\}$
     end for
     $g\leftarrow\left.\nabla_{\theta}\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi}(\hat{s},a)\right]\right\rvert_{\theta=\theta_{k}}$
     $b\leftarrow\left.\nabla_{\theta}\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi}_{D}(\hat{s},a)\right]\right\rvert_{\theta=\theta_{k}}$
     $c\leftarrow\mathcal{J}_{D}(\pi_{k})+2(H+1)\epsilon_{D}^{\pi}\sqrt{\delta/2}$
     $H\leftarrow\left.\nabla^{2}_{\theta}\mathbb{E}_{\hat{s}\sim\tau}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\right\rvert_{\theta=\theta_{k}}$
[…]
     ${}~{}~{}\frac{1}{2}(\theta-\theta_{k})^{\top}H(\theta-\theta_{k})\leq\delta,c+b^{\top}(\theta-\theta_{k})\leq 0$
     Get search direction $\Delta\theta^{*}\leftarrow\theta^{*}_{k+1}-\theta_{k}$
     for $j=0,1,2,\dots$ do $\triangleright$ Line search
         $\theta^{\prime}\leftarrow\theta_{k}+\xi^{j}\Delta\theta^{*}$
         if  $\mathbb{E}_{\hat{s}\sim\tau}[\mathcal{D}_{KL}(\pi_{\theta^{\prime}}\|\pi_{k})[\hat{s}]]\leq\delta$ and
$\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{\theta^{\prime}}}_{D}(\hat{s},a)-A^{\pi_{k}}_{D}(\hat{s},a)\right]\leq\mathrm{max}(-c,0)$ and
$(k\leq k_{\mathrm{safe}}~{}\mathrm{or}~{}\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{\theta^{\prime}}}(\hat{s},a)\right]\geq\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{k}}(\hat{s},a)\right])$ then
              $\theta_{k+1}\leftarrow\theta^{\prime}$triangleright Update policy
              break
         end if
     end for
end for
B.2 Additional Algorithms
The main body of AdamBA algorithm is summarized in Algorithm 2, and the main body of ISSA algorithm is summarized in Algorithm 3. The inputs for AdamBA are the approximation error bound ($\epsilon$), learning rate ($\beta$), reference control ($a^{r}$), gradient vector covariance ($\Sigma$), gradient vector number ($n$), reference gradient vector ($\vec{v}^{r}$), safety status of reference control ($S$), and the desired safety status of control solution ($S_{goal}$). The inputs for ISSA are the approximation error bound ($\epsilon$), the learning rate ($\beta$), gradient vector covariance ($\Sigma$), gradient vector number ($n$) and reference unsafe control ($a^{r}$).
Algorithm 2 Adaptive Momentum Boundary Approximation
1:procedure AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\vec{v}^{r},S,S_{goal}$)
2:     Initialize:
3:     if $\vec{v}^{r}$ is empty then
4:         Generate $n$ Gaussian distributed unit gradient vectors $\vec{v}_{i}\sim\mathcal{N}(0,\Sigma)$, $i=1,2,\ldots,n$
5:     else
6:         Initialize one unit gradient vector $\vec{v}_{1}=\frac{\vec{v}^{r}}{\|\vec{v}^{r}\|}$
7:     end if
8:     Approximation:
9:     for $i=1,2,\cdots,n$ do
10:         Initialize the approximated boundary point $P_{i}=a^{r}$, and stage = exponential outreach.
11:         while stage = exponential outreach do
12:              Set $P_{S}\leftarrow P_{i}$ and $P_{NS}\leftarrow P_{i}$
13:              $P_{i}=P_{i}+\vec{v}_{i}\beta$
14:              if $P_{i}$ is out of the control set then
15:                  break
16:              end if
17:              if $P_{i}$ safety status $\neq S$  then
18:                  Set $P_{NS}\leftarrow P_{i}$ , stage $\leftarrow$ exponential decay
19:                  break
20:              end if
21:              $\beta=2\beta$
22:         end while
23:         if stage = exponential decay then
24:              Apply Bisection method to locate boundary point until $\|P_{NS}-P_{S}\|<\epsilon$
25:              Set $P_{i}\leftarrow P_{S}$ if $S_{goal}=S$, $P_{i}\leftarrow P_{NS}$ otherwise
26:         end if
27:     end for
28:     Return Approximated Boundary Set $P$
29:end procedure
Algorithm 3 Implicit Safe Set Algorithm (ISSA)
1:procedure ISSA($\epsilon,\beta,\Sigma,n,a^{r}$)
2:     Phase 1: $\triangleright$ Phase 1
3:     Use AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\emptyset,\textit{UNSAFE},\textit{SAFE}$) to sample a collection $\mathbb{S}$ of safe control on the boundary of $\mathcal{A}_{S}^{D}$.
4:     if $\mathbb{S}=\emptyset$ then
5:         Enter Phase 2
6:     else
7:         For each primitive action $a_{i}\in\mathbb{S}$, compute the deviation $d_{i}=\|a_{i}-a^{r}\|^{2}$
8:         return $\mathop{\mathbf{argmin}}_{a_{i}}d_{i}$
9:     end if
10:
11:     Phase 2: $\triangleright$ Phase 2
12:     Use grid sampling by iteratively increasing sampling resolution to find an anchor safe control $a^{u}$, s.t. safety status of $a^{u}$ is SAFE.
13:     Use AdamBA($\epsilon,\frac{\|a^{r}-a^{u}\|}{4},\Sigma,1,a^{r},\frac{a^{u}-a^{r}}{\|a^{u}-a^{r}\|},\textit{UNSAFE},\textit{SAFE}$) to search for boundary point $a^{*}$
14:     if $a^{*}$ is not found then
15:         Use AdamBA($\epsilon,\frac{\|a^{r}-a^{u}\|}{4},\Sigma,1,a^{u},\frac{a^{r}-a^{u}}{\|a^{r}-a^{u}\|},\textit{SAFE},\textit{SAFE}$) to search for boundary point ${a^{u}}^{*}$
16:         Return ${a^{u}}^{*}$
17:     else
18:         Return $a^{*}$
19:     end if
20:end procedure
Appendix C Proof of Theorem 1
To prove the Theorem 1 we introduce two important corollaries to show that 1) the set of safe control is always nonempty if we choose a safety index that satisfies the design rule in (12); and 2) the proposed algorithm 3 is guaranteed to find a safe control if there exists one. With these two corollaries, it is then straightforward to prove the Theorem 1.
C.1 Feasibility of the Safety Index for Continuous-Time System
Corollary 1 (Non-emptiness of the set of safe control).
If 1) the dynamic system satisfies the conditions in Assumption 1; and 2) the safety index is designed according to the rule in Appendix A, then the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$.
Note that the set of safe control $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$ is non-empty if and only if it is non-empty in the following two cases: $\phi(s)-\eta<0$ or $\phi(s)-\eta\geq 0$. In the following discussion, we first show that the safety index design rule guarantees a non-empty set of safe control if there’s only one obstacle when $\phi(s)-\eta\geq 0$ (Lemma 1). Then we show that the set of safe control is non-empty if there’s only one obstacle when $\phi(s)-\eta<0$ (Lemma 2). Finally, we leverage Lemma 1 and Lemma 2 to show $\mathcal{A}_{S}^{D}(s)$ is non-empty if there’re multiple obstacles at any state.
Lemma 1.
If the dynamic system satisfies the conditions in Assumption 1 and there is only one obstacle in the environment, then the safety index design rule in Appendix A ensures that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for $x$ such that $\phi(s)-\eta\geq 0$.
Proof.
For $x$ such that $\phi(s)-\eta\geq 0$, the set of safe control becomes
$\mathcal{A}_{S}^{D}(s)=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\phi(s)-\eta\}$ (13)
According to the third condition in Assumption 1, we have $dt\rightarrow 0$. Therefore, the discrete-time approximation error approaches zero, i.e., $\phi(f(s,a))=\phi(s)+dt\cdot\dot{\phi}(s,a)+\Delta$, where $\Delta\rightarrow 0$. Then we can rewrite (13) as:
$\mathcal{A}_{S}^{D}(s)=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta/dt\}$ (14)
According to the definition of $\phi$, we have $\dot{\phi}=-nd^{n-1}\dot{d}-k\ddot{d}$. We ignored the subscript $i$ since it is assumed that there is only one obstacle. Therefore, the non-emptiness of $\mathcal{A}_{S}^{D}(s)$ in (14) is equivalent to the following condition
$\forall s\text{ s.t. }\phi(s)\geq\eta,\exists a,\text{s.t. }\ddot{d}\geq\frac{\eta/dt-nd^{n-1}\dot{d}}{k}.$ (15)
Note that in the 2D problem, $\ddot{d}=-z\cos(\alpha)+v\sin{(\alpha)}w$ and $\dot{d}=-v\cos(\alpha)$. According to Assumption 1, there is a surjection from $a$ to $(z,w)\in W:=\{(z,w)\mid z_{min}\leq z\leq z_{max},w_{min}\leq w\leq w_{max}\}$. Moreover, according to the definition of safety index, $\phi$ for the 2D problem only depends on $\alpha$, $v$, and $d$. Hence condition $\forall s\text{ s.t. }\phi(s)\geq\eta$ can be translated to $\forall(\alpha,v,d)\text{ s.t. }\sigma+d_{min}^{n}-d^{n}-kv\cos(\alpha)\geq\eta$. Denote the later set as
$\Phi:=\{(\alpha,v,d)\mid\sigma+d_{min}^{n}-d^{n}-kv\cos(\alpha)\geq\eta,v\in[0,v_{max}],d\geq 0,\alpha\in[0,2\pi)\}.$ (16)
Consequently, condition (6) is equivalent to the following condition
$\forall(\alpha,v,d)\in\Phi,\exists(z,w)\in W,\text{ s.t. }-z\cos(\alpha)+v\sin(\alpha)w$ $\geq\frac{\eta/dt+nd^{n-1}v\cos(\alpha)}{k}.$ (17)
According to the safety index design rule, we have $\eta=0$. Then we show (17) holds in different cases.
Case 1: $v=0$. In this case, we can simply choose $z=0$, then the inequality in (17) holds.
Case 2: $v\neq 0$ and $\cos(\alpha)\leq 0$. Note that velocity $v$ is always non-negative. Hence $v>0$. In this case, we just need to choose $z=w=0$, then the inequality in (17) holds, where the LHS becomes zero and the RHS becomes $\frac{nd^{n-1}v\cos(\alpha)}{k}$ which is non-positive.
Case 3: $v\neq 0$ and $\cos(\alpha)>0$. Dividing $v\cos(\alpha)$ on both sides of the inequality and rearranging the inequality, (17) is equivalent to
$\forall(\alpha,v,d)\in\Phi,\exists(z,w)\in W,\text{ s.t. }-\frac{z}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k}\geq 0,$ (18)
and (18) can be verified by showing:
$\min_{(\alpha,v,d)\in\Phi}\max_{(z,w)\in W}(-\frac{z}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k})\geq 0.$ (19)
Now let us expand the LHS of (19):
$\min_{(\alpha,v,d)\in\Phi}\max_{(z,w)\in W}(-\frac{z}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k})$ (20)
$=$ $\min_{(\alpha,v,d)\in\Phi}(-\frac{z_{min}}{v}+[\tan(\alpha)]_{+}w_{max}+[\tan(\alpha)]_{-}w_{min}-\frac{nd^{n-1}}{k})$ (21)
$=$ $\min_{\alpha\in[0,2\pi),v\in(0,v_{max}]}(-\frac{z_{min}}{v}+[\tan(\alpha)]_{+}w_{max}+[\tan(\alpha)]_{-}w_{min}-\frac{n(\sigma+d_{min}^{n}+kv\cos(\alpha))^{\frac{n-1}{n}}}{k})$ (22)
$=$ $-\frac{z_{min}}{v_{max}}-\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}.$ (23)
The first equality eliminates the inner maximization where $[\tan(\alpha)]_{+}:=\max\{\tan(\alpha),0\}$ and $[\tan(\alpha)]_{-}:=\min\{\tan(\alpha),0\}$. The second equality eliminates $d$ according to the constraint in $\Phi$. The third equality is achieved when $\alpha=0$ and $v=v_{max}$. According to the safety index design rule in Appendix A, (23) is greater than or equal to zero. Hence (19) holds, which then implies that the inequality in (17) holds.
The three cases cover all possible situations. Hence (17) always hold and the claim in the lemma is verified. ∎
Lemma 2.
If the dynamic system satisfies the contidions in Assumption 1 and there is only one obstacle in the environment, then the safety index design rule in Appendix A ensures that $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$ for any $s$ that $\phi(s)-\eta<0$.
Proof.
For $s$ such that $\phi(s)-\eta<0$, the set of safe control becomes
$\mathcal{A}_{S}^{D}(s)=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq 0\}$ (24)
According to the third assumption in Assumption 1, we have $dt\rightarrow 0$. Therefore, the discrete-time approximation error approaches zero, i.e., $\phi(f(s,a))=\phi(s)+dt\cdot\dot{\phi}(s,a)+\Delta$, where $\Delta\rightarrow 0$. Then we can rewrite (24) as:
$\mathcal{A}_{S}^{D}(s)=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\phi/dt\}$ (25)
Note that $\eta=0$ according to the safety index design rule, then $\phi(s)-\eta<0$ implies that $\phi(s)<0$. Hence $-\phi/dt\rightarrow\infty$ since $dt\rightarrow 0$. Then as long as $\dot{\phi}$ is bounded, $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$.
Now we show that $\dot{\phi}$ is bounded. According to the definition of safety index design rule, $\dot{\phi}=-nd^{n-1}\dot{d}-k\ddot{d}$. We ignored the subscript $i$ since it is assumed that there is only one obstacle. According to Assumption 1, we have the state space is bounded, thus both $d$ and $\dot{d}$ are bounded, which implies that $nd^{n-1}\dot{d}$ is bounded. Moreover, we have for all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$ according to Assumption 1, which indicates the mapping from $a$ to $(z,w)$ is surjective. Since $z$ and $w$ are bounded and both can achieve zeros according to Assumption 1, we have afor-$\forall a$, the corresponding $(z,w)$ are bounded. Since $\ddot{d}=-z\cos(\alpha)+v\sin(\alpha)w$, then $\ddot{d}$ is bounded. Hence $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$ any $s$ that $\phi(s)-\eta<0$ and the claim is true. ∎
Proof of Corollary 1.
Proof.
If there is one obstacle, then lemma 1 and lemma 2 have proved that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for all $s$. Now we need to consider the case where there are more than one obstacle but the environment is sparse in the sense that at any time step, there is at most one obstacle which is safety critical, i.e. $\phi_{i}\geq 0$. To show nonemptiness of $\mathcal{A}_{S}^{D}(s)$, we consider the following two cases. In the following discussion, we set $\eta=0$ according to the safety index design rule.
Case 1: $\phi(s)=\max_{i}\phi_{i}(s)\geq 0$. Denote $j:=\arg\max_{i}\phi_{i}(s)$. Since there is at most one obstacle that is safety critical, then $\phi_{j}(s)\geq 0$ and $\phi_{k}(s)<0$ for all $k\neq j$.
Denote ${\mathcal{A}_{S}^{D}}_{j}(s):=\{a\in\mathcal{A}\mid\phi_{j}(f(s,a))\leq\phi_{j}(s)\}$. Lemma 1 ensures that ${\mathcal{A}_{S}^{D}}_{j}(s)$ is nonempty. Denote ${\mathcal{A}_{S}^{D}}_{k}(s):=\{a\in\mathcal{A}\mid\phi_{k}(f(s,a))\leq 0\}$ where $k\neq j$. Since $\phi_{k}(s)<0$, lemma 2 ensures that ${\mathcal{A}_{S}^{D}}_{k}(s)=\mathcal{A}$.
Note that the set of safe control can be written as:
${\mathcal{A}_{S}^{D}}(s)$ :=$:=\{a\in\mathcal{A}\mid\max_{i}\phi_{i}(f(s,a))\leq\max_{i}\phi_{i}(s)\}$ (26a)
=$=\{a\in\mathcal{A}\mid\max_{i}\phi_{i}(f(s,a))\leq\phi_{j}(s)\}$ (26b)
$=\cap_{i}\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq\phi_{j}(s)\}$ (26c)
$={\mathcal{A}_{S}^{D}}_{j}(s)\neq\emptyset$ (26d)
Note that the last equality is due to the fact that $\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq\phi_{j}(s)\}\supseteq\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}=\mathcal{A}\supseteq{\mathcal{A}_{S}^{D}}_{j}(s)$ for $i\neq j$.
Case 2: $\phi(s)=\max_{i}\phi_{i}(s)<0$. Therefore, we have $\phi_{i}(s)<0$ for all $i$. According to Lemma 2, $\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}=\mathcal{A}$. Hence the set of safe control satisfies the following relationship
${\mathcal{A}_{S}^{D}}(s)$ :=$:=\{a\in\mathcal{A}\mid\max_{i}\phi_{i}(f(s,a))\leq 0\}$ (27a)
$=\cap_{i}\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}$ (27b)
$=\mathcal{A}\neq\emptyset$ (27c)
In summary, $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$ and the claim holds. ∎
C.2 Feasibility of ISSA
Corollary 2 (Feasibility of Algorithm 3).
If the set of safe control is non-empty, Algorithm 3 can always find a sub-optimal solution of (6) with a finite number of iterations.
Algorithm 3 executes two phases consecutively where the second phase will be executed if no solution of (6) is returned in the first phase. Hence, Algorithm 3 can always find a sub-optimal solution of (6) (safe control on the boundary of $\mathcal{A}_{S}^{D}$) if the solution of (6) can always be found in phase 2.
Note that Phase 2 first finds an anchor safe control $a^{u}$, then use it with AdamBA (Algorithm 2) to find the solution of (6). In the following discussion, we first show that the safety index design rule guarantees $a^{u}$ can be found with finite iterations (Lemma 3). Then we show that AdamBA will return a solution if it enters the exponential decay phase (Lemma 4). Subsequently, we show that the evoked AdamBA procedures in phase 2 will definitely enter exponential decay phase (Lemma 5). Finally, we provide a upper bound of the computation iterations for Algorithm 3.
Lemma 3 (Existence).
If the synthesized safety index can guarantee a non-empty set of safe control, then we can find an anchor point in phase 2 of Algorithm 3 with finite iterations (line 11 in algorithm 3).
Proof.
Figure 15: Illustration of the grid sampling to find an anchor control point.
If the synthesized safety index guarantees a non-empty set of safe control $\mathcal{A}^{D}_{S}$, then there exists a hypercube inside of $\mathcal{A}^{D}_{S}$, i.e. $\exists\mathcal{Q}\subset\mathcal{A}_{S}^{D}$, where $\mathcal{Q}$ is a $n_{u}$-dimensional hypercube with the same side length of $l>0$. Denote $\zeta_{[i]}=\max_{j,k}\|a_{[i]}^{j}-a_{[i]}^{k}\|$, where $a_{[i]}$ denotes the $i$-th dimension of control $a$, and $a^{j}\in\mathcal{A}^{D}_{S},a^{k}\in\mathcal{A}^{D}_{S}$.
By directly applying grid sampling in $\mathcal{A}^{D}_{s}$ with sample interval $l^{*}$ at each control dimension, such that $0<l^{*}<l$, the maximum sampling iteration $T^{a}$ for finding an anchor point in phase 2 satisfies the following condition:
$T^{a}<\prod\limits_{i=1}^{n_{u}}\lceil\frac{\zeta_{[i]}}{l^{*}}\rceil~{},$ (28)
where $T^{a}$ is a finite number since $l^{*}>0$. Then we have proved that we can find an anchor point in phase 2 of Algorithm 3 with finite iteration (i.e., finite sampling time). The grid sampling to find an anchor control point is illustrated in Figure 15.
Lemma 4 (Convergence).
If AdamBA enters the exponential decay phase, then it can always return a boundary point approximation (with desired safety status) where the approximation error is upper bounded by $\epsilon$.
Proof.
According to Algorithm 2, exponential decay phase applies Bisection method to locate the boundary point $P_{b}$ until $\|P_{NS}-P_{S}\|<\epsilon$. Denote the returned approximated boundary point as $P_{return}$, according to line 25, $P_{return}$ is either $P_{NS}$ or $P_{S}$, thus the approximation error satisfies:
$\|P_{return}-P_{b}\|$ $\leq\max\{\|P_{NS}-P_{b}\|,\|P_{S}-P_{b}\|\}$ (29)
$\leq\|P_{NS}-P_{S}\|$
$<\epsilon.$
Lemma 5 (Feasibility).
If we enter the phase 2 of Algorithm 3 with an anchor safe control being sampled, we can always find a local optimal solution of (6).
Proof.
According to line 13-15, after an anchor safe control is being sampled, phase 2 of Algorithm 3 will evoke at most two AdamBA processes. Hence, Lemma 5 can be proved by showing one of the two AdamBA will return a local optimal solution for (6). Next we show Lemma 5 holds in two cases.
Case 1:
line 13 of Algorithm 3 finds a solution.
In this case, the first AdamBA process finds a safe control $a^{u}$ solution (the return of AdamBA is a set, whereas the set here has at most one element). According to Algorithm 2, a solution will be returned only if AdamBA enters exponential decay stage. Hence, according to Lemma 4, $a^{u}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
Figure 16: Illustration of the case when it is unable to find $u^{*}$.
Case 2:
line 13 of Algorithm 3 fails to find a solution.
In this case, the second AdamBA process is evoked (line 15 of Algorithm 3). Since no solution is returned from the first AdamBA process (13 of Algorithm 3), where we start from $a^{r}$ and exponentially outreach along the direction $\vec{v}_{a}=\frac{a^{u}-a^{r}}{||a^{u}-a^{r}||}$, then all the searched control point along $\vec{v}_{a}$ is UNSAFE.
Specifically, we summarize the aforementioned scenario in Figure 16, the searched control points are represented as red dots along $\vec{v}_{a}$ (red arrow direction). Note that the exponential outreach starts with step size $\beta=\frac{\|a^{u}-a^{r}\|}{4}$, indicating two points $\{a_{s}^{1}$, $a_{s}^{2}\}$ are sampled between $a^{r}$ and $a^{u}$ such that
$\begin{cases}a_{s}^{1}=a^{r}+\frac{\|a^{u}-a^{r}\|}{4}\vec{v}_{a}=\frac{3a^{r}}{4}+\frac{a^{u}}{4}\\ a_{s}^{2}=a^{r}+\frac{3\|a^{u}-a^{r}\|}{4}\vec{v}_{a}=\frac{a^{r}}{4}+\frac{3a^{u}}{4}\end{cases}$ (30)
where the safety statuses of both $a_{s}^{1}$ and $a_{s}^{2}$ are UNSAFE.
During the second AdamBA process (line 15 of Algorithm 3), we start from $a^{u}$ and exponentially outreach along the direction $\vec{v}_{r}=\frac{a^{r}-a^{u}}{||a^{r}-a^{u}||}$ with initial step size $\beta=\frac{\|a^{r}-a^{u}\|}{4}$. Hence, denote $\bar{a}_{s}^{1}$ as the first sample point along $\vec{v}_{r}$, and $\bar{a}_{s}^{1}$ satisfies
$\bar{a}_{s}^{1}=a^{u}+\frac{\|a^{r}-a^{u}\|}{4}\vec{v}_{r}=\frac{a^{r}}{4}+\frac{3a^{u}}{4},$ (31)
which indicates $\bar{a}_{s}^{1}=a_{s}^{2}$, and the safety status of $\bar{a}_{s}^{1}$ is UNSAFE. Since the safe status of $a^{u}$ is SAFE and a UNSAFE point can be sampled during the exponential outreach stage, the second AdamBA process (line 15 of Algorithm 3) will enter exponential decay stage. Therefore, according to Lemma 4, ${a^{u}}^{*}$ will always be returned and ${a^{u}}^{*}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
The two cases cover all possible situations. Hence, after an anchor safe control $a^{u}$ is sampled, phase 2 of Algorithm 3 can always find a local optima of (6).
Proof of Corollary 2.
Proof.
According to Lemma 3 and Lemma 5, Algorithm 3 is able to find local optimal solution of (6). Next, we will prove Algorithm 3 can be finished within finite iterations.
According to Algorithm 3, ISSA include (i) one procedure to find anchor safe control $a_{a}$, and (ii) at most three AdamBA procedures. Firstly, based on Lemma 3, $a_{a}$ can be found within finite iterations. Secondly, each AdamBA procedure can be finished within finite iterations due to:
  • exponential outreach can be finished within finite iterations since the control space is bounded.
  • exponential decay can be finished within finite iterations since Bisection method will exit within finite iterations.
Therefore, ISSA can be finished within finite iterations.
C.3 Proof of Theorem 1
Before we prove Theorem 1, we first define set $\mathcal{S}_{S}^{D}:=\{s|\phi(s)\leq 0\}$, and we start with a preliminary result regarding $\mathcal{S}_{S}^{D}$ that is useful for proving the main theorem:
Lemma 6 (Forward Invariance of $\mathcal{S}_{S}^{D}$).
If the control system satisfies Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$.
Proof.
If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, then we can ensure the system has nonempty set of safe control at any state by Corollary 1. By Corollary 2, the implicit safe set algorithm can always find local optima solution of (6). The local optima solution always satisfies the constraint $\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}$, which indicates that 1) if $\phi(s_{t_{0}})\leq 0$, then $\phi(s_{t})\leq 0,\forall t\geq t_{0}$. Note that $\phi(s)\leq 0$ demonstrates that $s\in\mathcal{S}_{S}^{D}$. ∎
Proof the Theorem 1
Proof.
Leveraging Lemma 6, we then proceed to prove that the forward invariance to the set $\mathcal{S}_{S}^{D}$ guarantees the forward invariance to the set $\mathcal{\bar{S}}\subseteq\mathcal{S}_{S}$ which $\mathcal{\bar{S}}=\mathcal{S}_{S}\cap\mathcal{S}_{S}^{D}$. Depending on the relationship between $\mathcal{S}_{S}^{D}$ and $\mathcal{S}_{S}$, there are two cases in the proof which we will discuss below.
Case 1: $\mathcal{S}_{S}^{D}=\{s|\phi(s)\leq 0\}$ is a subset of $\mathcal{S}_{S}=\{s|\phi_{0}(s)\leq 0\}$.
In this case, $\mathcal{\bar{S}}=\mathcal{S}_{S}^{D}$. According to Lemma 6, If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$ and hence $\mathcal{\bar{S}}$.
Case 2: $\mathcal{S}_{S}^{D}=\{s|\phi(s)\leq 0\}$ is NOT a subset of $\mathcal{S}_{S}=\{s|\phi_{0}(s)\leq 0\}$.
In this case, if $s_{t}\in\mathcal{\bar{S}}$, we have $\phi_{0}(s_{t})=\max_{i}{\phi_{0}}_{i}(s_{t})\leq 0$, which indicates $\forall i,{\phi_{0}}_{i}\leq 0$.
Firstly, we consider the case where ${\phi_{0}}_{i}(s_{t})<0$. Note that ${\phi_{0}}_{i}(s_{t+1})={\phi_{0}}_{i}(s_{t})+\dot{\phi_{0}}_{i}(s_{t})dt+\frac{\ddot{\phi_{0}}_{i}(s_{t})dt^{2}}{2!}+\cdots$, since the state space and control space are both bounded, and $dt\rightarrow 0$ according to Assumption 1, we have ${\phi_{0}}_{i}(s_{t+1})\rightarrow{\phi_{0}}_{i}(s_{t})\leq 0$.
Secondly, we consider the case where ${\phi_{0}}_{i}(s_{t})=0$. Since $s_{t}\in\mathcal{\bar{S}}$, we have $\max_{i}\phi_{i}(s_{t})\leq 0$, which indicates $\forall i,\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}\leq 0$. Since ${\phi_{0}}_{i}(s_{t})=0$, we also have $d_{i}=d_{min}$. Therefore, the following condition holds:
$\sigma-k\dot{d}_{i}$ $\leq 0$ (32)
$\dot{d}_{i}$ $\geq\frac{\sigma}{k}$
According to the safety index design rule, we have $k,\sigma\in\mathbb{R}^{+}$, which indicates $\dot{d}_{i}>0$. Therefore, we have ${\phi_{0}}_{i}(s_{t+1})<0$.
Summarizing the above two cases, we have shown that if ${\phi_{0}}_{i}(s_{t})\leq 0$ then ${\phi_{0}}_{i}(s_{t+1})\leq 0$, which indicates if $\forall i,{\phi_{0}}_{i}(s_{t})\leq 0$ then $\forall i,{\phi_{0}}_{i}(s_{t+1})\leq 0$. Note that $\forall i,{\phi_{0}}_{i}(s_{t+1})\leq 0$ indicates that $\phi_{0}(s_{t+1})=\max_{i}{\phi_{0}}_{i}(s_{t+1})\leq 0$. Therefore, we have that if $s_{t}\in\mathcal{\bar{S}}$ then $s_{t+1}\in\mathcal{S}_{S}$. Thus, we also have $s_{t+1}\in\mathcal{\bar{S}}$ by Lemma 6. By induction, we have if $s_{t_{0}}\in\mathcal{\bar{S}}$, $s_{t}\in\mathcal{\bar{S}},\forall t>t_{0}$.
In summary, by discussing the two cases of whether $\mathcal{S}_{S}^{D}$ is the subset of $\mathcal{S}_{S}$, we have proven that if the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described in Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{\bar{S}}\subseteq\mathcal{S}_{S}$. Thus, if the initial state is safe, then the following state will always stay in $\mathcal{\bar{S}}$ which means ISSA can ensure the safety during training.
Appendix D Proof of Theorem 2
Mathematically, S-3PO requires the policy update at each iteration is bounded within a trust region, and updates policy via solving following optimization:
$\pi_{k+1}=\mathop{\underset{\pi\in\Pi_{\theta}}{\mathop{\mathbf{argmax}}}}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi_{k}}\\ a\sim\pi\end{subarray}}{\mathbb{E}}[A^{\pi_{k}}(\hat{s},a)]$ (33)
$~{}~{}~{}~{}~{}\mathop{\mathbf{s.t.}}~{}~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta,$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathcal{J}_{D}(\pi_{k})+\underset{\begin{subarray}{c}\hat{s}\sim\bar{d}^{\pi_{k}}\\ a\sim{\pi}\end{subarray}}{\mathbb{E}}\Bigg{[}A^{\pi_{k}}_{D}(\hat{s},a)\Bigg{]}+2(H+1)\epsilon_{D}^{\pi}\sqrt{\frac{1}{2}\delta}\leq 0$
where $\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$ is KL divergence between two policy $(\pi^{\prime},\pi)$ at state $\hat{s}$, the set $\{\pi\in\Pi_{\theta}~{}:~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta\}$ is called trust region, $d^{\pi_{k}}\doteq(1-\gamma)\sum_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, $\bar{d}^{\pi_{k}}\doteq\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, and $\epsilon^{{\pi}}_{D}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim{\pi}}[A^{\pi_{k}}_{D}(\hat{s},a)]|$. In practice, if ISSA is triggered, we take $D(\hat{s}_{t},a_{t},\hat{s}_{t+1})=max\{(\phi(f(\hat{s}_{t},a_{t}^{r}))-\phi(f(\hat{s}_{t},a_{t})))-M,0\}=max\{\Delta\phi_{t},0\}$. Otherwise we take $D(\hat{s}_{t},a_{t},\hat{s}_{t+1})=0$ directly.
D.1 Preliminaries
$\dot{d}^{\pi}$ we used is defined as
$\dot{d}^{\pi}(\hat{s})=\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi).$ (34)
which has a little difference with $d^{\pi}$ and is used to ensure the continuity of function we used for proof later. Then it allows us to express the expected discounted total reward or cost compactly as:
$\mathcal{J}_{g}(\pi)=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}\left[g(\hat{s},a,\hat{s}^{\prime})\right],$ (35)
where by $a\sim\pi$, we mean $a\sim\pi(\cdot|\hat{s})$, and by $\hat{s}^{\prime}\sim P$,we mean $\hat{s}^{\prime}\sim P(\cdot|\hat{s},a)$. $g(\hat{s},a,\hat{s}^{\prime})$ represents the cost or reward function. We drop the explicit notation for the sake of reducing clutter, but it should be clear from context that $a$ and $\hat{s}^{\prime}$ depend on $\hat{s}$.
[…]
$\dot{d}^{\pi}$ $=\sum\limits_{t=0}^{H}(\gamma P_{\pi})^{t}\hat{\mu}$ (36)
$=(I-(\gamma P_{\pi})^{H+1})(I-\gamma P_{\pi})^{-1}\hat{\mu}$
$=(I-\gamma P_{\pi})^{-1}\hat{\mu}$
Noticing that the finite MDP ends up at step $H$, thus $(P_{\pi})^{H+1}$ should be set to zero matrix.
This formulation helps us easily obtain the following lemma.
Lemma 7.
For any function $f:\hat{\mathcal{S}}\mapsto\mathbb{R}$ and any policy $\pi$,
$\underset{\hat{s}\sim\hat{\mu}}{\operatorname{E}}[f(\hat{s})]+\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\gamma f(\hat{s}^{\prime})]-\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}[f(\hat{s})]=0.$ (37)
Proof.
Multiply both sides of (36) by $(I-\gamma P_{\pi})$ and take the inner product with the vector $f\in\mathbb{R}^{|\hat{\mathcal{S}}|}$. ∎
Combining Lemma 7 with (35), we obtain the following, for any function $f$ and any policy $\pi$:
$\mathcal{J}_{g}(\pi)=\underset{\hat{s}\sim\hat{\mu}}{\operatorname{E}}\left[f(\hat{s})\right]+\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}\left[g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})\right]$ (38)
Lemma 8.
For any function $f\mapsto\mathbb{R}$ and any policies $\pi^{\prime}$ and $\pi$, define
$L_{\pi,f}(\pi^{\prime})\doteq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}\left[\left(\dfrac{\pi^{\prime}(a|\hat{s})}{\pi(a|\hat{s})}-1\right)(g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s}))\right],$ (39)
and $\epsilon_{f}^{\pi\prime}\doteq\max_{\hat{s}}|\mathrm{E}_{a\sim\pi^{\prime},\hat{s}^{\prime}\sim P}[g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})]|.$ Then the following bounds hold:
$\mathcal{J}_{g}(\pi^{\prime})-\mathcal{J}_{g}(\pi)\geq L_{\pi,f}(\pi^{\prime})-\epsilon_{f}^{\pi^{\prime}}\left\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\right\|_{1},$ (40)
$\mathcal{J}_{g}(\pi^{\prime})-\mathcal{J}_{g}(\pi)\leq L_{\pi,f}(\pi^{\prime})+\epsilon_{f}^{\pi^{\prime}}\left\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\right\|_{1},$ (41)
where $D_{TV}$ is the total variational divergence. Furthermore, the bounds are tight(when $\pi^{\prime}$ = $\pi$, the LHS and RHS are identically zero).
Proof.
First, for notational convenience, let $\delta_{f}(\hat{s},a,\hat{s}^{\prime})\doteq g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})$. By (38), we obtain the identity
$\mathcal{J}_{g}(\pi^{\prime})-\mathcal{J}_{g}(\pi)=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi^{\prime}}\\ a\sim{\pi^{\prime}}\\ \hat{s}^{\prime}\sim P\end{subarray}}{E}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]-\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{E}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$ (42)
Now, we restrict our attention to the first term in (42). Let $\dagger{\delta}_{f}^{\pi^{\prime}}\in\mathbb{R}^{|\hat{\mathcal{S}}|}$ denote the vector of components, where $\dagger{\delta}_{f}^{\pi^{\prime}}(\hat{s})=\mathbb{E}_{a\sim\pi^{\prime},\hat{s}^{\prime}\sim P}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})|\hat{s}]$. Observe that
$\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi^{\prime}}\\ a\sim{\pi^{\prime}}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$ $=\left\langle\dot{d}^{\pi^{\prime}},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle$
$=\left\langle\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle+\left\langle\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle$
With the Hölder’s inequality; for any $p,q\in[1,\infty]$ such that $\dfrac{1}{p}+\dfrac{1}{q}=1$, we have
$\left\langle\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle+\left\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\right\|_{p}\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{q}\geq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi^{\prime}}\\ a\sim{\pi^{\prime}}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]\geq\left\langle\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle-\left\|d^{\pi^{\prime}}-\dot{d}^{\pi}\right\|_{p}\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{q}$ (43)
We choose $p=1$ and $q=\infty$; With $\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{\infty}=\epsilon_{f}^{\pi^{\prime}}$, and by the importance sampling identity, we have
$\left\langle\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle$ $=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi^{\prime}}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$ (44)
$=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\left(\frac{\pi^{\prime}(a|\hat{s})}{\pi(a|\hat{s})}\right)\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$
After bringing (44), $\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{\infty}$ into (43), then substract $\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{E}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$, the bounds are obtained. The lower bound leads to (40), and the upper bound leads to (41). ∎
Lemma 9.
The divergence between discounted future state visitation distributions, $||\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}||_{1}$, is bounded by an average divergence of the policies $\pi^{\prime}$ and $\pi$:
$\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\|_{1}\leq 2\sum_{t=0}^{H}\gamma^{t+1}\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}\left[D_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right],$ (45)
where $D_{TV}(\pi^{\prime}||\pi)[\hat{s}]=\dfrac{1}{2}\sum_{a}|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|.$
Proof.
Firstly, we introduce an identity for the vector difference of the discounted future state visitation distributions on two different policies, $\pi^{\prime}$ and $\pi$. Define the matrices $G\doteq(I-\gamma P_{\pi})^{-1},\bar{G}\doteq(I-\gamma P_{\pi^{\prime}})^{-1}$, and $\Delta=P_{\pi^{\prime}}-P_{\pi}$. Then:
$G^{-1}-\bar{G}^{-1}$ $=(I-\gamma P_{\pi})-(I-\gamma P_{\pi^{\prime}})$ (46)
$=\gamma\Delta,$
left-multiplying by $G$ and right-multiplying by $\bar{G}$, we obtain
$\bar{G}-G=\gamma\bar{G}\Delta G.$ (47)
Thus, the following equality holds:
$\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}$ $=(1-\gamma)\left(\bar{G}-G\right)\hat{\mu}$ (48)
$=\gamma(1-\gamma)\bar{G}\Delta G\hat{\mu}$
$=\gamma\bar{G}\Delta\dot{d}^{\pi}.$
Using (48), we obtain
$\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\|_{1}$ $=\gamma\|\bar{G}\Delta d^{\pi}\|_{1}$ (49)
$\leq\gamma\|\bar{G}\|_{1}\|\Delta\dot{d}^{\pi}\|_{1},$
where $||\bar{G}||_{1}$ is bounded by:
$\|\bar{G}\|_{1}=\|(I-\gamma P_{\pi^{\prime}})^{-1}\|_{1}\leq\sum\limits_{t=0}^{\infty}\gamma^{t}\|P_{\pi^{\prime}}^{t}\|_{1}=\sum\limits_{t=0}^{H}\gamma^{t}.$ (50)
Next, we bound $\|\Delta\dot{d}^{\pi}_{1}\|$ as following:
$\|\Delta\dot{d}^{\pi}\|_{1}$ $=\quad\sum\limits_{\hat{s}^{\prime}}\left|\sum\limits_{\hat{s}}\Delta(\hat{s}^{\prime}|\hat{s})\dot{d}^{\pi}(\hat{s})\right|$ (51)
$\leq\quad\sum\limits_{\hat{s},\hat{s}^{\prime}}|\Delta(\hat{s}^{\prime}|\hat{s})|\dot{d}^{\pi}(\hat{s})$
$=\quad\sum_{\hat{s},\hat{s}^{\prime}}\left|\sum_{a}P(\hat{s}^{\prime}|\hat{s},a)\left(\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})\right)\right|\dot{d}^{\pi}(\hat{s})$
$\leq\quad\sum_{\hat{s},a,\hat{s}^{\prime}}P(\hat{s}^{\prime}|\hat{s},a)|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|\dot{d}^{\pi}(\hat{s})$
$=\quad\sum_{\hat{s},a}|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|\dot{d}^{\pi}(\hat{s})$
$=\quad 2\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}[D_{TV}(\pi^{\prime}||\pi)[\hat{s}]].$
By taking (51) and (50) into (49), this lemma is proved.
The new policy improvement bound follows immediately.
Lemma 10.
For any function $f:\hat{\mathcal{S}}\mapsto\mathbb{R}$ and any policies $\pi^{\prime}$ and $\pi$, define $\delta_{f}(\hat{s},a,\hat{s}^{\prime})\doteq g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})$,
$\begin{gathered}\epsilon_{f}^{\pi^{\prime}}\doteq\operatorname*{max}_{\hat{s}}|\mathrm{E}_{a\sim\pi^{\prime},\hat{s}^{\prime}\sim P}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]|,\\ L_{\pi,f}(\pi^{\prime})\doteq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{\mathrm{E}}\left[\left(\frac{\pi^{\prime}(a|\hat{s})}{\pi(a|\hat{s})}-1\right)\delta_{f}(\hat{s},a,\hat{s}^{\prime})\right],and\\ D_{\pi,f}^{\pm}(\pi^{\prime})\doteq L_{\pi,f}(\pi^{\prime})\pm 2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon_{f}^{\pi^{\prime}}\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}[D_{TV}(\pi^{\prime}||\pi)[\hat{s}]],\end{gathered}$
where $D_{TV}(\pi^{\prime}||\pi)[\hat{s}]=\frac{1}{2}\sum_{a}|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|$ is the total variational divergence between action distributions at $\hat{s}$. The following bounds hold:
$D_{\pi,f}^{+}(\pi^{\prime})\geq\mathcal{J}_{g}(\pi^{\prime})-\mathcal{J}_{g}(\pi)\geq D_{\pi,f}^{-}(\pi^{\prime}).$
Furthermore, the bounds are tight (when $\pi^{\prime}$ = $\pi$, all three expressions are identically zero)
Proof.
Begin with the bounds from lemma 8 and bound the divergence $D_{TV}(\dot{d}^{\pi^{\prime}}||\dot{d}^{\pi})$ by lemma 9. ∎
D.2 Safety invariance in expectation between S-3PO policies
Corollary 3 (SCPO Update Constraint Satisfaction).
Suppose $\mathcal{J}_{D}(\pi_{k})\leq 0$ and $\pi_{k},\pi_{k+1}$ are related by (33), then $D$-return for $\pi_{k+1}$ satisfies
$\forall i,\mathcal{J}_{D}(\pi_{k+1})\leq 0.$
Proof.
The choice of $f=\hat{V}_{D}^{\pi},g=D$ in lemma 10 leads to following inequality:
$\mathcal{\hat{J}}_{D}(\pi^{\prime})-\mathcal{\hat{J}}_{D}(\pi)\leq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[\hat{A}_{D}^{\pi}(\hat{s},a)+2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right].$ (52)
where $\mathcal{\hat{J}}_{D}(\pi)=\mathbb{E}_{\tau\thicksim\pi}\Bigg{[}\sum_{t=0}^{H}\gamma^{t}D\big{(}\hat{s}_{t},a_{t},\hat{s}_{t+1}\big{)}\Bigg{]}$, need to distinguish from $\mathcal{J}_{D}(\pi)$. And $\hat{V}_{D}^{\pi},\hat{A}_{D}^{\pi}$ are also the discounted version of $V_{D}^{\pi}$ and $A_{D}^{\pi}$. Note that according to Lemma 10 one can only get this the inequality holds when $\gamma\in(0,1)$.
Then we can define $\mathcal{F}(\gamma)=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[\hat{A}_{D}^{\pi}(\hat{s},a)+2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right]-\mathcal{\hat{J}}_{D}(\pi^{\prime})+\mathcal{\hat{J}}_{D}(\pi)$ with the following condition holds:
$\mathcal{F}(\gamma)\geq 0,\text{when}~{}\gamma\in(0,1)$ (53)
$\mathcal{F}(\gamma)\text{'s domain of definition is }\mathcal{R}$
$\mathcal{F}(\gamma)\text{ is a polynomial function}$
Because $\mathcal{F}(\gamma)$ is a polynomial function and coefficients are all limited, thus $\underset{\gamma\rightarrow 1^{-}}{\lim}\mathcal{F}(\gamma)$ exists and $\mathcal{F}(\gamma)$ is continuous at point $(1,\mathcal{F}(1)$). So $\mathcal{F}(1)=\underset{\gamma\rightarrow 1^{-}}{\lim}\mathcal{F}(\gamma)\geq 0$, which equals to:
$\mathcal{J}_{D}(\pi^{\prime})-\mathcal{J}_{D}(\pi)\leq\underset{\begin{subarray}{c}\hat{s}\sim\bar{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[A_{D}^{\pi}(\hat{s},a)+2(H+1)\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right].$
where $\bar{d}^{\pi}=\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|\pi)$. Thus, following the inequality (33), the Corollary 3 is proofed.
D.3 Proof of Theorem 2
If $s_{t}\in\mathcal{\bar{S}}\in\mathcal{S}_{S}^{D}$, then $\mathcal{J}_{D}(\pi_{t})\leq 0$ and $\phi(s_{t})\leq 0$. According to Corollary 3, we know that if policy $\pi_{t}$ is updated by solving (33), then we have $\mathcal{J}_{D}(\pi_{t+1})\leq 0$ which means:
$\underset{\tau\sim\pi_{t+1}}{\mathbb{E}}\left[\phi(s_{t+1})\right]=\mathcal{J}_{D}(\pi_{t+1})+\phi(s_{t})\leq 0$ (54)
Thus $s_{t+1}\in\mathcal{S}_{S}^{D}$ in expectation. According to the proof of Section C.3, we know that $s_{t+1}\in\mathcal{\bar{S}}\in\mathcal{S}_{S}^{D}$ in expectation.
Appendix E Proof of Theorem 3
E.1 KL Divergence Relationship Between $d^{\pi_{k}}$ and $\bar{d}^{\pi_{k}}$
Lemma 11.
$\underset{\hat{s}\sim d_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]<\underset{\hat{s}\sim\bar{d}_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]$
Proof.
$\underset{\hat{s}\sim d_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]$ $=\sum_{\hat{s}}(1-\gamma)\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi)\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$
$<\sum_{\hat{s}}\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi)\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$
$<\sum_{\hat{s}}\sum\limits_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|\pi)\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$
$=\underset{\hat{s}\sim\bar{d}_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]].$
E.2 Trust Region Update Performance
Lemma 12.
For any policies $\pi^{\prime},\pi$, with $\epsilon^{\pi^{\prime}}\doteq max_{\hat{s}}|E_{a\sim\pi^{\prime}}[A^{\pi}(\hat{s},a)]|$, and define $d^{\pi}=(1-\gamma)\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi)$ as the discounted augmented state distribution using $\pi$, then the following bound holds:
$\mathcal{J}(\pi^{\prime})-\mathcal{J}(\pi)>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim{\pi^{\prime}}\end{subarray}}{\mathbb{E}}\bigg{[}A^{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}\mathcal{D}_{TV}({\pi^{\prime}}\|\pi)[\hat{s}]\bigg{]}$ (55)
Proof.
The choice of $f=V_{\pi},g=R$ in lemma 10 leads to following inequality:
For any policies $\pi^{\prime},\pi$, with $\epsilon^{\pi^{\prime}}\doteq max_{\hat{s}}|E_{a\sim\pi^{\prime}}[A_{\pi}(\hat{s},a)]|$, the following bound holds:
$\mathcal{J}(\pi^{\prime})-\mathcal{J}(\pi)$ $\geq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathrm{E}}\left[A_{\pi}(\hat{s},a)-2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon^{\pi^{\prime}}D_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right]$
$>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathrm{E}}\left[A_{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}D_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right]$
At this point, the lemma 12 is proved.
E.3 Proof of Theorem 3
Proof.
If ISSA will constantly be triggered, the safeguard should be treated as a component of the environment, which indicates the environment is non-stationary for policy $\pi$ after safeguard is disabled. According to Algorithm 1, we assume that ISSA will not be triggered after $k_{safe}$ updates, meaning the environment is stationary for policy $\pi$ with or without safeguard. Therefore, we can infer S-3PO worst performance degradation after $k_{safe}$ updates following the trust region results of finite-horizon MDPs. Utilizing Lemma 12 and the relationship between the total variation divergence and the KL divergence, we have:
$\mathcal{J}(\pi^{\prime})-\mathcal{J}(\pi)>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\bigg{[}A^{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}\sqrt{\frac{1}{2}\mathbb{E}_{\hat{s}\sim d^{\pi}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]}\bigg{]}.$ (56)
In (33), the reward performance between two policies is associated with trust region, i.e.
$\pi_{k+1}=\mathop{\underset{\pi\in\Pi_{\theta}}{\mathop{\mathbf{argmax}}}}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi_{k}}\\ a\sim\pi\end{subarray}}{\mathbb{E}}[A^{\pi_{k}}(\hat{s},a)]$ (57)
$~{}~{}~{}~{}~{}\mathop{\mathbf{s.t.}}~{}~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta.$
Due to Lemma 11, if two policies are related with Equation 57, they are related with the following optimization:
$\pi_{k+1}=\mathop{\underset{\pi\in\Pi_{\theta}}{\mathop{\mathbf{argmax}}}}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi_{k}}\\ a\sim\pi\end{subarray}}{\mathbb{E}}[A^{\pi_{k}}(\hat{s},a)]$ (58)
$~{}~{}~{}~{}~{}\mathop{\mathbf{s.t.}}~{}~{}\mathbb{E}_{\hat{s}\sim d^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta.$
By (56) and (58), if $\pi_{k},\pi_{k+1}$ are related by (33), then performance return for $\pi_{k+1}$ satisfies
$\mathcal{J}(\pi_{k+1})-\mathcal{J}(\pi_{k})>-\frac{\sqrt{2\delta}\gamma\epsilon^{\pi_{k+1}}}{1-\gamma}.$
Appendix F Expeiment Details
F.1 Environment Settings
Goal Task
In the Goal task environments, the reward function is:
$\begin{split}&r(x_{t})=d^{g}_{t-1}-d^{g}_{t}+\mathbf{1}[d^{g}_{t}<R^{g}]~{},\\ \end{split}$
where $d^{g}_{t}$ is the distance from the robot to its closest goal and $R^{g}$ is the size (radius) of the goal. When a goal is achieved, the goal location is randomly reset to someplace new while keeping the rest of the layout the same. The test suites of our experiments are summarized in Table 1.
Hazard Constraint
In the Hazard constraint environments, the cost function is:
$\begin{split}&c(x_{t})=\max(0,R^{h}-d^{h}_{t})~{},\\ \end{split}$
where $d^{h}_{t}$ is the distance to the closest hazard and $R^{h}$ is the size (radius) of the hazard.
Pillar Constraint
In the Pillar constraint environments, the cost $c_{t}=1$ if the robot contacts with the pillar otherwise $c_{t}=0$.
Additional high dimensional link robot
To scale our method to high dimensional link robots. We additionally adopt Walker shown in 17 as a bipedal robot ($\mathcal{A}\subseteq\mathbb{R}^{10}$) in our experiments.
Figure 17: Walker
State Space
The state space is composed of two parts. The internal state spaces describe the state of the robots, which can be obtained from standard robot sensors (accelerometer, gyroscope, magnetometer, velocimeter, joint position sensor, joint velocity sensor and touch sensor). The details of the internal state spaces of the robots in our test suites are summarized in Table 2. The external state spaces are describe the state of the environment observed by the robots, which can be obtained from 2D lidar or 3D lidar (where each lidar sensor perceives objects of a single kind). The state spaces of all the test suites are summarized in Table 3. Note that Vase and Gremlin are two other constraints in Safety Gym Ray et al. (2019) and all the returns of vase lidar and gremlin lidar are zero vectors (i.e., $[0,0,\cdots,0]\in\mathbb{R}^{16}$) in our experiments since none of our test suites environments has vases.
Table 1: The test suites environments of our experiments
Task Setting Low dimension High dimension
Point Swimmer Walker Ant
Hazard-1
Hazard-4
Hazard-8
Pillar-1
Pillar-4
Pillar-8
Table 2: The internal state space components of different test suites environments.
Internal State Space Point Swimmer Walker Ant
Accelerometer ($\mathbb{R}^{3}$)
Gyroscope ($\mathbb{R}^{3}$)
Magnetometer ($\mathbb{R}^{3}$)
Velocimeter ($\mathbb{R}^{3}$)
Joint position sensor ($\mathbb{R}^{n}$) ${n=0}$ ${n=2}$ ${n=10}$ ${n=8}$
Joint velocity sensor ($\mathbb{R}^{n}$) ${n=0}$ ${n=2}$ ${n=10}$ ${n=8}$
Touch sensor ($\mathbb{R}^{n}$) ${n=0}$ ${n=4}$ ${n=2}$ ${n=8}$
Table 3: The external state space components of different test suites environments.
External State Space Goal-Hazard Goal-Pillar
Goal Compass ($\mathbb{R}^{3}$)
Goal Lidar ($\mathbb{R}^{16}$)
3D Goal Lidar ($\mathbb{R}^{60}$)
Hazard Lidar ($\mathbb{R}^{16}$)
3D Hazard Lidar ($\mathbb{R}^{60}$)
Pillar Lidar ($\mathbb{R}^{16}$)
Vase Lidar ($\mathbb{R}^{16}$)
Gremlin Lidar ($\mathbb{R}^{16}$)
Control Space
For all the experiments, the control space of all robots are continuous, and linearly scaled to [-1, +1].
F.2 Policy Settings
The hyper-parameters used in our experiments are listed in Table 4 as default.
Our experiments use separate multi-layer perception with ${tanh}$ activations for the policy network, value network and cost network. Each network consists of two hidden layers of size (64,64). All of the networks are trained using $Adam$ optimizer with learning rate of 0.01.
We apply an on-policy framework in our experiments. During each epoch the agent interact $B$ times with the environment and then perform a policy update based on the experience collected from the current epoch. The maximum length of the trajectory is set to 1000 and the total epoch number $N$ is set to 200 as default. In our experiments the Walker was trained for 250 epochs due to the high dimension.
The policy update step is based on the scheme of TRPO, which performs up to 100 steps of backtracking with a coefficient of 0.8 for line searching.
For all experiments, we use a discount factor of $\gamma=0.99$, an advantage discount factor $\lambda=0.95$, and a KL-divergence step size of $\delta_{KL}=0.02$.
For experiments which consider cost constraints we adopt a target cost $\delta_{c}=0.0$ to pursue a zero-violation policy.
Other unique hyper-parameters for each algorithms are hand-tuned to attain reasonable performance.
Each model is trained on a server with a 48-core Intel(R) Xeon(R) Silver 4214 CPU @ 2.2.GHz, Nvidia RTX A4000 GPU with 16GB memory, and Ubuntu 20.04.
Table 4: Important hyper-parameters of different algorithms in our experiments
Policy Parameter TRPO TRPO-ISSA TRPO-Lagrangian CPO PCPO SCPO S-3PO
Epochs $N$ 200 200 200 200 200 200 200
Steps per epoch $B$ 30000 30000 30000 30000 30000 30000 30000
Maximum length of trajectory $L$ 1000 1000 1000 1000 1000 1000
Policy network hidden layers (64, 64) (64, 64) (64, 64) (64, 64) (64, 64) (64, 64) (64, 64)
Discount factor $\gamma$ 0.99 0.99 0.99 0.99 0.99 0.99 0.99
Advantage discount factor $\lambda$ 0.97 0.97 0.97 0.97 0.97 0.97 0.97
TRPO backtracking steps 100 100 100 100 - 100 100
TRPO backtracking coefficient 0.8 0.8 0.8 0.8 - 0.8 0.8
Target KL $\delta_{KL}$ 0.02 0.02 0.02 0.02 0.02 0.02 0.02
Value network hidden layers (64, 64) (64, 64) (64, 64) (64, 64) (64, 64) (64, 64) (64, 64)
Value network iteration 80 80 80 80 80 80 80
Value network optimizer Adam Adam Adam Adam Adam Adam Adam
Value learning rate 0.001 0.001 0.001 0.001 0.001 0.001 0.001
Cost network hidden layers - - (64, 64) (64, 64) (64, 64) (64, 64) (64, 64)
Cost network iteration - - 80 80 80 80 80
Cost network optimizer - - Adam Adam Adam Adam Adam
Cost learning rate - - 0.001 0.001 0.001 0.001 0.001
Target Cost $\delta_{c}$ 0.0 0.0 0.0 0.0 0.0
Lagrangian learning rate - - 0.005 - - - -
Cost reduction - - - 0.0 - 0.0 0.0
F.3 Metrics Comparison
In Tables 5(c), 6(c) and 7, we report all the $9$ results of our test suites by three metrics:
  • The average episode return $J_{r}$.
  • The average episodic sum of costs $M_{c}$.
  • The average cost over the entirety of training $\rho_{c}$.
Both the evaluation performance and training performance are reported based on the above metrics. Besides, we also report the ISSA triger times as ISSA performance of TRPO-ISSA and S-3PO. All of the metrics were obtained from the final epoch after convergence. Each metric was averaged over two random seed. The evaluation performance curves of all experiments are shown in Figures 19, 25 and 31, the training performance curves of all experiments are shown in Figures 21, 27 and 33 and the ISSA performance curves of all experiments are shown in Figures 23, 29 and 35
Table 5: Metrics of three Point_Hazard environments obtained from the final epoch.
(a) Point_1Hazard
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.5738 0.5078 0.0082 2.5738 0.5078 0.0082 -
TRPO-Lagrangian 2.6313 0.5977 0.0058 2.6313 0.5977 0.0058 -
CPO 2.4988 0.1713 0.0045 2.4988 0.1713 0.0045 -
PCPO 2.4928 0.3765 0.0054 2.4928 0.3765 0.0054 -
SCPO 2.5457 0.0326 0.0022 2.5457 0.0326 0.0022 -
TRPO-ISSA 2.5113 0.0000 0.0000 2.5981 0.0000 0.0000 0.2714
S-3PO 2.4157 0.0000 0.0000 2.2878 0.0000 0.0000 0.0285
(b) Point_4Hazard
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.6098 0.2619 0.0037 2.6098 0.2619 0.0037 -
TRPO-Lagrangian 2.5494 0.2108 0.0034 2.5494 0.2108 0.0034 -
CPO 2.5924 0.1654 0.0024 2.5924 0.1654 0.0024 -
PCPO 2.5575 0.1824 0.0025 2.5575 0.1824 0.0025 -
SCPO 2.5535 0.0523 0.0009 2.5535 0.0523 0.0009 -
TRPO-ISSA 2.5014 0.0712 0.0000 2.5977 0.0135 0.0000 0.1781
S-3PO 2.3868 0.0000 0.0000 2.3550 0.0000 0.0000 0.0117
(c) Point_8Hazard
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.5535 0.5208 0.0074 2.5535 0.5208 0.0074 -
TRPO-Lagrangian 2.5851 0.5119 0.0064 2.5851 0.5119 0.0064 -
CPO 2.6440 0.2944 0.0041 2.6440 0.2944 0.0041 -
PCPO 2.6249 0.3843 0.0052 2.6249 0.3843 0.0052 -
SCPO 2.5126 0.0703 0.0020 2.5126 0.0703 0.0020 -
TRPO-ISSA 2.5862 0.0865 0.0000 2.5800 0.0152 0.0000 0.3431
S-3PO 2.4207 0.1710 0.0000 2.3323 0.0000 0.0000 0.0337
Table 6: Metrics of three Pillar_Hazard environments obtained from the final epoch.
(a) Point_1Pillar
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.6065 0.2414 0.0032 2.6065 0.2414 0.0032 -
TRPO-Lagrangian 2.5772 0.1218 0.0020 2.5772 0.1218 0.0020 -
CPO 2.5464 0.2342 0.0028 2.5464 0.2342 0.0028 -
PCPO 2.5857 0.2088 0.0025 2.5857 0.2088 0.0025 -
SCPO 2.5928 0.0040 0.0003 2.5928 0.0040 0.0003 -
TRPO-ISSA 2.5985 0.0000 0.0000 2.5909 0.0020 0.0000 0.3169
S-3PO 2.5551 0.0000 0.0000 2.5241 0.0000 0.0000 0.0060
(b) Point_4Pillar
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.5671 0.4112 0.0063 2.5671 0.4112 0.0063 -
TRPO-Lagrangian 2.6040 0.2786 0.0050 2.6040 0.2786 0.0050 -
CPO 2.5720 0.5523 0.0062 2.5720 0.5523 0.0062 -
PCPO 2.5709 0.3240 0.0052 2.5709 0.3240 0.0052 -
SCPO 2.5367 0.0064 0.0005 2.5367 0.0064 0.0005 -
TRPO-ISSA 2.5739 0.1198 0.0001 2.5881 0.0427 0.0001 0.2039
S-3PO 2.2513 0.0114 0.0000 2.3459 0.0000 0.0000 0.0116
(c) Point_8Pillar
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.6140 3.1552 0.0201 2.6140 3.1552 0.0201 -
TRPO-Lagrangian 2.6164 0.6632 0.0129 2.6164 0.6632 0.0129 -
CPO 2.6440 0.5655 0.0166 2.6440 0.5655 0.0166 -
PCPO 2.5704 6.6251 0.0219 2.5704 6.6251 0.0219 -
SCPO 2.4162 0.2589 0.0024 2.4162 0.2589 0.0024 -
TRPO-ISSA 2.6203 0.6910 0.0009 2.5921 0.0709 0.0009 0.3517
S-3PO 2.0325 0.0147 0.0002 2.3371 0.0000 0.0002 0.0231
Table 7: Metrics of three link robots environments obtained from the final epoch.
(a) Swimmer_1Hazard
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.7049 0.3840 0.0076 2.7049 0.3840 0.0076 -
TRPO-Lagrangian 2.6154 0.3739 0.0060 2.6154 0.3739 0.0060 -
CPO 2.5817 0.3052 0.0056 2.5817 0.3052 0.0056 -
PCPO 2.5418 0.6243 0.0055 2.5418 0.6243 0.0055 -
SCPO 2.6432 0.3919 0.0051 2.6432 0.3919 0.0051 -
TRPO-ISSA 2.5826 0.2595 0.0000 2.5955 0.0000 0.0000 0.1240
S-3PO 2.6032 0.0313 0.0000 2.6239 0.0001 0.0000 0.0378
(b) Ant_1Hazard
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.6390 0.3559 0.0074 2.6390 0.3559 0.0074 -
TRPO-Lagrangian 2.5866 0.2169 0.0044 2.5866 0.2169 0.0044 -
CPO 2.6175 0.2737 0.0072 2.6175 0.2737 0.0072 -
PCPO 2.6103 0.2289 0.0076 2.6103 0.2289 0.0076 -
SCPO 2.6341 0.2384 0.0065 2.6341 0.2384 0.0065 -
TRPO-ISSA 2.6509 0.3831 0.0032 2.6318 0.3516 0.0032 0.0279
S-3PO 2.2047 0.0000 0.0002 2.2031 0.0000 0.0002 0.0001
(c) Walker_1Hazard
Algorithm Evaluation Performance Training Performance ISSA Performance
$\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$ $\bar{J}_{r}$ $\bar{M}_{c}$ $\bar{\rho}_{c}$
TRPO 2.5812 0.2395 0.0075 2.5812 0.2395 0.0075 -
TRPO-Lagrangian 2.6227 0.1666 0.0041 2.6227 0.1666 0.0041 -
CPO 2.6035 0.3068 0.0062 2.6035 0.3068 0.0062 -
PCPO 2.5775 0.2414 0.0059 2.5775 0.2414 0.0059 -
SCPO 2.6352 0.1423 0.0051 2.6352 0.1423 0.0051 -
TRPO-ISSA 2.6419 0.3544 0.0037 2.5787 0.2060 0.0037 0.0316
S-3PO 2.6117 0.3437 0.0025 2.6055 0.2665 0.0025 0.0319
(a) Point_1Hazard
(b) Point_4Hazard
(c) Point_8Hazard
Figure 19: Evaluation performance of Point_Hazard
(a) Point_1Hazard
(b) Point_4Hazard
(c) Point_8Hazard
Figure 21: Training performance of Point_Hazard
(a) Point_1Hazard
(b) Point_4Hazard
(c) Point_8Hazard
Figure 23: ISSA performance of Point_Hazard
(a) Point_1Pillar
(b) Point_4Pillar
(c) Point_8Pillar
Figure 25: Evaluation performance of Point_Pillar
(a) Point_1Pillar
(b) Point_4Pillar
(c) Point_8Pillar
Figure 27: Training performance of Point_Pillar
(a) Point_1Pillar
(b) Point_4Pillar
(c) Point_8Pillar
Figure 29: ISSA performance of Point_Pillar
(a) Swimmer_1Hazard
(b) Ant_1Hazard
(c) Walker_1Hazard
Figure 31: Evaluation performance of link robots
(a) Swimmer_1Hazard
(b) Ant_1Hazard
(c) Walker_1Hazard
Figure 33: Training performance of link robots
(a) Swimmer_1Hazard
(b) Ant_1Hazard
(c) Walker_1Hazard
Figure 35: ISSA performance of link robots

### v4

@@
 Meanwhile, safe control methods to continuously meet stringent safety requirements in predictable environments are widely examined, with energy function-based methods being the most prevalent Khatib (1986); Ames et al.
-(2014); Liu & Tomizuka (2014); Gracia et al.
-(2013); Wei & Liu (2019).
+(2014); Liu and Tomizuka (2014); Gracia et al.
+(2013); Wei and Liu (2019).
 These methods establish energy functions assigning lower energy to safe states and project nominal control into energy dissipating control, hence persistently maintaining system safety.
@@
 The Implicit Safe Set Algorithm (ISSA) Zhao et al.
-(2021) addresses this problem using black-box optimization over black-box dynamics models.
+(2021, 2024) addresses this problem using black-box optimization over black-box dynamics models.
 Nevertheless, the reliance on continuous monitoring to ensure safety imposes significant computational overhead, and the RL policy itself does not inherently guarantee safe behavior.
@@
 3.2 Implicit Safe Set Algorithm
-Energy function-based methods  Wei & Liu (2019) achieve safe control by designing an energy function offline such that 1) the low energy states are safe and 2) there always exists a feasible control input to dissipate the energy.
-A typical design Liu & Tomizuka (2014) was proposed as $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$.
-It is shown in Liu & Tomizuka (2014) that if the control input is unbounded ($\mathcal{A}=\mathbb{R}^{n_{a}}$), then there always exist a control input that satisfies the constraint $\dot{\phi}\leq 0\text{ when }\phi=0$; and if the control input always satisfies that constraint, then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant.
-In practice, the actual control signal is computed through a quadratic projection of the nominal control $a^{r}$ to the control constraint
-$a=$ $\mathop{\mathbf{argmin}}_{a\in\mathcal{A}}\|a-a^{r}\|^{2}\text{ s.t. }\dot{\phi}\leq-\eta(\phi),$ (5)
-where $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$ according to SSA Liu & Tomizuka (2014).
-Here, we define the set of safe control as $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$.
-To match the notion of MDP, we should consider discrete-time safe control set $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$.
-Based on these, the implicit safe set algorithm (ISSA) Zhao et al.
-(2021) was proposed to construct the black-box dynamics based safeguard ensuring persistent satisfication of safety specification.
-The black-box dynamics can be a digital twin or neural network.
-Under some mild assumptions, it first synthesizes a safety index to make sure $\mathcal{A}_{S}(s)$ is nonempty for all $s$ (details of the mild assumptions and safety index synthesis are summarized in Appendix A).
-Then it manages to project the reference action generated from RL policy $\pi_{\theta}$ into $\mathcal{A}_{S}(s)$ during policy training.
-In detail, the nominal control $a_{t}^{r}$ needs to be projected to $\mathcal{A}_{S}^{D}(s_{t})$ by solving the following optimization:
+The Implicit Safe Set Algorithm (ISSA) Zhao et al.
+(2021, 2024) ensures the persistent satisfaction of safety specifications for systems with black-box dynamics (e.g., digital twins or neural networks) through energy function-based optimization.
+Leveraging energy function $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$ and theoretical results from SSA Liu and Tomizuka (2014), ISSA synthesizes a safety index to guarantee that the safe control set $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$ is nonempty.
+And then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant under mild assumptions.
+Here $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$.
+Intuitively, the energy function $\phi$ ensures that low-energy states are safe, while the control signal within $\mathcal{A}_{S}(s)$ reduces $\phi$ whenever the system state is outside the safe region.
+This further derives the forward invariance of $\bar{\mathcal{S}}$, meaning that once the system enters the safe region, its state will stays within the safe part of the state space.
+To match the notion of MDP, we consider discrete-time safe control set $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$.
+Thus ISSA ensures safety by projecting the nominal control $a_{t}^{r}$ generated from RL policy $\pi_{\theta}$ to $\mathcal{A}_{S}^{D}(s_{t})$ through solving the following optimization:
 $\begin{split}&\min_{a_{t}\in\mathcal{A}}\|a_{t}-a^{r}_{t}\|^{2}\\ &\text{s.t. }\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}.\end{split}$ (5)
-To solve (6), they first introduce a sample-efficient Adaptive Momentum Boundary Approximation (AdamBA) algorithm.
-Then, ISSA directly uses it to find the safe control with minimum deviation from the reference control along the boundaries of $\mathcal{A}_{S}^{D}(s)$.
-If the process fails to return a solution, grid sampling will be deployed to find a safe control; and AdamBA is deployed again to improve the solution optimality with respect to (6).
-We summarize details of AdamBA and ISSA in Algorithm 2 and Algorithm 3, respectively.
+The detailed mild assumoptions and safety index synthesis are summarized in Appendix A. And the algorithms for solving (5) are provided in Algorithms 2 and 3.
 3.3 State-wise Constrained Policy Optimization
@@
 The critical observation we rely upon is that:
-Observation 1.
+Observation 1
 Define $\Delta\phi_{t}=\Delta\phi(s_{t},a_{t},s_{t+1})\doteq\phi(f(s_{t},a^{r}_{t}))-\phi(f(s_{t},a_{t}))$, i.e.
@@
 $\max_{\theta}\mathcal{J}(\pi_{\theta}),\text{ s.t. }\pi_{\theta}\in\{\pi\in\Pi\big{|}~{}\forall\Delta\phi_{t}\sim\tau,\Delta\phi_{t}\leq 0\}.$ (6)
-Remark 1.
+Remark 4.1.
 Policies satisfying (6) ensure there is no imaginary safety violation for any possible $a_{t}^{r}$, making $\pi_{\theta}$ a safe policy as required by (4).
@@
 We denote those by $V_{D}^{\pi}$, $Q_{D}^{\pi}$ and $A_{D}^{\pi}$.
-Remark 2.
+Remark 4.2.
 Equation 6 is difficult to solve since there are as many constraints as the size of trajectory $\tau$.
@@
 $~{}~{}~{}~{}\mathcal{J}_{D}(\pi_{k})+\underset{\begin{subarray}{c}\hat{s}\sim\bar{d}^{\pi_{k}}\\ a\sim{\pi}\end{subarray}}{\mathbb{E}}\Bigg{[}A^{\pi_{k}}_{D}(\hat{s},a)\Bigg{]}+2(H+1)\epsilon_{D}^{\pi}\sqrt{\frac{1}{2}\delta}\leq 0.$
-where $\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$ is KL divergence between two policy $(\pi^{\prime},\pi)$ at state $\hat{s}$, the set $\{\pi\in\Pi_{\theta}~{}:~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta\}$ is called trust region, $d^{\pi_{k}}\doteq(1-\gamma)\sum_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, $\bar{d}^{\pi_{k}}\doteq\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$ and $\epsilon^{{\pi}}_{D}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim{\pi}}[A^{\pi_{k}}_{D}(\hat{s},a)]|$.
-Remark 3.
+where $\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$ is KL divergence between two policy $(\pi^{\prime},\pi)$ at state $\hat{s}$, the set $\{\pi\in\Pi_{\theta}~{}:\\ ~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta\}$ is called trust region, $d^{\pi_{k}}\doteq(1-\gamma)\sum_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, $\bar{d}^{\pi_{k}}\doteq\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$ and $\epsilon^{{\pi}}_{D}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim{\pi}}[A^{\pi_{k}}_{D}(\hat{s},a)]|$.
+Remark 4.3.
 Despite the complex forms, the objective and constraints in (10) can be interpreted in two steps.
@@
 We then show in Section 5 that S-3PO achieves (i) state-wise safety guarantee of satisfying (3), and (ii) bounded worst case performance degradation for policy update, by establishing new bounds on the difference in returns between two stochastic policies $\pi$ and $\pi^{\prime}$ for MMDPs.
-5 Practical Implementation
-In this section, we summarize implementation techniques that helps with S-3PO’s practical performance. The pseudocode of S-3PO is give as algorithm 1.
-Weighted loss for cost value targets
-A critical step in S-3PO requires fitting of the cost increment value functions, denoted as $V_{D}^{\pi}(\hat{s}_{t})$.
-By definition, $V_{D}^{\pi}(\hat{s}_{t})$ is equal to the maximum cost increment in any future state over the maximal state-wise cost so far.
-In other words, $V_{D}^{\pi}(\hat{s}_{t})$ forms a non-increasing stair shape along the trajectory.
-Here we visualize an example of $V_{D}^{\pi}(\hat{s}_{t})$ in Figure 2.
-To enhance the accuracy of fitting this stair shape function, a weighted loss strategy is adopted, capitalizing on its monotonic property.
-Specifically, we define a weighted loss $L_{weight}$:
-$\begin{split}L_{weight}&=L(\hat{y}_{t}-y_{t})*(1+w*\text{1}[(\hat{y}_{t}-y_{t-1})>0])\end{split}$
-Figure 2: $V_{[D_{i}]\pi}(\hat{s})$ target of five sampled episodes.
-where $L$ denotes Mean Squared Error (MSE), $\hat{y}_{t}$ is the prediction, ${y}_{t}$ is the fitting target and ${w}$ is the penalty weight.
-To account for the initial step (${t=0}$), we set ${y_{t-1}}$ to sufficiently large, thereby disregarding the weighted term associated with the first step.
-In essence, the rationale is to penalize any prediction that violates the non-increasing characteristics of the target sequence, thereby leading to an improved fitting quality.
-Line search scheduling
-Note that in (11), there are two constraints: (a) the trust region and (b) the bound on expected advantages.
-In practice, due to approximation errors, constraints in (11) might become infeasible.
-In that case, we perform a recovery update that only enforces the cost advantage $A_{D}^{\pi}$ to decrease starting from early training steps (in first $k_{\mathrm{safe}}$ updates), and starts to enforce reward improvements of $A^{\pi}$ towards the end of training.
-This is different from Zhao et al.
-(2023b), where the reward improvements are enforced at all times.
-This is because SCPO only guarantees safety (constraint satisfaction) after convergence, while S-3PO prioritizes constraining imaginary safety violation.
-With our line search scheduling, S-3PO is able to first grasp a safe policy, and then improve the reward performance.
-In that way, S-3PO achieves zero safety violation both during training and in testing with a worst-case performance degradation guarantee.
+4.3 Practical Implementation
+The pseudocode of S-3PO is give as algorithm 1.
+Here we summarize two techniques that helps with S-3PO’s practical performance.
+(i) Weighted loss for cost value targets: A critical step in S-3PO involves fitting the cost increment value function, $V_{D}^{\pi}(\hat{s}_{t})$, which represents the maximum future cost increment relative to the highest state-wise cost observed so far.
+This function follows a non-increasing staircase pattern along the trajectory.
+Thus, we adopt a weighted loss function, $L_{weight}$, to penalize predictions that violate the non-increasing property: $L_{weight}=L(\hat{y}_{t}-y_{t})*(1+w*\text{1}[(\hat{y}_{t}-y_{t-1})>0])$, where $L$ denotes Mean Squared Error, $\hat{y}_{t}$ is the prediction, ${y}_{t}$ is the fitting target and ${w}$ is the penalty weight.
+(ii) Line Search scheduling: Constraints in (10) might become infeasible due to approximation errors.
+In this case, we perform a recovery update, enforcing the cost advantage $A_{D}^{\pi}$ to decrease from early training steps $k_{\mathrm{safe}}$ while focusing on reward improvements of $A^{\pi}$ towards the end of training, prioritizing safety first and reward performance later.
+Please check Section F.3 for furthur details.
 5 Theoretical Results
-In this section, We present three theorems, including (i) zero violation exploration (Theorem 1), (ii) state-wise safety guarantee of S-3PO (Theorem 2), and (iii) worst case performance guarantee of S-3PO (Theorem 3).
+In this section, We present three theorems, including (i) zero violation exploration (Theorem 5.1), (ii) state-wise safety guarantee of S-3PO (Theorem 5.2), and (iii) worst case performance guarantee of S-3PO (Theorem 5.3).
 The proofs of the three theorems are summarized in Appendix C, Appendix D and Appendix E, respectively.
-Theorem 1 (Zero Training Time Violation).
+Theorem 5.1 (Zero Training Time Violation).
 If the system satisfies Assumption 1, and the safety index design follows the rule described in (11), the implicit safe set algorithm ensures the system is forward invariant in $\mathcal{\bar{S}}\subset\mathcal{S}_{S}$.
-Theorem 2 (Safety Guarantee of S-3PO).
+Theorem 5.2 (Safety Guarantee of S-3PO).
 If $\pi_{k},\pi_{k+1}$ are related by applying S-3PO and $s_{t}\in\mathcal{\bar{S}}$, then $s_{t+1}=f(s_{t},\pi_{k+1}(s_{t}))\in\mathcal{\bar{S}}$ in expectation.
-Theorem 3 (Worst-case Performance Degradation in S-3PO.).
+Theorem 5.3 (Worst-case Performance Degradation in S-3PO.).
 After $k_{safe}$ updates, suppose $\pi_{k},\pi_{k+1}$ are related by S-3PO update rules, with $\epsilon^{\pi_{k+1}}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim\pi_{k+1}}[A^{\pi_{k}}(\hat{s},a)]|$, then performance return for $\pi_{k+1}$ satisfies
 $\mathcal{J}(\pi_{k+1})-\mathcal{J}(\pi_{k})\geq-\frac{\sqrt{2\delta}\gamma\epsilon^{\pi_{k+1}}}{1-\gamma}.$
-Remark 4.
-Theorem 1 shows zero safety violation exploration by proving the system state will never leave $\mathcal{\bar{S}}$ under the safeguard of ISSA.
-Theorem 2 shows that by satisfying the constraint of (11), the new policy is guaranteed to generate safe action, i.e.
+Remark 5.4.
+Theorem 5.1 shows zero safety violation exploration by proving the system state will never leave $\mathcal{\bar{S}}$ under the safeguard of ISSA.
+Theorem 5.2 shows that by satisfying the constraint of (10), the new policy is guaranteed to generate safe action, i.e.
 $a\in\mathcal{A}_{S}^{D}$, in expectation.
-Intuitively, Theorem 3 shows that with enough training epochs, the reward performance of S-3PO will not deteriorate too much after each update.
+Intuitively, Theorem 5.3 shows that with enough training epochs, the reward performance of S-3PO will not deteriorate too much after each update.
 6 Experiments
-In our experiments, we aim to answer the following questions:
-Q1: Does S-3PO achieve zero-violation during the training?
+In our experiments, we aim to answer: Q1: Does S-3PO achieve zero-violation during the training?
 Q2: How does S-3PO without safeguard compare with other state-of-the-art safe RL methods?
@@
 To answer these questions, we conducted our experiments on the safe reinforcement learning benchmark GUARD Zhao et al. (2023a) which is based on Mujoco and Gym interface.
-(a) Point
-(b) Swimmer
-(c) Ant
-(d) Hazard
-(e) Pillar
-Figure 3: Robots and constraints for benchmark problems in our environment.
-Environment Setting
 […]
-Our experiments revolve around four distinct robotic entities that can be categorized into two primary types:
-  1. 1.
-    Wheel Robot: this category encompasses robots that use wheels for locomotion, maintaining a continuous and seamless interaction with their surrounding environment.
-An example is the Point in fig.
-3(a) which is designated as $\mathcal{A}\subseteq\mathbb{R}^{2}$.
-  2. 2.
-    Link Robot: this group includes robots composed of multiple connected links, which interact intermittently with their surroundings through the extremities of these links.
-Furthermore, the shape of these robots can undergo dynamic changes during these interactions.
-Specifically, in our environment, we have 1) Swimmer shown in fig.
-3(b) as a three-link robot ($\mathcal{A}\subseteq\mathbb{R}^{2}$); 2) Ant shown in fig.
-3(c) as a quadrupedal robot ($\mathcal{A}\subseteq\mathbb{R}^{8}$).
-Two different types of constraints are considered.
-  1. 1.
-    Hazard: Dangerous areas shown in fig. 3(d). Hazards are trespassable circles on the ground. The agent is penalized for entering them.
-  2. 2.
-    Pillar: Fixed obstacles shown in  fig. 3(e). The agent is penalized for hitting them.
-    More details about the experiments are discussed in Section F.1.
-(a) Point_1Hazard
-(b) Point_8Hazard
-(c) Point_1Pillar
-(d) Swimmer_1Hazard
-Figure 5: Zero violation performance from four representative test suites in low dimensional systems (evaluated with the safeguard).
-Comparison Group
-The methods in the comparison group include: (i) unconstrained RL algorithm TRPO (Schulman et al., 2015) and TRPO-ISSA.
+Comparison Group The methods in the comparison group include: (i) unconstrained RL algorithm TRPO (Schulman et al., 2015) and TRPO-ISSA.
 (ii) end-to-end constrained safe RL algorithms CPO (Achiam et al., 2017), TRPO-Lagrangian (Bohez et al., 2019), PCPO (Yang et al., 2020), SCPO (Zhao et al., 2023b).
 (iii) We select TRPO as our baseline method since it is state-of-the-art and already has safety-constrained derivatives that can be tested off-the-shelf.
-For all experiments, the policy $\pi$, the value $(V^{\pi},V_{D}^{\pi})$ are all encoded in feedforward neural networks using two hidden layers of size (64,64) with tanh activations.
 The full list of parameters of all methods compared can be found in Section F.2.
-Evaluation Metrics
-For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost and (iii) cost rate.
+Metrics For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost and (iii) cost rate.
 More details are provided in Section F.4.
 We set the limit of cost to 0 for all the safe RL algorithms since we aim to avoid any violation of the constraints.
-(a) Point_1Hazard
-(b) Point_8Hazard
-(c) Point_1Pillar
-(d) Swimmer_1Hazard
-Figure 7: Results from four representative test suites in low dimensional systems (evaluated without the safeguard).
-(a) Point_1Hazard
-(b) Point_8Hazard
-(c) Point_1Pillar
-(d) Swimmer_1Hazard
-Figure 9: Triggering frequency of the safeguard from four representative test suites in low dimensional systems.
 6.2 Evaluating S-3PO and Comparison Analysis
-Zero Violation During Training
-The training performance of four representative test suites are summarized in Figure 5, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Theorem 1.
+Zero Violation During Training The training performance of four representative test suites are summarized in LABEL:fig:zero_violation_performance, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Theorem 5.1.
 For more experiments, please check Appendix F.
 This superior performance is attributed to the safeguard mechanism within the S-3PO framework, which effectively corrects unsafe actions at every step, particularly during training.
-Furthermore, as demonstrated in Figure 7, the reward performance continues to improve throughout the learning process, remaining comparable to state-of-the-art baselines.
-This distinct capability of S-3PO ensures safe reinforcement learning with zero safety violations, even in real-world scenarios, addressing Q1.
-State-wise Safety Without Safety Monitor
-To assess the performance of the S-3PO policy without the safeguard, evaluations are conducted at the end of each epoch.
-During these evaluations, the S-3PO policy is tested over 10,000 steps without the safeguard.
+Furthermore, as demonstrated in LABEL:fig:comparison_results-main, the reward performance remains comparable to state-of-the-art baselines.
+This distinct capability of S-3PO ensures safe reinforcement learning with zero safety violations, addressing Q1.
+State-wise Safety Without Safety Monitor At the end of each epoch, the S-3PO policy is tested over 10,000 steps without the safeguard.
 This allows us to determine whether S-3PO effectively learns a state-wise safe policy through the guidance of the safe set-guided cost.
-The comparative results are presented in Figure 7.
-When contrasted with other baseline safe RL approaches, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance.
-These findings are consistent with Theorem 2 and Theorem 3, highlighting that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
-Learn to Act without Safeguard
-As highlighted in 1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation.
-To illustrate this, we visualize the average number of times the ISSA-based safeguard is triggered per step in Figure 9.
+As shown in in LABEL:fig:comparison_results-main, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance.
+These findings are consistent with Theorem 5.2 and Theorem 5.3, highlighting that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
+Learn to Act without Safeguard As highlighted in 1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation.
+To illustrate this, we visualize the average number of times the ISSA-based safeguard is triggered per step in LABEL:fig:comparison_results_ISSA-main.
 For comparison, TRPO-ISSA is included as a baseline, which relies continuously on the safeguard to maintain safe control.
-Figure 9 shows that S-3PO dramatically reduces the frequency of safeguard activations, approaching zero, indicating that a state-wise safe policy has been effectively learned, thus addressing Q3.
-Ablation on Weighted Loss for Fitting Cost Increment Value Targets
-As pointed in Section 5, fitting $V_{D_{i}}(\hat{s}_{t})$ is a critical step towards solving S-3PO, which is challenging due to non-increasing stair shape of the target sequence.
+LABEL:fig:comparison_results_ISSA-main shows that S-3PO dramatically reduces the frequency of safeguard activations, approaching zero, indicating that a state-wise safe policy has been effectively learned, thus addressing Q3.
+Ablation on Weighted Loss for Fitting Cost Increment Value Targets As pointed in Section 4.3, fitting $V_{D_{i}}(\hat{s}_{t})$ is a critical step towards solving S-3PO, which is challenging due to non-increasing stair shape of the target sequence.
 To elucidate the necessity of weighted loss for solving this challenge, we evaluate the cost rate of S-3PO under six distinct weight settings (0.0, 0.2, 0.4, 0.6, 0.8, 1.0) on Point_4Hazard test suite.
-The results shown in Figure 11(a) validates that a larger weight (hence higher penalty on predictions that violate the characteristics of value targets) results in better cost rate performance.
+The results shown in LABEL:fig:weight_performance validates that a larger weight (hence higher penalty on predictions that violate the characteristics of value targets) results in better cost rate performance.
 This ablation study answers Q4.
-(a) Comparison of cost rate performance with 6 different weights.
-(b) Comparison between “imaginary” cost and action correction cost.
-Figure 11: Comparison results of S-3PO
-Necessity of “Imaginary” Cost
-To understand the importance of the "imaginary" cost within the S-3PO framework, we compare it to another cost based on the magnitude of action correction Chen & Liu (2021).
+Necessity of “Imaginary” Cost To understand the importance of the "imaginary" cost within the S-3PO framework, we compare it to another cost based on the magnitude of action correction Chen and Liu (2021).
 This empirical analysis is conducted using the Point_4Hazard test suite.
-As shown in Figure 11(b), the "imaginary" cost yields superior cost rate performance.
+As shown in LABEL:fig:action_correction_performance, the "imaginary" cost yields superior cost rate performance.
 This suggests that the "imaginary" cost offers deeper insights into the complex dynamics between the robot and its environment, thereby addressing Q5.
-Figure 13: Cost and cost rate performance of Ant_1Hazard.
-Scale S-3PO to High-Dimensional Link Robots
-To showcase S-3PO’s scalability and performance with complex, high-dimensional link robots, we conducted additional tests on Ant_1Hazard featuring 8 dimensional control spaces.
-As shown in Figure 13, S-3PO effectively drives the cost to zero and rapidly reduces the cost rate, showcasing its clear superiority in high-dimensional safety policy learning and highlighting its exceptional scalability to more complex systems.
+Figure 5: Cost performance of Ant_1Hazard.
+Scale S-3PO to High-Dimensional Link Robots To showcase S-3PO’s scalability and performance with complex, high-dimensional link robots, we conducted additional tests on Ant_1Hazard featuring 8 dimensional control spaces.
+As shown in Figure 5, S-3PO effectively drives the cost to zero and rapidly reduces the cost rate, showcasing its superiority in high-dimensional safety policy learning and highlighting its exceptional scalability to more complex systems, which answers Q6.
 7 Conclusion and Future Prospectus
@@
 Subsequently, it integrates a novel “imaginary” safety cost to guide the RL agent towards optimal safe policies.
-S-3PO outperforms existing methods in complex high-dimensional robotics tasks.
-Nevertheless, a noteworthy limitation pertains to the potential costliness of acquiring a physical engine-based simulator (black-box dynamics model).
-A forward-looking perspective entails replacing the black-box dynamics model with a learned surrogate model, factoring in the nuances of errors in learned dynamics.
-This strategic move holds the promise of obliterating the final barrier impeding the seamless integration of safe RL training into real-world applications.
+In the following work, we will try to implement our algorithm in real robot implementation.
 Appendix A Implicit Safe Set Algorithm Details
@@
 Importantly, the synthesis of the safety index for 2D collision evasion remains independent of specific dynamic models, but under the following assumption:
-Assumption 1 (2D Collision Avoidance).
+Assumption 1 (2D Collision Avoidance)
 1) The state space is bounded, and the relative acceleration and angular velocity are bounded and both can achieve zeros, i.e., $w\in[w_{min},w_{max}]$ for $w_{min}\leq 0\leq w_{max}$ and $z\in[z_{min},z_{max}]$ for $z_{min}\leq 0\leq z_{max}$; 2) For all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$; 3) The discrete-time system time step $dt\to 0$; 4) At any given time, there can at most be one obstacle becoming safety critical, such that $\phi-\eta\geq 0$ (Sparse Obstacle Environment).
@@
 A.2 Safety Index Synthesis
-Following the rules in (Liu & Tomizuka, 2014), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$.
+Following the rules in (Liu and Tomizuka, 2014), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$.
 Our goal is to choose these parameters such that $\mathcal{A}_{S}^{D}(s)$ is always nonempty.
@@
      Get search direction $\Delta\theta^{*}\leftarrow\theta^{*}_{k+1}-\theta_{k}$
-     for $j=0,1,2,\dots$ do $\triangleright$ Line search
+     for $j=0,1,2,\dots$ do
         $\theta^{\prime}\leftarrow\theta_{k}+\xi^{j}\Delta\theta^{*}$
-         if  $\mathbb{E}_{\hat{s}\sim\tau}[\mathcal{D}_{KL}(\pi_{\theta^{\prime}}\|\pi_{k})[\hat{s}]]\leq\delta$ and
-$\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{\theta^{\prime}}}_{D}(\hat{s},a)-A^{\pi_{k}}_{D}(\hat{s},a)\right]\leq\mathrm{max}(-c,0)$ and
-$(k\leq k_{\mathrm{safe}}~{}\mathrm{or}~{}\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{\theta^{\prime}}}(\hat{s},a)\right]\geq\mathbb{E}_{\hat{s},a\sim\tau}\left[A^{\pi_{k}}(\hat{s},a)\right])$ then
-              $\theta_{k+1}\leftarrow\theta^{\prime}$triangleright Update policy
+[…]
+           $\theta_{k+1}\leftarrow\theta^{\prime}$Update policy}
            break
@@
 Algorithm 2 Adaptive Momentum Boundary Approximation
-1:procedure AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\vec{v}^{r},S,S_{goal}$)
+1:  procedure: AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\vec{v}^{r},S,S_{goal}$)
 2:     Initialize:
@@
 28:     Return Approximated Boundary Set $P$
-29:end procedure
+29:  end procudure
 Algorithm 3 Implicit Safe Set Algorithm (ISSA)
-1:procedure ISSA($\epsilon,\beta,\Sigma,n,a^{r}$)
-2:     Phase 1: $\triangleright$ Phase 1
+1:  procedure: ISSA($\epsilon,\beta,\Sigma,n,a^{r}$)
+2:     Phase 1: {Phase 1}
 3:     Use AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\emptyset,\textit{UNSAFE},\textit{SAFE}$) to sample a collection $\mathbb{S}$ of safe control on the boundary of $\mathcal{A}_{S}^{D}$.
@@
 10:
-11:     Phase 2: $\triangleright$ Phase 2
+11:     Phase 2: {Phase 2}
 12:     Use grid sampling by iteratively increasing sampling resolution to find an anchor safe control $a^{u}$, s.t. safety status of $a^{u}$ is SAFE.
@@
 19:     end if
-20:end procedure
-Appendix C Proof of Theorem 1
-To prove the Theorem 1 we introduce two important corollaries to show that 1) the set of safe control is always nonempty if we choose a safety index that satisfies the design rule in (12); and 2) the proposed algorithm 3 is guaranteed to find a safe control if there exists one.
-With these two corollaries, it is then straightforward to prove the Theorem 1.
+20:  end procudure
+Appendix C Proof of Theorem 5.1
+To prove the Theorem 5.1 we introduce two important corollaries to show that 1) the set of safe control is always nonempty if we choose a safety index that satisfies the design rule in (11); and 2) the proposed algorithm 3 is guaranteed to find a safe control if there exists one.
+With these two corollaries, it is then straightforward to prove the Theorem 5.1.
 C.1 Feasibility of the Safety Index for Continuous-Time System
-Corollary 1 (Non-emptiness of the set of safe control).
+Corollary C.1 (Non-emptiness of the set of safe control).
 If 1) the dynamic system satisfies the conditions in Assumption 1; and 2) the safety index is designed according to the rule in Appendix A, then the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$.
 Note that the set of safe control $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$ is non-empty if and only if it is non-empty in the following two cases: $\phi(s)-\eta<0$ or $\phi(s)-\eta\geq 0$.
-In the following discussion, we first show that the safety index design rule guarantees a non-empty set of safe control if there’s only one obstacle when $\phi(s)-\eta\geq 0$ (Lemma 1).
-Then we show that the set of safe control is non-empty if there’s only one obstacle when $\phi(s)-\eta<0$ (Lemma 2).
-Finally, we leverage Lemma 1 and Lemma 2 to show $\mathcal{A}_{S}^{D}(s)$ is non-empty if there’re multiple obstacles at any state.
-Lemma 1.
+In the following discussion, we first show that the safety index design rule guarantees a non-empty set of safe control if there’s only one obstacle when $\phi(s)-\eta\geq 0$ (Lemma C.2).
+Then we show that the set of safe control is non-empty if there’s only one obstacle when $\phi(s)-\eta<0$ (Lemma C.4).
+Finally, we leverage Lemma C.2 and Lemma C.4 to show $\mathcal{A}_{S}^{D}(s)$ is non-empty if there’re multiple obstacles at any state.
+Lemma C.2.
 If the dynamic system satisfies the conditions in Assumption 1 and there is only one obstacle in the environment, then the safety index design rule in Appendix A ensures that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for $x$ such that $\phi(s)-\eta\geq 0$.
-Proof.
+Proof C.3.
 For $x$ such that $\phi(s)-\eta\geq 0$, the set of safe control becomes
@@
 Hence (18) holds, which then implies that the inequality in (16) holds.
-The three cases cover all possible situations. Hence (17) always hold and the claim in the lemma is verified. ∎
-Lemma 2.
+The three cases cover all possible situations. Hence (16) always hold and the claim in the lemma is verified.
+Lemma C.4.
 If the dynamic system satisfies the contidions in Assumption 1 and there is only one obstacle in the environment, then the safety index design rule in Appendix A ensures that $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$ for any $s$ that $\phi(s)-\eta<0$.
-Proof.
+Proof C.5.
 For $s$ such that $\phi(s)-\eta<0$, the set of safe control becomes
@@
 Hence $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$ any $s$ that $\phi(s)-\eta<0$ and the claim is true.
-∎
-Proof of Corollary 1.
-Proof.
-If there is one obstacle, then lemma 1 and lemma 2 have proved that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for all $s$.
+Proof of Corollary C.1.
+Proof C.6.
+If there is one obstacle, then lemma C.2 and lemma C.4 have proved that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for all $s$.
 Now we need to consider the case where there are more than one obstacle but the environment is sparse in the sense that at any time step, there is at most one obstacle which is safety critical, i.e.
@@
 Denote ${\mathcal{A}_{S}^{D}}_{j}(s):=\{a\in\mathcal{A}\mid\phi_{j}(f(s,a))\leq\phi_{j}(s)\}$.
-Lemma 1 ensures that ${\mathcal{A}_{S}^{D}}_{j}(s)$ is nonempty.
+Lemma C.2 ensures that ${\mathcal{A}_{S}^{D}}_{j}(s)$ is nonempty.
 Denote ${\mathcal{A}_{S}^{D}}_{k}(s):=\{a\in\mathcal{A}\mid\phi_{k}(f(s,a))\leq 0\}$ where $k\neq j$.
-Since $\phi_{k}(s)<0$, lemma 2 ensures that ${\mathcal{A}_{S}^{D}}_{k}(s)=\mathcal{A}$.
+Since $\phi_{k}(s)<0$, lemma C.4 ensures that ${\mathcal{A}_{S}^{D}}_{k}(s)=\mathcal{A}$.
 Note that the set of safe control can be written as:
@@
 Therefore, we have $\phi_{i}(s)<0$ for all $i$.
-According to Lemma 2, $\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}=\mathcal{A}$.
+According to Lemma C.4, $\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}=\mathcal{A}$.
 Hence the set of safe control satisfies the following relationship
@@
 $=\mathcal{A}\neq\emptyset$ (26c)
-In summary, $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$ and the claim holds. ∎
+In summary, $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$ and the claim holds.
 C.2 Feasibility of ISSA
-Corollary 2 (Feasibility of Algorithm 3).
+Corollary C.7 (Feasibility of Algorithm 3).
 If the set of safe control is non-empty, Algorithm 3 can always find a sub-optimal solution of (5) with a finite number of iterations.
@@
 Note that Phase 2 first finds an anchor safe control $a^{u}$, then use it with AdamBA (Algorithm 2) to find the solution of (5).
-In the following discussion, we first show that the safety index design rule guarantees $a^{u}$ can be found with finite iterations (Lemma 3).
-Then we show that AdamBA will return a solution if it enters the exponential decay phase (Lemma 4).
-Subsequently, we show that the evoked AdamBA procedures in phase 2 will definitely enter exponential decay phase (Lemma 5).
+In the following discussion, we first show that the safety index design rule guarantees $a^{u}$ can be found with finite iterations (Lemma C.8).
+Then we show that AdamBA will return a solution if it enters the exponential decay phase (Lemma C.10).
+Subsequently, we show that the evoked AdamBA procedures in phase 2 will definitely enter exponential decay phase (Lemma C.12).
 Finally, we provide a upper bound of the computation iterations for Algorithm 3.
-Lemma 3 (Existence).
+Lemma C.8 (Existence).
 If the synthesized safety index can guarantee a non-empty set of safe control, then we can find an anchor point in phase 2 of Algorithm 3 with finite iterations (line 11 in algorithm 3).
-Proof.
+Proof C.9.
 Figure 7: Illustration of the grid sampling to find an anchor control point.
@@
 The grid sampling to find an anchor control point is illustrated in Figure 7.
-Lemma 4 (Convergence).
+Lemma C.10 (Convergence).
 If AdamBA enters the exponential decay phase, then it can always return a boundary point approximation (with desired safety status) where the approximation error is upper bounded by $\epsilon$.
-Proof.
+Proof C.11.
 According to Algorithm 2, exponential decay phase applies Bisection method to locate the boundary point $P_{b}$ until $\|P_{NS}-P_{S}\|<\epsilon$.
@@
 $<\epsilon.$
-Lemma 5 (Feasibility).
+Lemma C.12 (Feasibility).
 If we enter the phase 2 of Algorithm 3 with an anchor safe control being sampled, we can always find a local optimal solution of (5).
-Proof.
-According to line 13-15, after an anchor safe control is being sampled, phase 2 of Algorithm 3 will evoke at most two AdamBA processes.
-Hence, Lemma 5 can be proved by showing one of the two AdamBA will return a local optimal solution for (6).
-Next we show Lemma 5 holds in two cases.
+Proof C.13.
+According to line 13-16, after an anchor safe control is being sampled, phase 2 of Algorithm 3 will evoke at most two AdamBA processes.
+Hence, Lemma C.12 can be proved by showing one of the two AdamBA will return a local optimal solution for (5).
+Next we show Lemma C.12 holds in two cases.
 Case 1:
@@
 According to Algorithm 2, a solution will be returned only if AdamBA enters exponential decay stage.
-Hence, according to Lemma 4, $a^{u}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
+Hence, according to Lemma C.10, $a^{u}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
 Figure 8: Illustration of the case when it is unable to find $u^{*}$.
@@
 line 13 of Algorithm 3 fails to find a solution.
-In this case, the second AdamBA process is evoked (line 15 of Algorithm 3).
+In this case, the second AdamBA process is evoked (line 16 of Algorithm 3).
 Since no solution is returned from the first AdamBA process (13 of Algorithm 3), where we start from $a^{r}$ and exponentially outreach along the direction $\vec{v}_{a}=\frac{a^{u}-a^{r}}{||a^{u}-a^{r}||}$, then all the searched control point along $\vec{v}_{a}$ is UNSAFE.
@@
 where the safety statuses of both $a_{s}^{1}$ and $a_{s}^{2}$ are UNSAFE.
-During the second AdamBA process (line 15 of Algorithm 3), we start from $a^{u}$ and exponentially outreach along the direction $\vec{v}_{r}=\frac{a^{r}-a^{u}}{||a^{r}-a^{u}||}$ with initial step size $\beta=\frac{\|a^{r}-a^{u}\|}{4}$.
+During the second AdamBA process (line 16 of Algorithm 3), we start from $a^{u}$ and exponentially outreach along the direction $\vec{v}_{r}=\frac{a^{r}-a^{u}}{||a^{r}-a^{u}||}$ with initial step size $\beta=\frac{\|a^{r}-a^{u}\|}{4}$.
 Hence, denote $\bar{a}_{s}^{1}$ as the first sample point along $\vec{v}_{r}$, and $\bar{a}_{s}^{1}$ satisfies
@@
 which indicates $\bar{a}_{s}^{1}=a_{s}^{2}$, and the safety status of $\bar{a}_{s}^{1}$ is UNSAFE.
-Since the safe status of $a^{u}$ is SAFE and a UNSAFE point can be sampled during the exponential outreach stage, the second AdamBA process (line 15 of Algorithm 3) will enter exponential decay stage.
-Therefore, according to Lemma 4, ${a^{u}}^{*}$ will always be returned and ${a^{u}}^{*}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
+Since the safe status of $a^{u}$ is SAFE and a UNSAFE point can be sampled during the exponential outreach stage, the second AdamBA process (line 16 of Algorithm 3) will enter exponential decay stage.
+Therefore, according to Lemma C.10, ${a^{u}}^{*}$ will always be returned and ${a^{u}}^{*}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
 The two cases cover all possible situations. Hence, after an anchor safe control $a^{u}$ is sampled, phase 2 of Algorithm 3 can always find a local optima of (5).
-Proof of Corollary 2.
-Proof.
-According to Lemma 3 and Lemma 5, Algorithm 3 is able to find local optimal solution of (6). Next, we will prove Algorithm 3 can be finished within finite iterations.
+Proof of Corollary C.7.
+Proof C.14.
+According to Lemma C.8 and Lemma C.12, Algorithm 3 is able to find local optimal solution of (5). Next, we will prove Algorithm 3 can be finished within finite iterations.
 According to Algorithm 3, ISSA include (i) one procedure to find anchor safe control $a_{a}$, and (ii) at most three AdamBA procedures.
-Firstly, based on Lemma 3, $a_{a}$ can be found within finite iterations.
+Firstly, based on Lemma C.8, $a_{a}$ can be found within finite iterations.
 Secondly, each AdamBA procedure can be finished within finite iterations due to:
@@
 Therefore, ISSA can be finished within finite iterations.
-C.3 Proof of Theorem 1
-Before we prove Theorem 1, we first define set $\mathcal{S}_{S}^{D}:=\{s|\phi(s)\leq 0\}$, and we start with a preliminary result regarding $\mathcal{S}_{S}^{D}$ that is useful for proving the main theorem:
-Lemma 6 (Forward Invariance of $\mathcal{S}_{S}^{D}$).
+C.3 Proof of Theorem 5.1
+Before we prove Theorem 5.1, we first define set $\mathcal{S}_{S}^{D}:=\{s|\phi(s)\leq 0\}$, and we start with a preliminary result regarding $\mathcal{S}_{S}^{D}$ that is useful for proving the main theorem:
+Lemma C.15 (Forward Invariance of $\mathcal{S}_{S}^{D}$).
 If the control system satisfies Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$.
-Proof.
-If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, then we can ensure the system has nonempty set of safe control at any state by Corollary 1.
-By Corollary 2, the implicit safe set algorithm can always find local optima solution of (6).
+Proof C.16.
+If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, then we can ensure the system has nonempty set of safe control at any state by Corollary C.1.
+By Corollary C.7, the implicit safe set algorithm can always find local optima solution of (5).
 The local optima solution always satisfies the constraint $\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}$, which indicates that 1) if $\phi(s_{t_{0}})\leq 0$, then $\phi(s_{t})\leq 0,\forall t\geq t_{0}$.
 Note that $\phi(s)\leq 0$ demonstrates that $s\in\mathcal{S}_{S}^{D}$.
-∎
-Proof the Theorem 1
-Proof.
-Leveraging Lemma 6, we then proceed to prove that the forward invariance to the set $\mathcal{S}_{S}^{D}$ guarantees the forward invariance to the set $\mathcal{\bar{S}}\subseteq\mathcal{S}_{S}$ which $\mathcal{\bar{S}}=\mathcal{S}_{S}\cap\mathcal{S}_{S}^{D}$.
+Proof the Theorem 5.1
+Proof C.17.
+Leveraging Lemma C.15, we then proceed to prove that the forward invariance to the set $\mathcal{S}_{S}^{D}$ guarantees the forward invariance to the set $\mathcal{\bar{S}}\subseteq\mathcal{S}_{S}$ which $\mathcal{\bar{S}}=\mathcal{S}_{S}\cap\mathcal{S}_{S}^{D}$.
 Depending on the relationship between $\mathcal{S}_{S}^{D}$ and $\mathcal{S}_{S}$, there are two cases in the proof which we will discuss below.
@@
 In this case, $\mathcal{\bar{S}}=\mathcal{S}_{S}^{D}$.
-According to Lemma 6, If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$ and hence $\mathcal{\bar{S}}$.
+According to Lemma C.15, If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$ and hence $\mathcal{\bar{S}}$.
 Case 2: $\mathcal{S}_{S}^{D}=\{s|\phi(s)\leq 0\}$ is NOT a subset of $\mathcal{S}_{S}=\{s|\phi_{0}(s)\leq 0\}$.
@@
 Therefore, we have that if $s_{t}\in\mathcal{\bar{S}}$ then $s_{t+1}\in\mathcal{S}_{S}$.
-Thus, we also have $s_{t+1}\in\mathcal{\bar{S}}$ by Lemma 6.
+Thus, we also have $s_{t+1}\in\mathcal{\bar{S}}$ by Lemma C.15.
 By induction, we have if $s_{t_{0}}\in\mathcal{\bar{S}}$, $s_{t}\in\mathcal{\bar{S}},\forall t>t_{0}$.
@@
 Thus, if the initial state is safe, then the following state will always stay in $\mathcal{\bar{S}}$ which means ISSA can ensure the safety during training.
-Appendix D Proof of Theorem 2
+Appendix D Proof of Theorem 5.2
 Mathematically, S-3PO requires the policy update at each iteration is bounded within a trust region, and updates policy via solving following optimization:
@@
 This formulation helps us easily obtain the following lemma.
-Lemma 7.
+Lemma D.1.
 For any function $f:\hat{\mathcal{S}}\mapsto\mathbb{R}$ and any policy $\pi$,
 $\underset{\hat{s}\sim\hat{\mu}}{\operatorname{E}}[f(\hat{s})]+\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\gamma f(\hat{s}^{\prime})]-\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}[f(\hat{s})]=0.$ (36)
-Proof.
-Multiply both sides of (36) by $(I-\gamma P_{\pi})$ and take the inner product with the vector $f\in\mathbb{R}^{|\hat{\mathcal{S}}|}$. ∎
-Combining Lemma 7 with (35), we obtain the following, for any function $f$ and any policy $\pi$:
+Proof D.2.
+Multiply both sides of (35) by $(I-\gamma P_{\pi})$ and take the inner product with the vector $f\in\mathbb{R}^{|\hat{\mathcal{S}}|}$.
+Combining Lemma D.1 with (34), we obtain the following, for any function $f$ and any policy $\pi$:
 $\mathcal{J}_{g}(\pi)=\underset{\hat{s}\sim\hat{\mu}}{\operatorname{E}}\left[f(\hat{s})\right]+\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}\left[g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})\right]$ (37)
-Lemma 8.
+Lemma D.3.
 For any function $f\mapsto\mathbb{R}$ and any policies $\pi^{\prime}$ and $\pi$, define
@@
 where $D_{TV}$ is the total variational divergence. Furthermore, the bounds are tight(when $\pi^{\prime}$ = $\pi$, the LHS and RHS are identically zero).
-Proof.
+Proof D.4.
 First, for notational convenience, let $\delta_{f}(\hat{s},a,\hat{s}^{\prime})\doteq g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})$. By (37), we obtain the identity
@@
 The lower bound leads to (39), and the upper bound leads to (40).
-∎
-Lemma 9.
+Lemma D.5.
 The divergence between discounted future state visitation distributions, $||\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}||_{1}$, is bounded by an average divergence of the policies $\pi^{\prime}$ and $\pi$:
@@
 where $D_{TV}(\pi^{\prime}||\pi)[\hat{s}]=\dfrac{1}{2}\sum_{a}|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|.$
-Proof.
+Proof D.6.
 Firstly, we introduce an identity for the vector difference of the discounted future state visitation distributions on two different policies, $\pi^{\prime}$ and $\pi$.
@@
 The new policy improvement bound follows immediately.
-Lemma 10.
+Lemma D.7.
 For any function $f:\hat{\mathcal{S}}\mapsto\mathbb{R}$ and any policies $\pi^{\prime}$ and $\pi$, define $\delta_{f}(\hat{s},a,\hat{s}^{\prime})\doteq g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})$,
@@
 Furthermore, the bounds are tight (when $\pi^{\prime}$ = $\pi$, all three expressions are identically zero)
-Proof.
-Begin with the bounds from lemma 8 and bound the divergence $D_{TV}(\dot{d}^{\pi^{\prime}}||\dot{d}^{\pi})$ by lemma 9. ∎
+Proof D.8.
+Begin with the bounds from lemma D.3 and bound the divergence $D_{TV}(\dot{d}^{\pi^{\prime}}||\dot{d}^{\pi})$ by lemma D.5.
 D.2 Safety invariance in expectation between S-3PO policies
-Corollary 3 (SCPO Update Constraint Satisfaction).
+Corollary D.9 (SCPO Update Constraint Satisfaction).
 Suppose $\mathcal{J}_{D}(\pi_{k})\leq 0$ and $\pi_{k},\pi_{k+1}$ are related by (32), then $D$-return for $\pi_{k+1}$ satisfies
 $\forall i,\mathcal{J}_{D}(\pi_{k+1})\leq 0.$
-Proof.
-The choice of $f=\hat{V}_{D}^{\pi},g=D$ in lemma 10 leads to following inequality:
+Proof D.10.
+The choice of $f=\hat{V}_{D}^{\pi},g=D$ in lemma D.7 leads to following inequality:
 $\mathcal{\hat{J}}_{D}(\pi^{\prime})-\mathcal{\hat{J}}_{D}(\pi)\leq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[\hat{A}_{D}^{\pi}(\hat{s},a)+2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right].$ (51)
@@
 And $\hat{V}_{D}^{\pi},\hat{A}_{D}^{\pi}$ are also the discounted version of $V_{D}^{\pi}$ and $A_{D}^{\pi}$.
-Note that according to Lemma 10 one can only get this the inequality holds when $\gamma\in(0,1)$.
+Note that according to Lemma D.7 one can only get this the inequality holds when $\gamma\in(0,1)$.
 Then we can define $\mathcal{F}(\gamma)=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[\hat{A}_{D}^{\pi}(\hat{s},a)+2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right]-\mathcal{\hat{J}}_{D}(\pi^{\prime})+\mathcal{\hat{J}}_{D}(\pi)$ with the following condition holds:
@@
 $\mathcal{J}_{D}(\pi^{\prime})-\mathcal{J}_{D}(\pi)\leq\underset{\begin{subarray}{c}\hat{s}\sim\bar{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[A_{D}^{\pi}(\hat{s},a)+2(H+1)\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right].$
-where $\bar{d}^{\pi}=\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|\pi)$. Thus, following the inequality (33), the Corollary 3 is proofed.
-D.3 Proof of Theorem 2
+where $\bar{d}^{\pi}=\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|\pi)$. Thus, following the inequality (32), the Corollary D.9 is proofed.
+D.3 Proof of Theorem 5.2
 If $s_{t}\in\mathcal{\bar{S}}\in\mathcal{S}_{S}^{D}$, then $\mathcal{J}_{D}(\pi_{t})\leq 0$ and $\phi(s_{t})\leq 0$.
-According to Corollary 3, we know that if policy $\pi_{t}$ is updated by solving (33), then we have $\mathcal{J}_{D}(\pi_{t+1})\leq 0$ which means:
+According to Corollary D.9, we know that if policy $\pi_{t}$ is updated by solving (32), then we have $\mathcal{J}_{D}(\pi_{t+1})\leq 0$ which means:
 $\underset{\tau\sim\pi_{t+1}}{\mathbb{E}}\left[\phi(s_{t+1})\right]=\mathcal{J}_{D}(\pi_{t+1})+\phi(s_{t})\leq 0$ (53)
 Thus $s_{t+1}\in\mathcal{S}_{S}^{D}$ in expectation. According to the proof of Section C.3, we know that $s_{t+1}\in\mathcal{\bar{S}}\in\mathcal{S}_{S}^{D}$ in expectation.
-Appendix E Proof of Theorem 3
+Appendix E Proof of Theorem 5.3
 E.1 KL Divergence Relationship Between $d^{\pi_{k}}$ and $\bar{d}^{\pi_{k}}$
-Lemma 11.
+Lemma E.1.
 $\underset{\hat{s}\sim d_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]<\underset{\hat{s}\sim\bar{d}_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]$
-Proof.
+Proof E.2.
 $\underset{\hat{s}\sim d_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]$ $=\sum_{\hat{s}}(1-\gamma)\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi)\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$
@@
 E.2 Trust Region Update Performance
-Lemma 12.
+Lemma E.3.
 For any policies $\pi^{\prime},\pi$, with $\epsilon^{\pi^{\prime}}\doteq max_{\hat{s}}|E_{a\sim\pi^{\prime}}[A^{\pi}(\hat{s},a)]|$, and define $d^{\pi}=(1-\gamma)\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi)$ as the discounted augmented state distribution using $\pi$, then the following bound holds:
 $\mathcal{J}(\pi^{\prime})-\mathcal{J}(\pi)>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim{\pi^{\prime}}\end{subarray}}{\mathbb{E}}\bigg{[}A^{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}\mathcal{D}_{TV}({\pi^{\prime}}\|\pi)[\hat{s}]\bigg{]}$ (54)
-Proof.
-The choice of $f=V_{\pi},g=R$ in lemma 10 leads to following inequality:
+Proof E.4.
+The choice of $f=V_{\pi},g=R$ in lemma D.7 leads to following inequality:
 For any policies $\pi^{\prime},\pi$, with $\epsilon^{\pi^{\prime}}\doteq max_{\hat{s}}|E_{a\sim\pi^{\prime}}[A_{\pi}(\hat{s},a)]|$, the following bound holds:
@@
 $>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathrm{E}}\left[A_{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}D_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right]$
-At this point, the lemma 12 is proved.
-E.3 Proof of Theorem 3
-Proof.
+At this point, the lemma E.3 is proved.
+E.3 Proof of Theorem 5.3
+Proof E.5.
 If ISSA will constantly be triggered, the safeguard should be treated as a component of the environment, which indicates the environment is non-stationary for policy $\pi$ after safeguard is disabled.
@@
 Therefore, we can infer S-3PO worst performance degradation after $k_{safe}$ updates following the trust region results of finite-horizon MDPs.
-Utilizing Lemma 12 and the relationship between the total variation divergence and the KL divergence, we have:
+Utilizing Lemma E.3 and the relationship between the total variation divergence and the KL divergence, we have:
 $\mathcal{J}(\pi^{\prime})-\mathcal{J}(\pi)>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\bigg{[}A^{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}\sqrt{\frac{1}{2}\mathbb{E}_{\hat{s}\sim d^{\pi}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]}\bigg{]}.$ (55)
@@
 $~{}~{}~{}~{}~{}\mathop{\mathbf{s.t.}}~{}~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta.$
-Due to Lemma 11, if two policies are related with Equation 57, they are related with the following optimization:
+Due to Lemma E.1, if two policies are related with Equation 56, they are related with the following optimization:
 $\pi_{k+1}=\mathop{\underset{\pi\in\Pi_{\theta}}{\mathop{\mathbf{argmax}}}}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi_{k}}\\ a\sim\pi\end{subarray}}{\mathbb{E}}[A^{\pi_{k}}(\hat{s},a)]$ (57)
@@
 Additional high dimensional link robot
-To scale our method to high dimensional link robots. We additionally adopt Walker shown in 17 as a bipedal robot ($\mathcal{A}\subseteq\mathbb{R}^{10}$) in our experiments.
+To scale our method to high dimensional link robots. We additionally adopt Walker shown in 13 as a bipedal robot ($\mathcal{A}\subseteq\mathbb{R}^{10}$) in our experiments.
 Figure 13: Walker
@@
 Cost reduction - - - 0.0 - 0.0 0.0
-F.3 Metrics Comparison
-In Tables 5(c), 6(c) and 7, we report all the $9$ results of our test suites by three metrics:
+F.3 Practical Implementation Details
+In this section, we summarize implementation techniques that helps with S-3PO’s practical performance. The pseudocode of S-3PO is give as algorithm 1.
+Weighted loss for cost value targets
+A critical step in S-3PO requires fitting of the cost increment value functions, denoted as $V_{D}^{\pi}(\hat{s}_{t})$.
+By definition, $V_{D}^{\pi}(\hat{s}_{t})$ is equal to the maximum cost increment in any future state over the maximal state-wise cost so far.
+In other words, $V_{D}^{\pi}(\hat{s}_{t})$ forms a non-increasing stair shape along the trajectory.
+Here we visualize an example of $V_{D}^{\pi}(\hat{s}_{t})$ in Figure 14.
+To enhance the accuracy of fitting this stair shape function, a weighted loss strategy is adopted, capitalizing on its monotonic property.
+Specifically, we define a weighted loss $L_{weight}$:
+$\begin{split}L_{weight}&=L(\hat{y}_{t}-y_{t})*(1+w*\text{1}[(\hat{y}_{t}-y_{t-1})>0])\end{split}$
+Figure 14: $V_{[D_{i}]\pi}(\hat{s})$ target of five sampled episodes.
+where $L$ denotes Mean Squared Error (MSE), $\hat{y}_{t}$ is the prediction, ${y}_{t}$ is the fitting target and ${w}$ is the penalty weight.
+To account for the initial step (${t=0}$), we set ${y_{t-1}}$ to sufficiently large, thereby disregarding the weighted term associated with the first step.
+In essence, the rationale is to penalize any prediction that violates the non-increasing characteristics of the target sequence, thereby leading to an improved fitting quality.
+Line search scheduling
+Note that in (10), there are two constraints: (a) the trust region and (b) the bound on expected advantages.
+In practice, due to approximation errors, constraints in (10) might become infeasible.
+In that case, we perform a recovery update that only enforces the cost advantage $A_{D}^{\pi}$ to decrease starting from early training steps (in first $k_{\mathrm{safe}}$ updates), and starts to enforce reward improvements of $A^{\pi}$ towards the end of training.
+This is different from Zhao et al.
+(2023b), where the reward improvements are enforced at all times.
+This is because SCPO only guarantees safety (constraint satisfaction) after convergence, while S-3PO prioritizes constraining imaginary safety violation.
+With our line search scheduling, S-3PO is able to first grasp a safe policy, and then improve the reward performance.
+In that way, S-3PO achieves zero safety violation both during training and in testing with a worst-case performance degradation guarantee.
+F.4 Metrics Comparison
+In Tables 5, 6 and 7, we report all the $9$ results of our test suites by three metrics:
   • The average episode return $J_{r}$.
@@
 Each metric was averaged over two random seed.
-The evaluation performance curves of all experiments are shown in Figures 19, 25 and 31, the training performance curves of all experiments are shown in Figures 21, 27 and 33 and the ISSA performance curves of all experiments are shown in Figures 23, 29 and 35
+The evaluation performance curves of all experiments are shown in LABEL:fig:exp-point-hazard-eval, LABEL:fig:exp-point-pillar-eval and LABEL:fig:exp-link-robot-eval, the training performance curves of all experiments are shown in LABEL:fig:exp-point-hazard-train, LABEL:fig:exp-point-pillar-train and LABEL:fig:exp-link-robot-train and the ISSA performance curves of all experiments are shown in LABEL:fig:exp-point-hazard-ISSA, LABEL:fig:exp-point-pillar-ISSA and LABEL:fig:exp-link-robot-ISSA
 Table 5: Metrics of three Point_Hazard environments obtained from the final epoch.
-(a) Point_1Hazard
+Point_1Hazard
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 S-3PO 2.4157 0.0000 0.0000 2.2878 0.0000 0.0000 0.0285
-(b) Point_4Hazard
+Point_4Hazard
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 S-3PO 2.3868 0.0000 0.0000 2.3550 0.0000 0.0000 0.0117
-(c) Point_8Hazard
+Point_8Hazard
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 Table 6: Metrics of three Pillar_Hazard environments obtained from the final epoch.
-(a) Point_1Pillar
+Point_1Pillar
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 S-3PO 2.5551 0.0000 0.0000 2.5241 0.0000 0.0000 0.0060
-(b) Point_4Pillar
+Point_4Pillar
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 S-3PO 2.2513 0.0114 0.0000 2.3459 0.0000 0.0000 0.0116
-(c) Point_8Pillar
+Point_8Pillar
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 Table 7: Metrics of three link robots environments obtained from the final epoch.
-(a) Swimmer_1Hazard
+Swimmer_1Hazard
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 S-3PO 2.6032 0.0313 0.0000 2.6239 0.0001 0.0000 0.0378
-(b) Ant_1Hazard
+Ant_1Hazard
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 S-3PO 2.2047 0.0000 0.0002 2.2031 0.0000 0.0002 0.0001
-(c) Walker_1Hazard
+Walker_1Hazard
 Algorithm Evaluation Performance Training Performance ISSA Performance
@@
 S-3PO 2.6117 0.3437 0.0025 2.6055 0.2665 0.0025 0.0319
-(a) Point_1Hazard
-(b) Point_4Hazard
-(c) Point_8Hazard
-Figure 19: Evaluation performance of Point_Hazard
-(a) Point_1Hazard
-(b) Point_4Hazard
-(c) Point_8Hazard
-Figure 21: Training performance of Point_Hazard
-(a) Point_1Hazard
-(b) Point_4Hazard
-(c) Point_8Hazard
-Figure 23: ISSA performance of Point_Hazard
-(a) Point_1Pillar
-(b) Point_4Pillar
-(c) Point_8Pillar
-Figure 25: Evaluation performance of Point_Pillar
-(a) Point_1Pillar
-(b) Point_4Pillar
-(c) Point_8Pillar
-Figure 27: Training performance of Point_Pillar
-(a) Point_1Pillar
-(b) Point_4Pillar
-(c) Point_8Pillar
-Figure 29: ISSA performance of Point_Pillar
-(a) Swimmer_1Hazard
-(b) Ant_1Hazard
-(c) Walker_1Hazard
-Figure 31: Evaluation performance of link robots
-(a) Swimmer_1Hazard
-(b) Ant_1Hazard
-(c) Walker_1Hazard
-Figure 33: Training performance of link robots
-(a) Swimmer_1Hazard
-(b) Ant_1Hazard
-(c) Walker_1Hazard
-Figure 35: ISSA performance of link robots

### v5

@@
 Abstract
-Deep reinforcement learning (RL) excels in various control tasks, yet the absence of safety guarantees hampers its real-world applicability.
-In particular, explorations during learning usually results in safety violations, while the RL agent learns from those mistakes.
-On the other hand, safe control techniques ensure persistent safety satisfaction but demand strong priors on system dynamics, which is usually hard to obtain in practice.
-To address these problems, we present Safe Set Guided State-wise Constrained Policy Optimization (S-3PO), a pioneering algorithm generating state-wise safe optimal policies with zero training violations, i.e., learning without mistakes.
-S-3PO first employs a safety-oriented monitor with black-box dynamics to ensure safe exploration.
-It then enforces an "imaginary" cost for the RL agent to converge to optimal behaviors within safety constraints.
-S-3PO outperforms existing methods in high-dimensional robotics tasks, managing state-wise constraints with zero training violation.
-This innovation marks a significant stride towards real-world safe RL deployment.
+Deep reinforcement learning (RL) has achieved remarkable success across various control tasks.
+However, its reliance on exploration through trial and error often results in safety violations during training.
+To mitigate this, safety filters are commonly employed to correct unsafe actions generated by the RL policy.
+Yet, a key challenge remains: how to enable safety-filter-guided learning to produce a policy that remains optimally safe even after the filter is removed.
+In this paper, we propose Safe Set Guided State-wise Constrained Policy Optimization (S-3PO) — a novel algorithm designed to generate optimal safe RL policies with zero training violations while maintaining safety during evaluation even without any safety filter.
+S-3PO integrates a safety-oriented monitor operating on black-box dynamics to ensure safe exploration and introduces an imaginary cost mechanism that guides the safe RL agent toward optimal behavior under safety constraints.
+This imaginary cost inherits the interpretability of the safety filter while outperforming conventional imitation-based cost designs.
+Equipped with state-of-the-art components, S-3PO demonstrates superior performance on high-dimensional robotic control tasks, effectively handling expected state-wise constraints and ensuring safety throughout the training process.
 1 Introduction
-Reinforcement Learning (RL) has showcased remarkable advancements in domains like control and games.
-However, its focus on reward maximization sometimes neglects safety, potentially leading to catastrophic outcomes (Gu et al., 2022).
-To rectify this, the concept of safe RL emerged, aiming to ensure safety throughout or after training.
-Initial endeavors centered on Constrained Markov Decision Processes, often emphasizing cumulative or chance constraints (Ray et al., 2019; Achiam et al., 2017; Liu et al., 2021).
-While effective, these approaches lack instantaneous safety guarantees, which is crucial for managing emergencies like collision avoidance in autonomous vehicles (Zhao et al., 2023c; He et al., 2023a).
-Recent strides Zhao et al.
-(2023b) leveraged the Maximum Markov Decision Process to ensure instantaneous safety by bounding violations while integrating trust region techniques for policy enhancement, resulting in simultaneous improvement of worst-case performance and adherence to cost constraints.
-However, these approaches still cannot ensure safety during learning due to potentially unsafe explorative behaviors.
-Meanwhile, safe control methods to continuously meet stringent safety requirements in predictable environments are widely examined, with energy function-based methods being the most prevalent Khatib (1986); Ames et al.
-(2014); Liu and Tomizuka (2014); Gracia et al.
-(2013); Wei and Liu (2019).
-These methods establish energy functions assigning lower energy to safe states and project nominal control into energy dissipating control, hence persistently maintaining system safety.
-However, these approahces heavily rely on accessible white-box analytical models of system dynamics.
-The Implicit Safe Set Algorithm (ISSA) Zhao et al.
-(2021, 2024) addresses this problem using black-box optimization over black-box dynamics models.
-Nevertheless, the reliance on continuous monitoring to ensure safety imposes significant computational overhead, and the RL policy itself does not inherently guarantee safe behavior.
-This combination presents a major obstacle to the successful deployment of the algorithm in real-time, safety-critical applications.
+Safe Reinforcement Learning (safe RL) has emerged as a powerful approach in domains such as games and robotic control, where ensuring safety during or after training is critical.
+While objective-based methods aim to optimize reward, they often lack formal guarantees on safety performance [Bohez et al.
+(2019)].
+To address this, many approaches enforce hard constraints [Bouvier et al.
+(2024b), Bouvier et al.
+(2024a)]; however, these methods are typically effective only in low-dimensional systems.
+More recent advances [Zhao et al.
+(2023b), Zhao et al.
+(2024b)] leverage trust-region techniques combined with Maximum Markov Decision Processes (MMDP) to enforce simultaneous improvement of worst-case performance and adherence to cost constraints.
+Despite these developments, RL-based methods fundamentally depend on trial-and-error exploration, making it difficult to guarantee safety throughout the training process.
+A common strategy to mitigate this issue is the use of safety filters [Alshiekh et al.
+(2018)], which is designed to correct unsafe actions generated by the RL policy.
+These safety filters are often constructed using principles from safe control theories [Shao et al.
+(2021)], where energy function-based methods remain the most widely adopted approach [Khatib (1986), Ames et al.
+(2014), Liu and Tomizuka (2014), Gracia et al.
+(2013), Wei and Liu (2019)].
+Although safety filters can ensure safety during training by overriding unsafe actions, they also prevent the policy from learning how to avoid unsafe behaviors on its own.
+This creates a fundamental dilemma: how can an agent learn to avoid unsafe scenarios if it is always shielded from experiencing them?
+To enable the policy to learn from the safety filter and generate safe actions by itself, rather than only being protected by the safety filter,  [Cheng et al.
+(2019)] use Gaussian Process (GP) models to estimate unknown system dynamics and construct a safety filter that implicitly guides the policy updates based on the history of safety filter interventions.
+To explicitly imitate the effect of the safety filter at each step, other works have considered using reward penalties to learn from safety filters.
+[Krasowski et al.
+(2022)] introduce a constant penalty when the safety filter is activated.
+Yet, this approach does not account for how unsafe the proposed action was.
+In contrast,  [Wabersich and Zeilinger (2021)] penalize the reward with magnitude of the action correction applied by the safety filter.
+However, the magnitude of the correction alone may not accurately capture the true impact of the action on system safety, and the reward penalty can only impose a soft constraint during the learning process.
 Figure 1: Overview of the principles of the S-3PO algorithm.
-In summary, a pressing need arises for a method leveraging both the zero-training-time-violation ability of safe control and optimal convergence of safe RL.
-To bridge this gap, we introduce a novel Safe Set Guided State-wise Constrained Policy Optimization (S-3PO) algorithm.
-S-3PO safeguards the exploration of immature policies with a black-box safe control method and derives a novel formulation where RL learns an optimal safe policy by constraining the state-wise “imaginary" safety violations.
-An overview of the prescribed S-3PO principles is presented in Figure 1.
-Empirical validation underscores S-3PO’s efficacy in training neural network policies encompassing thousands of parameters for high-dimensional simulated robot locomotion tasks.
-Our contribution marks a substantial advancement in the realm of practical safe RL algorithms, poised to find applications in a multitude of real-world challenges.
+These limitations highlight a critical gap: to enable a policy to maintain safety after the safety filter is removed, it is essential to introduce a cost term that explicitly measures how the policy’s actions influence the system’s safety level if the safe filter is removed.
+To this end, we introduce Safe Set Guided State-wise Constrained Policy Optimization (S-3PO).
+S-3PO safeguards the exploration of immature policies through a black-box safe control mechanism and formulates a novel constrained optimization framework where RL learns an optimal safe policy by constraining imaginary safety violations—violations that would have occurred without the filter.
 2 Problem Formulation
-2.1 Preliminaries
+2.1 Assumptions
 Dynamics
@@
 where $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is a deterministic function that maps the current robot state and control to the robot’s state in the next time step.
-To maintain simplicity, our approach focuses on deterministic dynamics, although it is worth noting that the proposed method can be readily extended to accommodate stochastic dynamics Zhao et al.
-(2021); Noren et al.
-(2021).
-Additionally, we assume the access to the dynamics model $f$ is only in the training phase, and is restricted to an implicit black-box form, as exemplified by an implicit digital twin simulator or a deep neural network model Zhao et al.
-(2021).
-We also assume there is no model mismatch, while model mismatch can be addressed by robust safe control Wei et al.
-(2022) and is left for future work.
-Post training, the knowledge of the dynamics model is concealed, aligning with practical scenarios where digital twins of real-world environments are too costly to access during model deployment.
+To maintain simplicity, our approach focuses on deterministic dynamics, although the proposed method can be easily extended to stochastic dynamics [Zhao et al.
+(2021)].
+Additionally, we assume the access to the dynamics model $f$ is only in the training phase and restricted to an black-box form, such as an implicit digital twin simulator or a deep neural network model [Zhao et al.
+(2021)].
+We also assume there is no model mismatch.
+Model mismatch can be addressed by robust safe control [Wei et al.
+(2022)] and is left for future work.
+Post training, the knowledge of the dynamics model is concealed—a benefit of using "imaginary cost"—aligning with practical scenarios where digital twins of real-world environments are too costly to access during deployment.
+Based on these, our core target is to figure out how can safety-filter-guided learning be used to produce a policy that remains optimally safe even without the filter.
 Markov Decision Process
@@
 An MDP is defined by a tuple $(\mathcal{S},\mathcal{A},\gamma,R,P,\mu)$.
-The reward function is denoted by $R:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$, the discount factor by $10\leq\gamma<$ 1, the initial state distribution by $\mu:\mathcal{S}\mapsto\mathbb{R}$, and the transition probability function by $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$.
+The reward function is denoted by $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$, the discount factor by $10\leq\gamma<$ 1, the initial state distribution by $\mu:\mathcal{S}\rightarrow\mathbb{R}$, and the transition probability function by $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$.
 The transition probability $P(s^{\prime}|s,a)$ represents the likelihood of transitioning to state $s^{\prime}$ when the previous state was $s$, and the agent executed action $a$ at state $s$.
@@
 The safety specification requires that the system state remains within a closed subset in the state space, denoted as the “safe set” $\mathcal{S}_{S}$.
-This safe set is defined by the zero-sublevel set of a continuous and piecewise smooth function $\phi_{0}:\mathbb{R}^{n_{s}}\rightarrow\mathbb{R}$, where $\mathcal{S}_{S}=\{s\mid\phi_{0}(s)\leq 0\}$.
-Users directly specify both $\mathcal{S}_{S}$ and $\phi_{0}$, which is easy to specify.
-For instance, for collision avoidance, $\phi_{0}$ can be specified as the negative closest distance between the robot and environmental obstacles.
+This safe set is defined by the zero-sublevel set of a continuous and piecewise smooth function $\phi_{0}:\mathbb{R}^{n_{s}}\rightarrow\mathbb{R}$, where $\mathcal{S}_{S}=\{s\mid\phi_{0}(s)\leq 0\}$, usually specified by users.
+For instance, for collision avoidance, $\phi_{0}$ can be specified as $d_{min}-d$ where $d$ is the closest distance between the robot and environmental obstacles and $d_{min}$ is the distance margin.
 2.2 Problem
 We are interested in the safety imperative of averting collisions for mobile robots navigating 2D planes.
-We aim to persistently satisfy safety specifications at every time step while solving MDP, following the intuition of State-wise Constrained Markov Decision Process (SCMDP) Zhao et al.
-(2023c).
+We aim to persistently satisfy safety specifications at every time step while solving MDP, following the intuition of State-wise Constrained Markov Decision Process (SCMDP) [Zhao et al.
+(2023c)].
 Formally, the set of feasible stationary policies for SCMDP is defined as
@@
 The primary focus of this paper centers on solving (4), i.e., ensuring no safety violation during the training process, while achieving convergence of the policy to the optimal solution of (4).
-3 Prior Works
-3.1 Safe Reinforcement Learning
-Existing safe RL approaches either consider safety after convergence or safety during training Zhao et al.
-(2023c).
-End-to-end approaches are usually used to ensure safety after convergence Liang et al.
-(2018); Tessler et al.
-(2018); Bohez et al.
-(2019); Ma et al.
-(2021); He et al.
-(2023b).
-However, these approaches cannot avoid unsafe explorations.
-Safety in training is achieved by hierarchical approaches which uses a safeguard to filter out unsafe explorative actions.
-The safeguard relies on the knowledge on the system dynamics, which can be either learned dynamics Dalal et al.
-(2018); Thananjeyan et al.
-(2021); Zhao et al.
-(2022), white-box dynamics Fisac et al.
-(2018); Shao et al.
-(2021), or black-box dynamics Zhao et al.
-(2021).
-Nevertheless, these methods cannot guarantee convergence of the RL policy.
-The proposed approach is the first to address safety both during training and after convergence, which can potentially serve as a general framework to bridge these two types of approaches.
-In the following, we choose the most advanced method in each category to form the proposed approach, so that it relies on the least assumptions.
+3 Preliminary
 3.1 Implicit Safe Set Algorithm
-The Implicit Safe Set Algorithm (ISSA) Zhao et al.
-(2021, 2024) ensures the persistent satisfaction of safety specifications for systems with black-box dynamics (e.g., digital twins or neural networks) through energy function-based optimization.
-Leveraging energy function $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$ and theoretical results from SSA Liu and Tomizuka (2014), ISSA synthesizes a safety index to guarantee that the safe control set $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$ is nonempty.
-And then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant under mild assumptions.
+As a deterministic safety filter, Implicit Safe Set Algorithm (ISSA) [Zhao et al.
+(2021), Zhao et al.
+(2024a)] ensures the persistent satisfaction of safety specifications for systems with black-box dynamics (e.g., digital twins or neural networks) through energy function-based optimization.
+Leveraging energy function $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$ and theoretical results from SSA [Liu and Tomizuka (2014)], ISSA synthesizes a safety index to guarantee that the safe control set $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$ is nonempty.
+And then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant under Assumption 1.
 Here $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$.
-Intuitively, the energy function $\phi$ ensures that low-energy states are safe, while the control signal within $\mathcal{A}_{S}(s)$ reduces $\phi$ whenever the system state is outside the safe region.
-This further derives the forward invariance of $\bar{\mathcal{S}}$, meaning that once the system enters the safe region, its state will stays within the safe part of the state space.
-To match the notion of MDP, we consider discrete-time safe control set $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$.
-Thus ISSA ensures safety by projecting the nominal control $a_{t}^{r}$ generated from RL policy $\pi_{\theta}$ to $\mathcal{A}_{S}^{D}(s_{t})$ through solving the following optimization:
-$\begin{split}&\min_{a_{t}\in\mathcal{A}}\|a_{t}-a^{r}_{t}\|^{2}\\ &\text{s.t. }\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}.\end{split}$ (5)
-The detailed mild assumoptions and safety index synthesis are summarized in Appendix A. And the algorithms for solving (5) are provided in Algorithms 2 and 3.
+Assumption 1
+The system (1) is a second-order mobile platform that avoids 2D obstacles.
+1) The state space is bounded, and the relative acceleration $w$ and angular velocity $z$ to the obstacle are bounded and both can achieve zeros, i.e., $w\in[w_{min},w_{max}]$ for $w_{min}\leq 0\leq w_{max}$ and $z\in[z_{min},z_{max}]$ for $z_{min}\leq 0\leq z_{max}$; 2) For all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$; 3) The discrete-time system time step $dt\to 0$; 4) At any given time, there can at most be one obstacle becoming safety critical (Sparse Obstacle Environment).
+Remark 3.1.
+The bounds in the first assumption will be directly used to synthesize $\phi$.
+The second assumption enables us to turn the question on whether there exists a feasible control in $\mathcal{A}_{S}^{D}$ to the question on whether there exists $z$ and $w$ to decrease $\phi$.
+The third assumption ensures that the discrete time approximation error is small.
+The last assumption ensures that the safety index design rule is applicable to multiple moving obstacles.
+The deterministic ISSA could be used as a safety filter in the discrete-time MDP and is treated as part of the environment during training.
+We define the discrete-time safe control set as $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$.
+The ISSA mechanism ensures safety by projecting the nominal control action $a_{t}$, proposed by the RL policy $\pi_{\theta}$, onto the safe control set $\mathcal{A}_{S}^{D}(s_{t})$ by solving the optimization problem:
+$\begin{split}&\min_{a_{t}^{*}\in\mathcal{A}}\|a_{t}^{*}-a_{t}\|^{2}\\ &\text{s.t. }\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}.\end{split}$ (5)
 3.2 State-wise Constrained Policy Optimization
 Safe RL algorithms under the framework of Constrained Markov Decision Process (CMDP) do not consider state-wise constraints.
-To address this gap, State-wise Constrained Policy Optimization (SCPO) was proposed Zhao et al.
-(2023b) to provide guarantees for state-wise constraint satisfaction in expectation, which is under the framework of State-wise CMDP (SCMDP).
+To address this gap, State-wise Constrained Policy Optimization (SCPO) was proposed [Zhao et al.
+(2023b)] to provide guarantees for state-wise constraint satisfaction in expectation, which is under the framework of State-wise CMDP (SCMDP).
 To achieve this, SCPO directly constrain the expected maximum state-wise cost along the trajectory.
@@
 4 Safety Index Guided State-wise Constrained Policy Optimization
-The core idea of S-3PO is to enforce zero safety violation during training by projecting unsafe actions to the safe set, and then constrain the "imaginary" safety violation (i.e., what if the projection is not done) to ensure convergence of the policy to an optimal safe policy.
+Figure 2: (a) S-3PO pipeline during training. (b) S-3PO pipeline during evaluation.
+The core idea of S-3PO is to enforce zero safety violations during training by projecting unsafe actions to the safe set, and then constraining the "imaginary" safety violations to ensure convergence of the policy to an optimal safe policy.
+As shown in Figure 2, the safety filter is part of the training environment, while the evaluation environment does not have the safety filter.
+4.1 Learn with Imaginary Cost
 Zero Violation Exploration
-To ensure zero violation exploration, we safeguard nominal control via solving (5) at every time step during policy training. In this paper, we adopt ISSA to solve (5).
-For the 2D collision avoidance problem considered in this paper, we choose the same safety index synthesis rule as [Section 4.1, (Zhao et al., 2021)], which is summarized in Appendix A.
-We show in Section 5 that with the safety index synthesis rule, ISSA is guaranteed to find a feasible solution of (5), making the system forward invariance in the set $\mathcal{\bar{S}}$.
+To ensure zero violation exploration, we adopt ISSA as the safety filter and safeguard nominal control via solving (5) at every time step during policy training.
+With the safety index synthesis rule in [Zhao et al.
+(2021)], ISSA is guaranteed to find a feasible solution of (5), making the system forward invariance in the set $\mathcal{\bar{S}}$.
+It is worth mentioning that any energy-function-based method that ensures forward invariance could be used as the safety filter, and the scalar energy function could be used to evaluate the imaginary cost.
+The integration with other energy-function-based methods will be left for future work.
 Learning Safety Measures Safely
-While eliminating safety violations during training is good, this also brings inevitable challenges for the policy learning as it directly eliminates the unsafe experience, making the policy unable to distinguish between safe and unsafe actions.
-To overcome this challenge, the main intuition we have is that instead of directly experiencing unsafe states, i.e.
-$s\notin\mathcal{S}_{S}$, policy can learn to act safely from “imagination", i.e., how unsafe it will be if the safeguard had not been triggered?
-The critical observation we rely upon is that:
-Observation 1
-Define $\Delta\phi_{t}=\Delta\phi(s_{t},a_{t},s_{t+1})\doteq\phi(f(s_{t},a^{r}_{t}))-\phi(f(s_{t},a_{t}))$, i.e.
-the degree of required correction to safeguard $a_{t}^{r}$.
-Therefore, $\Delta\phi_{t}$ can be treated as an imagination on how unsafe the reference action would be, where $\Delta\phi_{t}\leq 0$ means $a_{t}^{r}\in\mathcal{A}_{S}^{D}(s_{t})$.
-Following 1, Equation 4 can be translated to:
+While eliminating safety violations during training is beneficial, it also poses challenges for RL policy training, as RL relies on a trial-and-error process.
+To address this, our key insight is that instead of directly encountering unsafe states ($s\notin\mathcal{S}_{S}$), the policy can leverage an "imaginary cost" to learn about unsafe scenarios without actually experiencing them.
+Imaginary Cost
+Define "imaginary cost" as $\Delta\phi_{t}=\Delta\phi(s_{t},a_{t},s_{t+1})\doteq\phi(f(s_{t},a_{t}))-\phi(f(s_{t},a_{t}^{*}))$, i.e.
+the degree of required correction to safeguard $a_{t}$.
+Here $a_{t}^{*}$ is the projected action by ISSA.
+Therefore, $\Delta\phi_{t}$ can be treated as an imagination on how unsafe the reference action would be, where $\Delta\phi_{t}\leq 0$ means $a_{t}\in\mathcal{A}_{S}^{D}(s_{t})$.
+Following the definition, Equation 4 can be translated to:
 $\max_{\theta}\mathcal{J}(\pi_{\theta}),\text{ s.t. }\pi_{\theta}\in\{\pi\in\Pi\big{|}~{}\forall\Delta\phi_{t}\sim\tau,\Delta\phi_{t}\leq 0\}.$ (6)
 Remark 4.1.
-Policies satisfying (6) ensure there is no imaginary safety violation for any possible $a_{t}^{r}$, making $\pi_{\theta}$ a safe policy as required by (4).
+Policies satisfying (6) ensure there is no imaginary safety violation in expectation for any possible $a_{t}$, making $\pi_{\theta}$ a safe policy as required by (4), to be proved by lemma 5.1.
 4.2 Transfrom State-wise Constraint into Maximum Constraint
 For (6), each state-action transition pair corresponds to a constraint, which is intractable to solve.
-Inspired by Zhao et al.
-(2023c), we constrain the expected maximum state-wise $\Delta\phi$ along the trajectory instead of individual state-action transition $\Delta\phi$.
-Next, by treating $\Delta\phi_{t}$ as an “imaginary” cost, we define a MMDP Zhao et al.
-(2023c) by introducing (i) an up-to-now maximum state-wise cost $M$ within $\mathcal{M}\subset\mathbb{R}$, and (ii) a "cost increment" function $D$, where $D:(\mathcal{S},\mathcal{M})\times\mathcal{A}\times\mathcal{S}\mapsto[0,\mathbb{R}^{+}]$ maps the augmented state-action transition tuple to non-negative cost increments.
+Inspired by [Zhao et al.
+(2023c)], we constrain the expected maximum state-wise $\Delta\phi$ along the trajectory instead of individual state-action transition $\Delta\phi$.
+Next, by treating $\Delta\phi_{t}$ as an “imaginary” cost, we define a MMDP [Zhao et al.
+(2023c)] by introducing (i) an up-to-now maximum state-wise cost $M$ within $\mathcal{M}\subset\mathbb{R}$, and (ii) a "cost increment" function $D$, where $D:(\mathcal{S},\mathcal{M})\times\mathcal{A}\times(\mathcal{S},\mathcal{M})\rightarrow[0,\mathbb{R}^{+}]$ maps the augmented state-action transition tuple to non-negative cost increments.
 We define the augmented state $\hat{s}=(s,M)\in(\mathcal{S},\mathcal{M})\doteq\hat{{\mathcal{S}}}$, where $\hat{{\mathcal{S}}}$ is the augmented state space.
@@
 By setting $D\big{(}\hat{s}_{0},a_{0},\hat{s}_{1}\big{)}=\Delta\phi(s_{0},a_{0},s_{1})$, we have $M=\sum_{k=0}^{t-1}D\big{(}\hat{s}_{k},a_{k},\hat{s}_{k+1}\big{)}$ for $t\geq 1$.
-Hence, we define expected maximum state-wise cost (or $D$-return) for $\pi$:
+Hence, we define expected maximum state-wise cost (or $D$-return) for S-3PO policy $\pi$:
 $\mathcal{J}_{D}(\pi)=\mathbb{E}_{\tau\thicksim\pi}\Bigg{[}\sum_{t=0}^{H}D\big{(}\hat{s}_{t},a_{t},\hat{s}_{t+1}\big{)}\Bigg{]}.$ (8)
@@
 4.3 S-3PO
-To solve (9), we propose S-3PO inspired by recent trust region optimization methods Schulman et al.
-(2015).
+To solve (9), we propose S-3PO under the framework of trust region optimization methods [Schulman et al.
+(2015)].
 S-3PO uses KL divergence distance to restrict the policy search in (9) within a trust region around the most recent policy $\pi_{k}$.
@@
 First, maximizing the objective (expected reward advantage) within the trust region (marked by the KL divergence constraint) theoretically guarantees the worst performance degradation.
-Second, constraining the current cost advantage $A_{D}^{\pi_{k}}$ based on the previous value $\mathcal{J}_{D}(\pi_{k})$ guarantees that the worst-case “imagionary” cost is non-positive at all steps as in (6).
-In turn, the original safety constraint (4) is satisifed.
-We then show in Section 5 that S-3PO achieves (i) state-wise safety guarantee of satisfying (3), and (ii) bounded worst case performance degradation for policy update, by establishing new bounds on the difference in returns between two stochastic policies $\pi$ and $\pi^{\prime}$ for MMDPs.
+Second, $J_{D}(\pi)$ can not be computed at step $k+1$ since the state $s_{k+1}$ is inaccessible, thus we leverage a surrogate function to upper bound the $J_{D}(\pi)$ to guarantee the worst-case “imagionary” cost is non-positive at all steps as in (6).
 4.4 Practical Implementation
@@
 In this case, we perform a recovery update, enforcing the cost advantage $A_{D}^{\pi}$ to decrease from early training steps $k_{\mathrm{safe}}$ while focusing on reward improvements of $A^{\pi}$ towards the end of training, prioritizing safety first and reward performance later.
-Please check Section F.3 for furthur details.
+Check Section C.3 for furthur details.
 5 Theoretical Results
-In this section, We present three theorems, including (i) zero violation exploration (Theorem 5.1), (ii) state-wise safety guarantee of S-3PO (Theorem 5.2), and (iii) worst case performance guarantee of S-3PO (Theorem 5.3).
-The proofs of the three theorems are summarized in Appendix C, Appendix D and Appendix E, respectively.
-Theorem 5.1 (Zero Training Time Violation).
-If the system satisfies Assumption 1, and the safety index design follows the rule described in (11), the implicit safe set algorithm ensures the system is forward invariant in $\mathcal{\bar{S}}\subset\mathcal{S}_{S}$.
-Theorem 5.2 (Safety Guarantee of S-3PO).
-If $\pi_{k},\pi_{k+1}$ are related by applying S-3PO and $s_{t}\in\mathcal{\bar{S}}$, then $s_{t+1}=f(s_{t},\pi_{k+1}(s_{t}))\in\mathcal{\bar{S}}$ in expectation.
-Theorem 5.3 (Worst-case Performance Degradation in S-3PO.).
-After $k_{safe}$ updates, suppose $\pi_{k},\pi_{k+1}$ are related by S-3PO update rules, with $\epsilon^{\pi_{k+1}}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim\pi_{k+1}}[A^{\pi_{k}}(\hat{s},a)]|$, then performance return for $\pi_{k+1}$ satisfies
+In this section, we first present the lemma to show the equivalence between constraining the imaginary cost and constraining the safety violation. Then we present the main conclusion for S3PO.
+Lemma 5.1 (Safety Equivalence under Imaginary Cost).
+For a given policy $\pi$ and any initially safe state $s_{0}$ ($\phi(s_{0})\leq 0$), the following two conditions are equivalent: 1) the corresponding trajectory in the training environment has zero imaginary cost $\max_{t}\Delta\phi_{t}\leq 0$; 2) the corresponding trajectory in the evaluation environment has zero safety violation $\max_{t}{\phi_{t}}\leq 0$.
+And if either condition holds, the two trajectories are the same.
+Theorem 5.2 (Safety and Optimality of S-3PO).
+S-3PO will converge in the training environment to a policy $\pi$ with no imaginary cost in expectation, and bounded worst case reward performance.
+In particular, if $\pi_{k}$ and $\pi_{k+1}$ are related by applying S-3PO (10), then with $\epsilon^{\pi_{k+1}}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim\pi_{k+1}}[A^{\pi_{k}}(\hat{s},a)]|$, the performance of $\pi_{k+1}$ in the training environment satisfies:
 $\mathcal{J}(\pi_{k+1})-\mathcal{J}(\pi_{k})\geq-\frac{\sqrt{2\delta}\gamma\epsilon^{\pi_{k+1}}}{1-\gamma}.$
-Remark 5.4.
-Theorem 5.1 shows zero safety violation exploration by proving the system state will never leave $\mathcal{\bar{S}}$ under the safeguard of ISSA.
-Theorem 5.2 shows that by satisfying the constraint of (10), the new policy is guaranteed to generate safe action, i.e.
-$a\in\mathcal{A}_{S}^{D}$, in expectation.
-Intuitively, Theorem 5.3 shows that with enough training epochs, the reward performance of S-3PO will not deteriorate too much after each update.
+Since the training environment is deterministic and all the assumptions in the theoretical results for SCPO are satisfied, the proof for the theorem directly follows from Proposition 2 from SCPO [Zhao et al.
+(2023b)].
+By lemma 5.1, no imaginary cost in the training environment implies no safety violation in the evaluation environment.
+The theorem then implies that the converged policy could achieve zero safety violation in expectation in the evaluation environment.
+Nevertheless, formally establishing their equivalence in the probability space (e.g., in expectation) will be left for future work.
 6 Experiments
 In our experiments, we aim to answer: Q1: Does S-3PO achieve zero-violation during the training?
-Q2: How does S-3PO without safeguard compare with other state-of-the-art safe RL methods?
+Q2: How does S-3PO without safeguard compare with other advanced safe RL methods?
 Q3: Does S-3PO learn to act without safeguard?
 Q4: How does weighted loss trick impact the performance of S-3PO?
-Q5: Is “imaginary” cost necessary to achieve zero violation?
+Q5: Is “imaginary” cost necessary to make the RL policy learn to achieve zero violation by itself?
 Q6: How does S-3PO scale to high dimensional robots?
 6.1 Experiments Setup
-To answer these questions, we conducted our experiments on the safe reinforcement learning benchmark GUARD Zhao et al. (2023a) which is based on Mujoco and Gym interface.
+To answer these questions, we conducted experiments on the safe reinforcement learning benchmark GUARD [Zhao et al. (2023a)] which is based on Mujoco and Gym interface.
 […]
@@
 (ii) end-to-end constrained safe RL algorithms CPO (Achiam et al., 2017), TRPO-Lagrangian (Bohez et al., 2019), PCPO (Yang et al., 2020), SCPO (Zhao et al., 2023b).
-(iii) We select TRPO as our baseline method since it is state-of-the-art and already has safety-constrained derivatives that can be tested off-the-shelf.
-The full list of parameters of all methods compared can be found in Section F.2.
-Metrics For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost and (iii) cost rate.
-More details are provided in Section F.4.
-We set the limit of cost to 0 for all the safe RL algorithms since we aim to avoid any violation of the constraints.
+(iii) We select TRPO as our baseline method since it already has safety-constrained derivatives that can be tested off-the-shelf.
+The full list of parameters of all methods compared can be found in Section C.2.
+Metrics For comparison, we evaluate algorithm performance based on (i) reward performance, (ii) average episode cost, and (iii) cost rate.
+More details are provided in Section C.4.
+We set the limit of cost to 0 for all safe RL algorithms since no violation of the constraints is allowed.
 6.2 Evaluating S-3PO and Comparison Analysis
-Zero Violation During Training The training performance of four representative test suites are summarized in LABEL:fig:zero_violation_performance, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Theorem 5.1.
-For more experiments, please check Appendix F.
+Zero Violation During Training The training performance of four representative test suites are summarized in LABEL:fig:zero_violation_performance, where the S-3PO algorithm clearly outperforms other baseline methods by achieving zero violations, consistent with the safety guarantee outlined in Zhao et al.
+(2021).
+For more experiments, please check Appendix C.
 This superior performance is attributed to the safeguard mechanism within the S-3PO framework, which effectively corrects unsafe actions at every step, particularly during training.
-Furthermore, as demonstrated in LABEL:fig:comparison_results-main, the reward performance remains comparable to state-of-the-art baselines.
+Furthermore, as demonstrated in LABEL:fig:comparison_results-main, the reward performance remains comparable to advanced baselines.
 This distinct capability of S-3PO ensures safe reinforcement learning with zero safety violations, addressing Q1.
@@
 This allows us to determine whether S-3PO effectively learns a state-wise safe policy through the guidance of the safe set-guided cost.
-As shown in in LABEL:fig:comparison_results-main, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance.
-These findings are consistent with Theorem 5.2 and Theorem 5.3, highlighting that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
-Learn to Act without Safeguard As highlighted in 1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation.
+As shown in LABEL:fig:comparison_results-main, S-3PO demonstrates superior performance even without the safeguard, achieving (i) near-zero average episode cost and (ii) significantly reduced cost rates, all while maintaining competitive reward performance.
+These findings highlights that by minimizing imaginary safety violations, the policy rapidly learns to act safely, which addresses Q2.
+Learn to Act without Safeguard As highlighted in Section 4.1, the key concept behind penalizing imaginary safety violations is to minimize the activation of the safeguard, thereby significantly reducing its computational complexity and enabling real-time implementation.
 To illustrate this, we visualize the average number of times the ISSA-based safeguard is triggered per step in LABEL:fig:comparison_results_ISSA-main.
@@
 In the following work, we will try to implement our algorithm in real robot implementation.
-Appendix A Implicit Safe Set Algorithm Details
-Figure 6: Notations.
-A.1 Circumstances and Assumptions
-In our treatment, both the robot and obstacles assume the form of point-mass circles, confined by finite collision radii.
-The safety criterion adopts the form $\phi_{0}=\max_{i}{\phi_{0}}_{i}$, with ${\phi_{0}}_{i}=d_{min}-d_{i}$.
-Here, $d_{i}$ captures the separation between the robot’s center and the $i$-th obstacle, encompassing both static and non-static entities.
-In this context, we introduce $z$ and $w$ to signify the relative acceleration and relative angular velocity of the robot, respectively, within the obstacle’s frame, as depicted in Figure 6.
-Importantly, the synthesis of the safety index for 2D collision evasion remains independent of specific dynamic models, but under the following assumption:
-Assumption 1 (2D Collision Avoidance)
-1) The state space is bounded, and the relative acceleration and angular velocity are bounded and both can achieve zeros, i.e., $w\in[w_{min},w_{max}]$ for $w_{min}\leq 0\leq w_{max}$ and $z\in[z_{min},z_{max}]$ for $z_{min}\leq 0\leq z_{max}$; 2) For all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$; 3) The discrete-time system time step $dt\to 0$; 4) At any given time, there can at most be one obstacle becoming safety critical, such that $\phi-\eta\geq 0$ (Sparse Obstacle Environment).
-The bounds in the first assumption will be directly used to synthesize $\phi$.
-The second assumption enables us to turn the question on whether these exists a feasible control in $\mathcal{A}_{S}^{D}$ to the question on whether there exists $z$ and $w$ to decrease $\phi$.
-The third assumption ensures that the discrete time approximation error is small.
-The last assumption enables safety index design rule applicable with multiple moving obstacles.
-A.2 Safety Index Synthesis
-Following the rules in (Liu and Tomizuka, 2014), we parameterize the safety index as $\phi=\max_{i}\phi_{i}$, and $\phi_{i}=\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}$, where all $\phi_{i}$ share the same set of tunable parameters $\sigma,n,k,\eta\in\mathbb{R}^{+}$.
-Our goal is to choose these parameters such that $\mathcal{A}_{S}^{D}(s)$ is always nonempty.
-By setting $\eta=0$, the parameterization rule of safety index design rule is defined as:
-$\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}\leq\frac{-z_{min}}{v_{max}}$ (11)
-where $v_{max}$ is the maximum relative velocity that the robot can achieve in the obstacle frame.
-Note that the kinematic constraints $v_{max},z_{min}$ can be obtained by sampling the environment Zhao et al.
-(2021).
-Appendix B Algorithms
-B.1 S-3PO Algorithm
+Acknowledgement
+This work is supported by the National Science Foundation under grant No. 2144489.
+Appendix A Algorithms
+A.1 S-3PO Algorithm
 Algorithm 1 S-3PO
@@
      for $t=0,1,2,\dots$ do
-        Sample nominal action $a^{r}_{t}\sim\pi_{k}(s_{t})$
-        Compute and execute $a_{t}=\mathrm{ISSA}(s_{t},a^{r}_{t})$
+        Sample nominal action $a_{t}\sim\pi_{k}(s_{t})$
+        Compute and execute $a_{t}^{*}=\mathrm{ISSA}(s_{t},a^{r}_{t})$
         Log $\tau\leftarrow\tau\cup\{(s_{t},a_{t},r_{t},s_{t+1},\Delta\phi_{t})\}$
@@
   end for
-B.2 Additional Algorithms
-The main body of AdamBA algorithm is summarized in Algorithm 2, and the main body of ISSA algorithm is summarized in Algorithm 3.
-The inputs for AdamBA are the approximation error bound ($\epsilon$), learning rate ($\beta$), reference control ($a^{r}$), gradient vector covariance ($\Sigma$), gradient vector number ($n$), reference gradient vector ($\vec{v}^{r}$), safety status of reference control ($S$), and the desired safety status of control solution ($S_{goal}$).
-The inputs for ISSA are the approximation error bound ($\epsilon$), the learning rate ($\beta$), gradient vector covariance ($\Sigma$), gradient vector number ($n$) and reference unsafe control ($a^{r}$).
-Algorithm 2 Adaptive Momentum Boundary Approximation
-1:  procedure: AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\vec{v}^{r},S,S_{goal}$)
-2:     Initialize:
-3:     if $\vec{v}^{r}$ is empty then
-4:        Generate $n$ Gaussian distributed unit gradient vectors $\vec{v}_{i}\sim\mathcal{N}(0,\Sigma)$, $i=1,2,\ldots,n$
-5:     else
-6:        Initialize one unit gradient vector $\vec{v}_{1}=\frac{\vec{v}^{r}}{\|\vec{v}^{r}\|}$
-7:     end if
-8:     Approximation:
-9:     for $i=1,2,\cdots,n$ do
-10:        Initialize the approximated boundary point $P_{i}=a^{r}$, and stage = exponential outreach.
-11:        while stage = exponential outreach do
-12:           Set $P_{S}\leftarrow P_{i}$ and $P_{NS}\leftarrow P_{i}$
-13:           $P_{i}=P_{i}+\vec{v}_{i}\beta$
-14:           if $P_{i}$ is out of the control set then
-15:              break
-16:           end if
-17:           if $P_{i}$ safety status $\neq S$  then
-18:              Set $P_{NS}\leftarrow P_{i}$ , stage $\leftarrow$ exponential decay
-19:              break
-20:           end if
-21:           $\beta=2\beta$
-22:        end while
-23:        if stage = exponential decay then
-24:           Apply Bisection method to locate boundary point until $\|P_{NS}-P_{S}\|<\epsilon$
-25:           Set $P_{i}\leftarrow P_{S}$ if $S_{goal}=S$, $P_{i}\leftarrow P_{NS}$ otherwise
-26:        end if
-27:     end for
-28:     Return Approximated Boundary Set $P$
-29:  end procudure
-Algorithm 3 Implicit Safe Set Algorithm (ISSA)
-1:  procedure: ISSA($\epsilon,\beta,\Sigma,n,a^{r}$)
-2:     Phase 1: {Phase 1}
-3:     Use AdamBA($\epsilon,\beta,\Sigma,n,a^{r},\emptyset,\textit{UNSAFE},\textit{SAFE}$) to sample a collection $\mathbb{S}$ of safe control on the boundary of $\mathcal{A}_{S}^{D}$.
-4:     if $\mathbb{S}=\emptyset$ then
-5:        Enter Phase 2
-6:     else
-7:        For each primitive action $a_{i}\in\mathbb{S}$, compute the deviation $d_{i}=\|a_{i}-a^{r}\|^{2}$
-8:        return $\mathop{\mathbf{argmin}}_{a_{i}}d_{i}$
-9:     end if
-10:
-11:     Phase 2: {Phase 2}
-12:     Use grid sampling by iteratively increasing sampling resolution to find an anchor safe control $a^{u}$, s.t. safety status of $a^{u}$ is SAFE.
-13:     Use AdamBA($\epsilon,\frac{\|a^{r}-a^{u}\|}{4},\Sigma,1,a^{r},\frac{a^{u}-a^{r}}{\|a^{u}-a^{r}\|},\textit{UNSAFE},\textit{SAFE}$) to search for boundary point $a^{*}$
-14:     if $a^{*}$ is not found then
-15:        Use AdamBA($\epsilon,\frac{\|a^{r}-a^{u}\|}{4},\Sigma,1,a^{u},\frac{a^{r}-a^{u}}{\|a^{r}-a^{u}\|},\textit{SAFE},\textit{SAFE}$) to search for boundary point ${a^{u}}^{*}$
-16:        Return ${a^{u}}^{*}$
-17:     else
-18:        Return $a^{*}$
-19:     end if
-20:  end procudure
-Appendix C Proof of Theorem 5.1
-To prove the Theorem 5.1 we introduce two important corollaries to show that 1) the set of safe control is always nonempty if we choose a safety index that satisfies the design rule in (11); and 2) the proposed algorithm 3 is guaranteed to find a safe control if there exists one.
-With these two corollaries, it is then straightforward to prove the Theorem 5.1.
-C.1 Feasibility of the Safety Index for Continuous-Time System
-Corollary C.1 (Non-emptiness of the set of safe control).
-If 1) the dynamic system satisfies the conditions in Assumption 1; and 2) the safety index is designed according to the rule in Appendix A, then the robot system in 2D plane has nonempty set of safe control at any state, i.e., $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$.
-Note that the set of safe control $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$ is non-empty if and only if it is non-empty in the following two cases: $\phi(s)-\eta<0$ or $\phi(s)-\eta\geq 0$.
-In the following discussion, we first show that the safety index design rule guarantees a non-empty set of safe control if there’s only one obstacle when $\phi(s)-\eta\geq 0$ (Lemma C.2).
-Then we show that the set of safe control is non-empty if there’s only one obstacle when $\phi(s)-\eta<0$ (Lemma C.4).
-Finally, we leverage Lemma C.2 and Lemma C.4 to show $\mathcal{A}_{S}^{D}(s)$ is non-empty if there’re multiple obstacles at any state.
-Lemma C.2.
-If the dynamic system satisfies the conditions in Assumption 1 and there is only one obstacle in the environment, then the safety index design rule in Appendix A ensures that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for $x$ such that $\phi(s)-\eta\geq 0$.
-Proof C.3.
-For $x$ such that $\phi(s)-\eta\geq 0$, the set of safe control becomes
-$\mathcal{A}_{S}^{D}(s)=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\phi(s)-\eta\}$ (12)
-According to the third condition in Assumption 1, we have $dt\rightarrow 0$.
-Therefore, the discrete-time approximation error approaches zero, i.e., $\phi(f(s,a))=\phi(s)+dt\cdot\dot{\phi}(s,a)+\Delta$, where $\Delta\rightarrow 0$.
-Then we can rewrite (12) as:
-$\mathcal{A}_{S}^{D}(s)=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta/dt\}$ (13)
-According to the definition of $\phi$, we have $\dot{\phi}=-nd^{n-1}\dot{d}-k\ddot{d}$.
-We ignored the subscript $i$ since it is assumed that there is only one obstacle.
-Therefore, the non-emptiness of $\mathcal{A}_{S}^{D}(s)$ in (13) is equivalent to the following condition
-$\forall s\text{ s.t. }\phi(s)\geq\eta,\exists a,\text{s.t. }\ddot{d}\geq\frac{\eta/dt-nd^{n-1}\dot{d}}{k}.$ (14)
-Note that in the 2D problem, $\ddot{d}=-z\cos(\alpha)+v\sin{(\alpha)}w$ and $\dot{d}=-v\cos(\alpha)$.
-According to Assumption 1, there is a surjection from $a$ to $(z,w)\in W:=\{(z,w)\mid z_{min}\leq z\leq z_{max},w_{min}\leq w\leq w_{max}\}$.
-Moreover, according to the definition of safety index, $\phi$ for the 2D problem only depends on $\alpha$, $v$, and $d$.
-Hence condition $\forall s\text{ s.t.
-}\phi(s)\geq\eta$ can be translated to $\forall(\alpha,v,d)\text{ s.t.
-}\sigma+d_{min}^{n}-d^{n}-kv\cos(\alpha)\geq\eta$.
-Denote the later set as
-$\Phi:=\{(\alpha,v,d)\mid\sigma+d_{min}^{n}-d^{n}-kv\cos(\alpha)\geq\eta,v\in[0,v_{max}],d\geq 0,\alpha\in[0,2\pi)\}.$ (15)
-Consequently, condition (5) is equivalent to the following condition
-$\forall(\alpha,v,d)\in\Phi,\exists(z,w)\in W,\text{ s.t. }-z\cos(\alpha)+v\sin(\alpha)w$ $\geq\frac{\eta/dt+nd^{n-1}v\cos(\alpha)}{k}.$ (16)
-According to the safety index design rule, we have $\eta=0$. Then we show (16) holds in different cases.
-Case 1: $v=0$. In this case, we can simply choose $z=0$, then the inequality in (16) holds.
-Case 2: $v\neq 0$ and $\cos(\alpha)\leq 0$.
-Note that velocity $v$ is always non-negative.
-Hence $v>0$.
-In this case, we just need to choose $z=w=0$, then the inequality in (16) holds, where the LHS becomes zero and the RHS becomes $\frac{nd^{n-1}v\cos(\alpha)}{k}$ which is non-positive.
-Case 3: $v\neq 0$ and $\cos(\alpha)>0$. Dividing $v\cos(\alpha)$ on both sides of the inequality and rearranging the inequality, (16) is equivalent to
-$\forall(\alpha,v,d)\in\Phi,\exists(z,w)\in W,\text{ s.t. }-\frac{z}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k}\geq 0,$ (17)
-and (17) can be verified by showing:
-$\min_{(\alpha,v,d)\in\Phi}\max_{(z,w)\in W}(-\frac{z}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k})\geq 0.$ (18)
-Now let us expand the LHS of (18):
-$\min_{(\alpha,v,d)\in\Phi}\max_{(z,w)\in W}(-\frac{z}{v}+\tan(\alpha)w-\frac{nd^{n-1}}{k})$ (19)
-$=$ $\min_{(\alpha,v,d)\in\Phi}(-\frac{z_{min}}{v}+[\tan(\alpha)]_{+}w_{max}+[\tan(\alpha)]_{-}w_{min}-\frac{nd^{n-1}}{k})$ (20)
-$=$ $\min_{\alpha\in[0,2\pi),v\in(0,v_{max}]}(-\frac{z_{min}}{v}+[\tan(\alpha)]_{+}w_{max}+[\tan(\alpha)]_{-}w_{min}-\frac{n(\sigma+d_{min}^{n}+kv\cos(\alpha))^{\frac{n-1}{n}}}{k})$ (21)
-$=$ $-\frac{z_{min}}{v_{max}}-\frac{n(\sigma+d_{min}^{n}+kv_{max})^{\frac{n-1}{n}}}{k}.$ (22)
-The first equality eliminates the inner maximization where $[\tan(\alpha)]_{+}:=\max\{\tan(\alpha),0\}$ and $[\tan(\alpha)]_{-}:=\min\{\tan(\alpha),0\}$.
-The second equality eliminates $d$ according to the constraint in $\Phi$.
-The third equality is achieved when $\alpha=0$ and $v=v_{max}$.
-According to the safety index design rule in Appendix A, (22) is greater than or equal to zero.
-Hence (18) holds, which then implies that the inequality in (16) holds.
-The three cases cover all possible situations. Hence (16) always hold and the claim in the lemma is verified.
-Lemma C.4.
-If the dynamic system satisfies the contidions in Assumption 1 and there is only one obstacle in the environment, then the safety index design rule in Appendix A ensures that $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$ for any $s$ that $\phi(s)-\eta<0$.
-Proof C.5.
-For $s$ such that $\phi(s)-\eta<0$, the set of safe control becomes
-$\mathcal{A}_{S}^{D}(s)=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq 0\}$ (23)
-According to the third assumption in Assumption 1, we have $dt\rightarrow 0$.
-Therefore, the discrete-time approximation error approaches zero, i.e., $\phi(f(s,a))=\phi(s)+dt\cdot\dot{\phi}(s,a)+\Delta$, where $\Delta\rightarrow 0$.
-Then we can rewrite (23) as:
-$\mathcal{A}_{S}^{D}(s)=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\phi/dt\}$ (24)
-Note that $\eta=0$ according to the safety index design rule, then $\phi(s)-\eta<0$ implies that $\phi(s)<0$.
-Hence $-\phi/dt\rightarrow\infty$ since $dt\rightarrow 0$.
-Then as long as $\dot{\phi}$ is bounded, $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$.
-Now we show that $\dot{\phi}$ is bounded.
-According to the definition of safety index design rule, $\dot{\phi}=-nd^{n-1}\dot{d}-k\ddot{d}$.
-We ignored the subscript $i$ since it is assumed that there is only one obstacle.
-According to Assumption 1, we have the state space is bounded, thus both $d$ and $\dot{d}$ are bounded, which implies that $nd^{n-1}\dot{d}$ is bounded.
-Moreover, we have for all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$ according to Assumption 1, which indicates the mapping from $a$ to $(z,w)$ is surjective.
-Since $z$ and $w$ are bounded and both can achieve zeros according to Assumption 1, we have afor-$\forall a$, the corresponding $(z,w)$ are bounded.
-Since $\ddot{d}=-z\cos(\alpha)+v\sin(\alpha)w$, then $\ddot{d}$ is bounded.
-Hence $\mathcal{A}_{S}^{D}(s)=\mathcal{A}$ any $s$ that $\phi(s)-\eta<0$ and the claim is true.
-Proof of Corollary C.1.
-Proof C.6.
-If there is one obstacle, then lemma C.2 and lemma C.4 have proved that $\mathcal{A}_{S}^{D}(s)\neq\emptyset$ for all $s$.
-Now we need to consider the case where there are more than one obstacle but the environment is sparse in the sense that at any time step, there is at most one obstacle which is safety critical, i.e.
-$\phi_{i}\geq 0$.
-To show nonemptiness of $\mathcal{A}_{S}^{D}(s)$, we consider the following two cases.
-In the following discussion, we set $\eta=0$ according to the safety index design rule.
-Case 1: $\phi(s)=\max_{i}\phi_{i}(s)\geq 0$.
-Denote $j:=\arg\max_{i}\phi_{i}(s)$.
-Since there is at most one obstacle that is safety critical, then $\phi_{j}(s)\geq 0$ and $\phi_{k}(s)<0$ for all $k\neq j$.
-Denote ${\mathcal{A}_{S}^{D}}_{j}(s):=\{a\in\mathcal{A}\mid\phi_{j}(f(s,a))\leq\phi_{j}(s)\}$.
-Lemma C.2 ensures that ${\mathcal{A}_{S}^{D}}_{j}(s)$ is nonempty.
-Denote ${\mathcal{A}_{S}^{D}}_{k}(s):=\{a\in\mathcal{A}\mid\phi_{k}(f(s,a))\leq 0\}$ where $k\neq j$.
-Since $\phi_{k}(s)<0$, lemma C.4 ensures that ${\mathcal{A}_{S}^{D}}_{k}(s)=\mathcal{A}$.
-Note that the set of safe control can be written as:
-${\mathcal{A}_{S}^{D}}(s)$ :=$:=\{a\in\mathcal{A}\mid\max_{i}\phi_{i}(f(s,a))\leq\max_{i}\phi_{i}(s)\}$ (25a)
-=$=\{a\in\mathcal{A}\mid\max_{i}\phi_{i}(f(s,a))\leq\phi_{j}(s)\}$ (25b)
-$=\cap_{i}\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq\phi_{j}(s)\}$ (25c)
-$={\mathcal{A}_{S}^{D}}_{j}(s)\neq\emptyset$ (25d)
-Note that the last equality is due to the fact that $\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq\phi_{j}(s)\}\supseteq\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}=\mathcal{A}\supseteq{\mathcal{A}_{S}^{D}}_{j}(s)$ for $i\neq j$.
-Case 2: $\phi(s)=\max_{i}\phi_{i}(s)<0$.
-Therefore, we have $\phi_{i}(s)<0$ for all $i$.
-According to Lemma C.4, $\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}=\mathcal{A}$.
-Hence the set of safe control satisfies the following relationship
-${\mathcal{A}_{S}^{D}}(s)$ :=$:=\{a\in\mathcal{A}\mid\max_{i}\phi_{i}(f(s,a))\leq 0\}$ (26a)
-$=\cap_{i}\{a\in\mathcal{A}\mid\phi_{i}(f(s,a))\leq 0\}$ (26b)
-$=\mathcal{A}\neq\emptyset$ (26c)
-In summary, $\mathcal{A}_{S}^{D}(s)\neq\emptyset,\forall s$ and the claim holds.
-C.2 Feasibility of ISSA
-Corollary C.7 (Feasibility of Algorithm 3).
-If the set of safe control is non-empty, Algorithm 3 can always find a sub-optimal solution of (5) with a finite number of iterations.
-Algorithm 3 executes two phases consecutively where the second phase will be executed if no solution of (5) is returned in the first phase.
-Hence, Algorithm 3 can always find a sub-optimal solution of (5) (safe control on the boundary of $\mathcal{A}_{S}^{D}$) if the solution of (5) can always be found in phase 2.
-Note that Phase 2 first finds an anchor safe control $a^{u}$, then use it with AdamBA (Algorithm 2) to find the solution of (5).
-In the following discussion, we first show that the safety index design rule guarantees $a^{u}$ can be found with finite iterations (Lemma C.8).
-Then we show that AdamBA will return a solution if it enters the exponential decay phase (Lemma C.10).
-Subsequently, we show that the evoked AdamBA procedures in phase 2 will definitely enter exponential decay phase (Lemma C.12).
-Finally, we provide a upper bound of the computation iterations for Algorithm 3.
-Lemma C.8 (Existence).
-If the synthesized safety index can guarantee a non-empty set of safe control, then we can find an anchor point in phase 2 of Algorithm 3 with finite iterations (line 11 in algorithm 3).
-Proof C.9.
-Figure 7: Illustration of the grid sampling to find an anchor control point.
-If the synthesized safety index guarantees a non-empty set of safe control $\mathcal{A}^{D}_{S}$, then there exists a hypercube inside of $\mathcal{A}^{D}_{S}$, i.e.
-$\exists\mathcal{Q}\subset\mathcal{A}_{S}^{D}$, where $\mathcal{Q}$ is a $n_{u}$-dimensional hypercube with the same side length of $l>0$.
-Denote $\zeta_{[i]}=\max_{j,k}\|a_{[i]}^{j}-a_{[i]}^{k}\|$, where $a_{[i]}$ denotes the $i$-th dimension of control $a$, and $a^{j}\in\mathcal{A}^{D}_{S},a^{k}\in\mathcal{A}^{D}_{S}$.
-By directly applying grid sampling in $\mathcal{A}^{D}_{s}$ with sample interval $l^{*}$ at each control dimension, such that $0<l^{*}<l$, the maximum sampling iteration $T^{a}$ for finding an anchor point in phase 2 satisfies the following condition:
-$T^{a}<\prod\limits_{i=1}^{n_{u}}\lceil\frac{\zeta_{[i]}}{l^{*}}\rceil~{},$ (27)
-where $T^{a}$ is a finite number since $l^{*}>0$.
-Then we have proved that we can find an anchor point in phase 2 of Algorithm 3 with finite iteration (i.e., finite sampling time).
-The grid sampling to find an anchor control point is illustrated in Figure 7.
-Lemma C.10 (Convergence).
-If AdamBA enters the exponential decay phase, then it can always return a boundary point approximation (with desired safety status) where the approximation error is upper bounded by $\epsilon$.
-Proof C.11.
-According to Algorithm 2, exponential decay phase applies Bisection method to locate the boundary point $P_{b}$ until $\|P_{NS}-P_{S}\|<\epsilon$.
-Denote the returned approximated boundary point as $P_{return}$, according to line 25, $P_{return}$ is either $P_{NS}$ or $P_{S}$, thus the approximation error satisfies:
-$\|P_{return}-P_{b}\|$ $\leq\max\{\|P_{NS}-P_{b}\|,\|P_{S}-P_{b}\|\}$ (28)
-$\leq\|P_{NS}-P_{S}\|$
-$<\epsilon.$
-Lemma C.12 (Feasibility).
-If we enter the phase 2 of Algorithm 3 with an anchor safe control being sampled, we can always find a local optimal solution of (5).
-Proof C.13.
-According to line 13-16, after an anchor safe control is being sampled, phase 2 of Algorithm 3 will evoke at most two AdamBA processes.
-Hence, Lemma C.12 can be proved by showing one of the two AdamBA will return a local optimal solution for (5).
-Next we show Lemma C.12 holds in two cases.
-Case 1:
-line 13 of Algorithm 3 finds a solution.
-In this case, the first AdamBA process finds a safe control $a^{u}$ solution (the return of AdamBA is a set, whereas the set here has at most one element).
-According to Algorithm 2, a solution will be returned only if AdamBA enters exponential decay stage.
-Hence, according to Lemma C.10, $a^{u}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
-Figure 8: Illustration of the case when it is unable to find $u^{*}$.
-Case 2:
-line 13 of Algorithm 3 fails to find a solution.
-In this case, the second AdamBA process is evoked (line 16 of Algorithm 3).
-Since no solution is returned from the first AdamBA process (13 of Algorithm 3), where we start from $a^{r}$ and exponentially outreach along the direction $\vec{v}_{a}=\frac{a^{u}-a^{r}}{||a^{u}-a^{r}||}$, then all the searched control point along $\vec{v}_{a}$ is UNSAFE.
-Specifically, we summarize the aforementioned scenario in Figure 8, the searched control points are represented as red dots along $\vec{v}_{a}$ (red arrow direction).
-Note that the exponential outreach starts with step size $\beta=\frac{\|a^{u}-a^{r}\|}{4}$, indicating two points $\{a_{s}^{1}$, $a_{s}^{2}\}$ are sampled between $a^{r}$ and $a^{u}$ such that
-$\begin{cases}a_{s}^{1}=a^{r}+\frac{\|a^{u}-a^{r}\|}{4}\vec{v}_{a}=\frac{3a^{r}}{4}+\frac{a^{u}}{4}\\ a_{s}^{2}=a^{r}+\frac{3\|a^{u}-a^{r}\|}{4}\vec{v}_{a}=\frac{a^{r}}{4}+\frac{3a^{u}}{4}\end{cases}$ (29)
-where the safety statuses of both $a_{s}^{1}$ and $a_{s}^{2}$ are UNSAFE.
-During the second AdamBA process (line 16 of Algorithm 3), we start from $a^{u}$ and exponentially outreach along the direction $\vec{v}_{r}=\frac{a^{r}-a^{u}}{||a^{r}-a^{u}||}$ with initial step size $\beta=\frac{\|a^{r}-a^{u}\|}{4}$.
-Hence, denote $\bar{a}_{s}^{1}$ as the first sample point along $\vec{v}_{r}$, and $\bar{a}_{s}^{1}$ satisfies
-$\bar{a}_{s}^{1}=a^{u}+\frac{\|a^{r}-a^{u}\|}{4}\vec{v}_{r}=\frac{a^{r}}{4}+\frac{3a^{u}}{4},$ (30)
-which indicates $\bar{a}_{s}^{1}=a_{s}^{2}$, and the safety status of $\bar{a}_{s}^{1}$ is UNSAFE.
-Since the safe status of $a^{u}$ is SAFE and a UNSAFE point can be sampled during the exponential outreach stage, the second AdamBA process (line 16 of Algorithm 3) will enter exponential decay stage.
-Therefore, according to Lemma C.10, ${a^{u}}^{*}$ will always be returned and ${a^{u}}^{*}$ is close to the boundary of the set of safe control with approximation error upper bounded by $\epsilon$.
-The two cases cover all possible situations. Hence, after an anchor safe control $a^{u}$ is sampled, phase 2 of Algorithm 3 can always find a local optima of (5).
-Proof of Corollary C.7.
-Proof C.14.
-According to Lemma C.8 and Lemma C.12, Algorithm 3 is able to find local optimal solution of (5). Next, we will prove Algorithm 3 can be finished within finite iterations.
-According to Algorithm 3, ISSA include (i) one procedure to find anchor safe control $a_{a}$, and (ii) at most three AdamBA procedures.
-Firstly, based on Lemma C.8, $a_{a}$ can be found within finite iterations.
-Secondly, each AdamBA procedure can be finished within finite iterations due to:
-  • exponential outreach can be finished within finite iterations since the control space is bounded.
-  • exponential decay can be finished within finite iterations since Bisection method will exit within finite iterations.
-Therefore, ISSA can be finished within finite iterations.
-C.3 Proof of Theorem 5.1
-Before we prove Theorem 5.1, we first define set $\mathcal{S}_{S}^{D}:=\{s|\phi(s)\leq 0\}$, and we start with a preliminary result regarding $\mathcal{S}_{S}^{D}$ that is useful for proving the main theorem:
-Lemma C.15 (Forward Invariance of $\mathcal{S}_{S}^{D}$).
-If the control system satisfies Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$.
-Proof C.16.
-If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, then we can ensure the system has nonempty set of safe control at any state by Corollary C.1.
-By Corollary C.7, the implicit safe set algorithm can always find local optima solution of (5).
-The local optima solution always satisfies the constraint $\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}$, which indicates that 1) if $\phi(s_{t_{0}})\leq 0$, then $\phi(s_{t})\leq 0,\forall t\geq t_{0}$.
-Note that $\phi(s)\leq 0$ demonstrates that $s\in\mathcal{S}_{S}^{D}$.
-Proof the Theorem 5.1
-Proof C.17.
-Leveraging Lemma C.15, we then proceed to prove that the forward invariance to the set $\mathcal{S}_{S}^{D}$ guarantees the forward invariance to the set $\mathcal{\bar{S}}\subseteq\mathcal{S}_{S}$ which $\mathcal{\bar{S}}=\mathcal{S}_{S}\cap\mathcal{S}_{S}^{D}$.
-Depending on the relationship between $\mathcal{S}_{S}^{D}$ and $\mathcal{S}_{S}$, there are two cases in the proof which we will discuss below.
-Case 1: $\mathcal{S}_{S}^{D}=\{s|\phi(s)\leq 0\}$ is a subset of $\mathcal{S}_{S}=\{s|\phi_{0}(s)\leq 0\}$.
-In this case, $\mathcal{\bar{S}}=\mathcal{S}_{S}^{D}$.
-According to Lemma C.15, If the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{S}_{S}^{D}$ and hence $\mathcal{\bar{S}}$.
-Case 2: $\mathcal{S}_{S}^{D}=\{s|\phi(s)\leq 0\}$ is NOT a subset of $\mathcal{S}_{S}=\{s|\phi_{0}(s)\leq 0\}$.
-In this case, if $s_{t}\in\mathcal{\bar{S}}$, we have $\phi_{0}(s_{t})=\max_{i}{\phi_{0}}_{i}(s_{t})\leq 0$, which indicates $\forall i,{\phi_{0}}_{i}\leq 0$.
-Firstly, we consider the case where ${\phi_{0}}_{i}(s_{t})<0$.
-Note that ${\phi_{0}}_{i}(s_{t+1})={\phi_{0}}_{i}(s_{t})+\dot{\phi_{0}}_{i}(s_{t})dt+\frac{\ddot{\phi_{0}}_{i}(s_{t})dt^{2}}{2!}+\cdots$, since the state space and control space are both bounded, and $dt\rightarrow 0$ according to Assumption 1, we have ${\phi_{0}}_{i}(s_{t+1})\rightarrow{\phi_{0}}_{i}(s_{t})\leq 0$.
-Secondly, we consider the case where ${\phi_{0}}_{i}(s_{t})=0$.
-Since $s_{t}\in\mathcal{\bar{S}}$, we have $\max_{i}\phi_{i}(s_{t})\leq 0$, which indicates $\forall i,\sigma+d_{min}^{n}-d_{i}^{n}-k\dot{d}_{i}\leq 0$.
-Since ${\phi_{0}}_{i}(s_{t})=0$, we also have $d_{i}=d_{min}$.
-Therefore, the following condition holds:
-$\sigma-k\dot{d}_{i}$ $\leq 0$ (31)
-$\dot{d}_{i}$ $\geq\frac{\sigma}{k}$
-According to the safety index design rule, we have $k,\sigma\in\mathbb{R}^{+}$, which indicates $\dot{d}_{i}>0$. Therefore, we have ${\phi_{0}}_{i}(s_{t+1})<0$.
-Summarizing the above two cases, we have shown that if ${\phi_{0}}_{i}(s_{t})\leq 0$ then ${\phi_{0}}_{i}(s_{t+1})\leq 0$, which indicates if $\forall i,{\phi_{0}}_{i}(s_{t})\leq 0$ then $\forall i,{\phi_{0}}_{i}(s_{t+1})\leq 0$.
-Note that $\forall i,{\phi_{0}}_{i}(s_{t+1})\leq 0$ indicates that $\phi_{0}(s_{t+1})=\max_{i}{\phi_{0}}_{i}(s_{t+1})\leq 0$.
-Therefore, we have that if $s_{t}\in\mathcal{\bar{S}}$ then $s_{t+1}\in\mathcal{S}_{S}$.
-Thus, we also have $s_{t+1}\in\mathcal{\bar{S}}$ by Lemma C.15.
-By induction, we have if $s_{t_{0}}\in\mathcal{\bar{S}}$, $s_{t}\in\mathcal{\bar{S}},\forall t>t_{0}$.
-In summary, by discussing the two cases of whether $\mathcal{S}_{S}^{D}$ is the subset of $\mathcal{S}_{S}$, we have proven that if the control system satisfies the assumptions in Assumption 1, and the safety index design follows the rule described in Appendix A, the implicit safe set algorithm guarantees the forward invariance to the set $\mathcal{\bar{S}}\subseteq\mathcal{S}_{S}$.
-Thus, if the initial state is safe, then the following state will always stay in $\mathcal{\bar{S}}$ which means ISSA can ensure the safety during training.
-Appendix D Proof of Theorem 5.2
-Mathematically, S-3PO requires the policy update at each iteration is bounded within a trust region, and updates policy via solving following optimization:
-$\pi_{k+1}=\mathop{\underset{\pi\in\Pi_{\theta}}{\mathop{\mathbf{argmax}}}}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi_{k}}\\ a\sim\pi\end{subarray}}{\mathbb{E}}[A^{\pi_{k}}(\hat{s},a)]$ (32)
-$~{}~{}~{}~{}~{}\mathop{\mathbf{s.t.}}~{}~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta,$
-$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathcal{J}_{D}(\pi_{k})+\underset{\begin{subarray}{c}\hat{s}\sim\bar{d}^{\pi_{k}}\\ a\sim{\pi}\end{subarray}}{\mathbb{E}}\Bigg{[}A^{\pi_{k}}_{D}(\hat{s},a)\Bigg{]}+2(H+1)\epsilon_{D}^{\pi}\sqrt{\frac{1}{2}\delta}\leq 0$
-where $\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$ is KL divergence between two policy $(\pi^{\prime},\pi)$ at state $\hat{s}$, the set $\{\pi\in\Pi_{\theta}~{}:~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta\}$ is called trust region, $d^{\pi_{k}}\doteq(1-\gamma)\sum_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, $\bar{d}^{\pi_{k}}\doteq\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|{\pi_{k}})$, and $\epsilon^{{\pi}}_{D}\doteq\mathbf{max}_{\hat{s}}|\mathbb{E}_{a\sim{\pi}}[A^{\pi_{k}}_{D}(\hat{s},a)]|$.
-In practice, if ISSA is triggered, we take $D(\hat{s}_{t},a_{t},\hat{s}_{t+1})=max\{(\phi(f(\hat{s}_{t},a_{t}^{r}))-\phi(f(\hat{s}_{t},a_{t})))-M,0\}=max\{\Delta\phi_{t},0\}$.
-Otherwise we take $D(\hat{s}_{t},a_{t},\hat{s}_{t+1})=0$ directly.
-D.1 Preliminaries
-$\dot{d}^{\pi}$ we used is defined as
-$\dot{d}^{\pi}(\hat{s})=\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi).$ (33)
-which has a little difference with $d^{\pi}$ and is used to ensure the continuity of function we used for proof later.
-Then it allows us to express the expected discounted total reward or cost compactly as:
-$\mathcal{J}_{g}(\pi)=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}\left[g(\hat{s},a,\hat{s}^{\prime})\right],$ (34)
-where by $a\sim\pi$, we mean $a\sim\pi(\cdot|\hat{s})$, and by $\hat{s}^{\prime}\sim P$,we mean $\hat{s}^{\prime}\sim P(\cdot|\hat{s},a)$.
-$g(\hat{s},a,\hat{s}^{\prime})$ represents the cost or reward function.
-We drop the explicit notation for the sake of reducing clutter, but it should be clear from context that $a$ and $\hat{s}^{\prime}$ depend on $\hat{s}$.
-[…]
-$\dot{d}^{\pi}$ $=\sum\limits_{t=0}^{H}(\gamma P_{\pi})^{t}\hat{\mu}$ (35)
-$=(I-(\gamma P_{\pi})^{H+1})(I-\gamma P_{\pi})^{-1}\hat{\mu}$
-$=(I-\gamma P_{\pi})^{-1}\hat{\mu}$
-Noticing that the finite MDP ends up at step $H$, thus $(P_{\pi})^{H+1}$ should be set to zero matrix.
-This formulation helps us easily obtain the following lemma.
-Lemma D.1.
-For any function $f:\hat{\mathcal{S}}\mapsto\mathbb{R}$ and any policy $\pi$,
-$\underset{\hat{s}\sim\hat{\mu}}{\operatorname{E}}[f(\hat{s})]+\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\gamma f(\hat{s}^{\prime})]-\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}[f(\hat{s})]=0.$ (36)
-Proof D.2.
-Multiply both sides of (35) by $(I-\gamma P_{\pi})$ and take the inner product with the vector $f\in\mathbb{R}^{|\hat{\mathcal{S}}|}$.
-Combining Lemma D.1 with (34), we obtain the following, for any function $f$ and any policy $\pi$:
-$\mathcal{J}_{g}(\pi)=\underset{\hat{s}\sim\hat{\mu}}{\operatorname{E}}\left[f(\hat{s})\right]+\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}\left[g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})\right]$ (37)
-Lemma D.3.
-For any function $f\mapsto\mathbb{R}$ and any policies $\pi^{\prime}$ and $\pi$, define
-$L_{\pi,f}(\pi^{\prime})\doteq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}\left[\left(\dfrac{\pi^{\prime}(a|\hat{s})}{\pi(a|\hat{s})}-1\right)(g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s}))\right],$ (38)
-and $\epsilon_{f}^{\pi\prime}\doteq\max_{\hat{s}}|\mathrm{E}_{a\sim\pi^{\prime},\hat{s}^{\prime}\sim P}[g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})]|.$ Then the following bounds hold:
-$\mathcal{J}_{g}(\pi^{\prime})-\mathcal{J}_{g}(\pi)\geq L_{\pi,f}(\pi^{\prime})-\epsilon_{f}^{\pi^{\prime}}\left\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\right\|_{1},$ (39)
-$\mathcal{J}_{g}(\pi^{\prime})-\mathcal{J}_{g}(\pi)\leq L_{\pi,f}(\pi^{\prime})+\epsilon_{f}^{\pi^{\prime}}\left\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\right\|_{1},$ (40)
-where $D_{TV}$ is the total variational divergence. Furthermore, the bounds are tight(when $\pi^{\prime}$ = $\pi$, the LHS and RHS are identically zero).
-Proof D.4.
-First, for notational convenience, let $\delta_{f}(\hat{s},a,\hat{s}^{\prime})\doteq g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})$. By (37), we obtain the identity
-$\mathcal{J}_{g}(\pi^{\prime})-\mathcal{J}_{g}(\pi)=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi^{\prime}}\\ a\sim{\pi^{\prime}}\\ \hat{s}^{\prime}\sim P\end{subarray}}{E}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]-\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{E}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$ (41)
-Now, we restrict our attention to the first term in (41).
-Let $\dagger{\delta}_{f}^{\pi^{\prime}}\in\mathbb{R}^{|\hat{\mathcal{S}}|}$ denote the vector of components, where $\dagger{\delta}_{f}^{\pi^{\prime}}(\hat{s})=\mathbb{E}_{a\sim\pi^{\prime},\hat{s}^{\prime}\sim P}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})|\hat{s}]$.
-Observe that
-$\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi^{\prime}}\\ a\sim{\pi^{\prime}}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$ $=\left\langle\dot{d}^{\pi^{\prime}},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle$
-$=\left\langle\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle+\left\langle\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle$
-With the Hölder’s inequality; for any $p,q\in[1,\infty]$ such that $\dfrac{1}{p}+\dfrac{1}{q}=1$, we have
-$\left\langle\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle+\left\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\right\|_{p}\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{q}\geq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi^{\prime}}\\ a\sim{\pi^{\prime}}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]\geq\left\langle\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle-\left\|d^{\pi^{\prime}}-\dot{d}^{\pi}\right\|_{p}\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{q}$ (42)
-We choose $p=1$ and $q=\infty$; With $\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{\infty}=\epsilon_{f}^{\pi^{\prime}}$, and by the importance sampling identity, we have
-$\left\langle\dot{d}^{\pi},\dagger{\delta}_{f}^{\pi^{\prime}}\right\rangle$ $=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi^{\prime}}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$ (43)
-$=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim{\pi}\\ \hat{s}^{\prime}\sim P\end{subarray}}{\operatorname{E}}[\left(\frac{\pi^{\prime}(a|\hat{s})}{\pi(a|\hat{s})}\right)\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$
-After bringing (43), $\left\|\dagger{\delta}_{f}^{\pi^{\prime}}\right\|_{\infty}$ into (42), then substract $\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{E}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]$, the bounds are obtained.
-The lower bound leads to (39), and the upper bound leads to (40).
-Lemma D.5.
-The divergence between discounted future state visitation distributions, $||\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}||_{1}$, is bounded by an average divergence of the policies $\pi^{\prime}$ and $\pi$:
-$\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\|_{1}\leq 2\sum_{t=0}^{H}\gamma^{t+1}\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}\left[D_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right],$ (44)
-where $D_{TV}(\pi^{\prime}||\pi)[\hat{s}]=\dfrac{1}{2}\sum_{a}|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|.$
-Proof D.6.
-Firstly, we introduce an identity for the vector difference of the discounted future state visitation distributions on two different policies, $\pi^{\prime}$ and $\pi$.
-Define the matrices $G\doteq(I-\gamma P_{\pi})^{-1},\bar{G}\doteq(I-\gamma P_{\pi^{\prime}})^{-1}$, and $\Delta=P_{\pi^{\prime}}-P_{\pi}$.
-Then:
-$G^{-1}-\bar{G}^{-1}$ $=(I-\gamma P_{\pi})-(I-\gamma P_{\pi^{\prime}})$ (45)
-$=\gamma\Delta,$
-left-multiplying by $G$ and right-multiplying by $\bar{G}$, we obtain
-$\bar{G}-G=\gamma\bar{G}\Delta G.$ (46)
-Thus, the following equality holds:
-$\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}$ $=(1-\gamma)\left(\bar{G}-G\right)\hat{\mu}$ (47)
-$=\gamma(1-\gamma)\bar{G}\Delta G\hat{\mu}$
-$=\gamma\bar{G}\Delta\dot{d}^{\pi}.$
-Using (47), we obtain
-$\|\dot{d}^{\pi^{\prime}}-\dot{d}^{\pi}\|_{1}$ $=\gamma\|\bar{G}\Delta d^{\pi}\|_{1}$ (48)
-$\leq\gamma\|\bar{G}\|_{1}\|\Delta\dot{d}^{\pi}\|_{1},$
-where $||\bar{G}||_{1}$ is bounded by:
-$\|\bar{G}\|_{1}=\|(I-\gamma P_{\pi^{\prime}})^{-1}\|_{1}\leq\sum\limits_{t=0}^{\infty}\gamma^{t}\|P_{\pi^{\prime}}^{t}\|_{1}=\sum\limits_{t=0}^{H}\gamma^{t}.$ (49)
-Next, we bound $\|\Delta\dot{d}^{\pi}_{1}\|$ as following:
-$\|\Delta\dot{d}^{\pi}\|_{1}$ $=\quad\sum\limits_{\hat{s}^{\prime}}\left|\sum\limits_{\hat{s}}\Delta(\hat{s}^{\prime}|\hat{s})\dot{d}^{\pi}(\hat{s})\right|$ (50)
-$\leq\quad\sum\limits_{\hat{s},\hat{s}^{\prime}}|\Delta(\hat{s}^{\prime}|\hat{s})|\dot{d}^{\pi}(\hat{s})$
-$=\quad\sum_{\hat{s},\hat{s}^{\prime}}\left|\sum_{a}P(\hat{s}^{\prime}|\hat{s},a)\left(\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})\right)\right|\dot{d}^{\pi}(\hat{s})$
-$\leq\quad\sum_{\hat{s},a,\hat{s}^{\prime}}P(\hat{s}^{\prime}|\hat{s},a)|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|\dot{d}^{\pi}(\hat{s})$
-$=\quad\sum_{\hat{s},a}|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|\dot{d}^{\pi}(\hat{s})$
-$=\quad 2\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}[D_{TV}(\pi^{\prime}||\pi)[\hat{s}]].$
-By taking (50) and (49) into (48), this lemma is proved.
-The new policy improvement bound follows immediately.
-Lemma D.7.
-For any function $f:\hat{\mathcal{S}}\mapsto\mathbb{R}$ and any policies $\pi^{\prime}$ and $\pi$, define $\delta_{f}(\hat{s},a,\hat{s}^{\prime})\doteq g(\hat{s},a,\hat{s}^{\prime})+\gamma f(\hat{s}^{\prime})-f(\hat{s})$,
-$\begin{gathered}\epsilon_{f}^{\pi^{\prime}}\doteq\operatorname*{max}_{\hat{s}}|\mathrm{E}_{a\sim\pi^{\prime},\hat{s}^{\prime}\sim P}[\delta_{f}(\hat{s},a,\hat{s}^{\prime})]|,\\ L_{\pi,f}(\pi^{\prime})\doteq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi\\ \hat{s}^{\prime}\sim P\end{subarray}}{\mathrm{E}}\left[\left(\frac{\pi^{\prime}(a|\hat{s})}{\pi(a|\hat{s})}-1\right)\delta_{f}(\hat{s},a,\hat{s}^{\prime})\right],and\\ D_{\pi,f}^{\pm}(\pi^{\prime})\doteq L_{\pi,f}(\pi^{\prime})\pm 2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon_{f}^{\pi^{\prime}}\underset{\hat{s}\sim\dot{d}^{\pi}}{\operatorname{E}}[D_{TV}(\pi^{\prime}||\pi)[\hat{s}]],\end{gathered}$
-where $D_{TV}(\pi^{\prime}||\pi)[\hat{s}]=\frac{1}{2}\sum_{a}|\pi^{\prime}(a|\hat{s})-\pi(a|\hat{s})|$ is the total variational divergence between action distributions at $\hat{s}$.
-The following bounds hold:
-$D_{\pi,f}^{+}(\pi^{\prime})\geq\mathcal{J}_{g}(\pi^{\prime})-\mathcal{J}_{g}(\pi)\geq D_{\pi,f}^{-}(\pi^{\prime}).$
-Furthermore, the bounds are tight (when $\pi^{\prime}$ = $\pi$, all three expressions are identically zero)
-Proof D.8.
-Begin with the bounds from lemma D.3 and bound the divergence $D_{TV}(\dot{d}^{\pi^{\prime}}||\dot{d}^{\pi})$ by lemma D.5.
-D.2 Safety invariance in expectation between S-3PO policies
-Corollary D.9 (SCPO Update Constraint Satisfaction).
-Suppose $\mathcal{J}_{D}(\pi_{k})\leq 0$ and $\pi_{k},\pi_{k+1}$ are related by (32), then $D$-return for $\pi_{k+1}$ satisfies
-$\forall i,\mathcal{J}_{D}(\pi_{k+1})\leq 0.$
-Proof D.10.
-The choice of $f=\hat{V}_{D}^{\pi},g=D$ in lemma D.7 leads to following inequality:
-$\mathcal{\hat{J}}_{D}(\pi^{\prime})-\mathcal{\hat{J}}_{D}(\pi)\leq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[\hat{A}_{D}^{\pi}(\hat{s},a)+2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right].$ (51)
-where $\mathcal{\hat{J}}_{D}(\pi)=\mathbb{E}_{\tau\thicksim\pi}\Bigg{[}\sum_{t=0}^{H}\gamma^{t}D\big{(}\hat{s}_{t},a_{t},\hat{s}_{t+1}\big{)}\Bigg{]}$, need to distinguish from $\mathcal{J}_{D}(\pi)$.
-And $\hat{V}_{D}^{\pi},\hat{A}_{D}^{\pi}$ are also the discounted version of $V_{D}^{\pi}$ and $A_{D}^{\pi}$.
-Note that according to Lemma D.7 one can only get this the inequality holds when $\gamma\in(0,1)$.
-Then we can define $\mathcal{F}(\gamma)=\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[\hat{A}_{D}^{\pi}(\hat{s},a)+2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right]-\mathcal{\hat{J}}_{D}(\pi^{\prime})+\mathcal{\hat{J}}_{D}(\pi)$ with the following condition holds:
-$\mathcal{F}(\gamma)\geq 0,\text{when}~{}\gamma\in(0,1)$ (52)
-$\mathcal{F}(\gamma)\text{'s domain of definition is }\mathcal{R}$
-$\mathcal{F}(\gamma)\text{ is a polynomial function}$
-Because $\mathcal{F}(\gamma)$ is a polynomial function and coefficients are all limited, thus $\underset{\gamma\rightarrow 1^{-}}{\lim}\mathcal{F}(\gamma)$ exists and $\mathcal{F}(\gamma)$ is continuous at point $(1,\mathcal{F}(1)$).
-So $\mathcal{F}(1)=\underset{\gamma\rightarrow 1^{-}}{\lim}\mathcal{F}(\gamma)\geq 0$, which equals to:
-$\mathcal{J}_{D}(\pi^{\prime})-\mathcal{J}_{D}(\pi)\leq\underset{\begin{subarray}{c}\hat{s}\sim\bar{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\left[A_{D}^{\pi}(\hat{s},a)+2(H+1)\epsilon_{D}^{\pi^{\prime}}\mathcal{D}_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right].$
-where $\bar{d}^{\pi}=\sum_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|\pi)$. Thus, following the inequality (32), the Corollary D.9 is proofed.
-D.3 Proof of Theorem 5.2
-If $s_{t}\in\mathcal{\bar{S}}\in\mathcal{S}_{S}^{D}$, then $\mathcal{J}_{D}(\pi_{t})\leq 0$ and $\phi(s_{t})\leq 0$.
-According to Corollary D.9, we know that if policy $\pi_{t}$ is updated by solving (32), then we have $\mathcal{J}_{D}(\pi_{t+1})\leq 0$ which means:
-$\underset{\tau\sim\pi_{t+1}}{\mathbb{E}}\left[\phi(s_{t+1})\right]=\mathcal{J}_{D}(\pi_{t+1})+\phi(s_{t})\leq 0$ (53)
-Thus $s_{t+1}\in\mathcal{S}_{S}^{D}$ in expectation. According to the proof of Section C.3, we know that $s_{t+1}\in\mathcal{\bar{S}}\in\mathcal{S}_{S}^{D}$ in expectation.
-Appendix E Proof of Theorem 5.3
-E.1 KL Divergence Relationship Between $d^{\pi_{k}}$ and $\bar{d}^{\pi_{k}}$
-Lemma E.1.
-$\underset{\hat{s}\sim d_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]<\underset{\hat{s}\sim\bar{d}_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]$
-Proof E.2.
-$\underset{\hat{s}\sim d_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]$ $=\sum_{\hat{s}}(1-\gamma)\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi)\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$
-$<\sum_{\hat{s}}\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi)\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$
-$<\sum_{\hat{s}}\sum\limits_{t=0}^{H}P(\hat{s}_{t}=\hat{s}|\pi)\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]$
-$=\underset{\hat{s}\sim\bar{d}_{\pi}}{\mathrm{E}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]].$
-E.2 Trust Region Update Performance
-Lemma E.3.
-For any policies $\pi^{\prime},\pi$, with $\epsilon^{\pi^{\prime}}\doteq max_{\hat{s}}|E_{a\sim\pi^{\prime}}[A^{\pi}(\hat{s},a)]|$, and define $d^{\pi}=(1-\gamma)\sum\limits_{t=0}^{H}\gamma^{t}P(\hat{s}_{t}=\hat{s}|\pi)$ as the discounted augmented state distribution using $\pi$, then the following bound holds:
-$\mathcal{J}(\pi^{\prime})-\mathcal{J}(\pi)>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim{\pi^{\prime}}\end{subarray}}{\mathbb{E}}\bigg{[}A^{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}\mathcal{D}_{TV}({\pi^{\prime}}\|\pi)[\hat{s}]\bigg{]}$ (54)
-Proof E.4.
-The choice of $f=V_{\pi},g=R$ in lemma D.7 leads to following inequality:
-For any policies $\pi^{\prime},\pi$, with $\epsilon^{\pi^{\prime}}\doteq max_{\hat{s}}|E_{a\sim\pi^{\prime}}[A_{\pi}(\hat{s},a)]|$, the following bound holds:
-$\mathcal{J}(\pi^{\prime})-\mathcal{J}(\pi)$ $\geq\underset{\begin{subarray}{c}\hat{s}\sim\dot{d}^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathrm{E}}\left[A_{\pi}(\hat{s},a)-2(\sum_{t=0}^{H}\gamma^{t+1})\epsilon^{\pi^{\prime}}D_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right]$
-$>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathrm{E}}\left[A_{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}D_{TV}(\pi^{\prime}||\pi)[\hat{s}]\right]$
-At this point, the lemma E.3 is proved.
-E.3 Proof of Theorem 5.3
-Proof E.5.
-If ISSA will constantly be triggered, the safeguard should be treated as a component of the environment, which indicates the environment is non-stationary for policy $\pi$ after safeguard is disabled.
-According to Algorithm 1, we assume that ISSA will not be triggered after $k_{safe}$ updates, meaning the environment is stationary for policy $\pi$ with or without safeguard.
-Therefore, we can infer S-3PO worst performance degradation after $k_{safe}$ updates following the trust region results of finite-horizon MDPs.
-Utilizing Lemma E.3 and the relationship between the total variation divergence and the KL divergence, we have:
-$\mathcal{J}(\pi^{\prime})-\mathcal{J}(\pi)>\frac{1}{1-\gamma}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi}\\ a\sim\pi^{\prime}\end{subarray}}{\mathbb{E}}\bigg{[}A^{\pi}(\hat{s},a)-\frac{2\gamma\epsilon^{\pi^{\prime}}}{1-\gamma}\sqrt{\frac{1}{2}\mathbb{E}_{\hat{s}\sim d^{\pi}}[\mathcal{D}_{KL}(\pi^{\prime}\|\pi)[\hat{s}]]}\bigg{]}.$ (55)
-In (32), the reward performance between two policies is associated with trust region, i.e.
-$\pi_{k+1}=\mathop{\underset{\pi\in\Pi_{\theta}}{\mathop{\mathbf{argmax}}}}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi_{k}}\\ a\sim\pi\end{subarray}}{\mathbb{E}}[A^{\pi_{k}}(\hat{s},a)]$ (56)
-$~{}~{}~{}~{}~{}\mathop{\mathbf{s.t.}}~{}~{}\mathbb{E}_{\hat{s}\sim\bar{d}^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta.$
-Due to Lemma E.1, if two policies are related with Equation 56, they are related with the following optimization:
-$\pi_{k+1}=\mathop{\underset{\pi\in\Pi_{\theta}}{\mathop{\mathbf{argmax}}}}\underset{\begin{subarray}{c}\hat{s}\sim d^{\pi_{k}}\\ a\sim\pi\end{subarray}}{\mathbb{E}}[A^{\pi_{k}}(\hat{s},a)]$ (57)
-$~{}~{}~{}~{}~{}\mathop{\mathbf{s.t.}}~{}~{}\mathbb{E}_{\hat{s}\sim d^{\pi_{k}}}[\mathcal{D}_{KL}(\pi\|\pi_{k})[\hat{s}]]\leq\delta.$
-By (55) and (57), if $\pi_{k},\pi_{k+1}$ are related by (32), then performance return for $\pi_{k+1}$ satisfies
-$\mathcal{J}(\pi_{k+1})-\mathcal{J}(\pi_{k})>-\frac{\sqrt{2\delta}\gamma\epsilon^{\pi_{k+1}}}{1-\gamma}.$
-Appendix F Expeiment Details
-F.1 Environment Settings
+Appendix B Proof of Lemma 5.1
+Proof B.1.
+To prove that the two conditions are equivalent, we show that condition 2 can be derived from condition 1, and vice versa.
+(Condition 1 $\Rightarrow$ Condition 2):
+By the property of ISSA, $\Delta\phi>0$ only when the safety filter is triggered.
+Under condition 1, we have $\Delta\phi\leq 0$ throughout the trajectory in the training environment, which implies that ISSA is never activated during training.
+As a result, the trajectory remains entirely within the safe state set, satisfying $\max_{t}\phi_{t}\leq 0$.
+Moreover, since ISSA is inactive, the applied action coincides with the nominal policy action, i.e., $a_{t}^{*}=a_{t}$ for all $t$.
+Therefore, the same sequence of actions will be executed in the evaluation environment (which uses the same initial state), producing the same trajectory that stays within the safe set, with $\max_{t}\phi_{t}\leq 0$ and without triggering ISSA.
+(Condition 2 $\Rightarrow$ Condition 1):
+Conversely, if a given policy starting from a certain initial state generates a trajectory in the evaluation environment that never triggers ISSA and satisfies $\max_{t}\phi_{t}\leq 0$, the same trajectory will be produced in the training environment when initialized from the same state.
+Along this trajectory, ISSA remains inactive, implying that the imaginary cost remains zero, i.e., $\Delta\phi_{t}\leq 0$ for all $t$.
+Since both conditions ensure that ISSA is never triggered along the entire trajectory, the actions remain unchanged, $a_{t}^{*}=a_{t}$ for all $t$, and the resulting trajectories are identical in both the training and evaluation environments.
+It is important to note that the guarantee in Lemma 5.1 holds only when both conditions are strictly satisfied over the entire trajectory distribution induced by the policy.
+In contrast, for the probability space where the conditions are satisfied in expectation, the trajectories in the evaluation environment may deviate from those in the training environment, even when initialized from the same initial state.
+Such deviation can occur if either the imaginary cost or the safety index $\phi$ becomes positive at any step along the trajectory.
+The deviation may lead to discrepancies in the policy performance during evaluation, which will be analyzed in future work.
+Appendix C Expeiment Details
+C.1 Environment Settings
 Goal Task
@@
 Additional high dimensional link robot
-To scale our method to high dimensional link robots. We additionally adopt Walker shown in 13 as a bipedal robot ($\mathcal{A}\subseteq\mathbb{R}^{10}$) in our experiments.
+To scale our method to high dimensional link robots. We additionally adopt Walker shown in 11 as a bipedal robot ($\mathcal{A}\subseteq\mathbb{R}^{10}$) in our experiments.
 Figure 11: Walker
@@
 For all the experiments, the control space of all robots are continuous, and linearly scaled to [-1, +1].
-F.2 Policy Settings
+C.2 Policy Settings
 The hyper-parameters used in our experiments are listed in Table 4 as default.
@@
 Cost reduction - - - 0.0 - 0.0 0.0
-F.3 Practical Implementation Details
+C.3 Practical Implementation Details
 In this section, we summarize implementation techniques that helps with S-3PO’s practical performance. The pseudocode of S-3PO is give as algorithm 1.
@@
 In that way, S-3PO achieves zero safety violation both during training and in testing with a worst-case performance degradation guarantee.
-F.4 Metrics Comparison
+C.4 Metrics Comparison
 In Tables 5, 6 and 7, we report all the $9$ results of our test suites by three metrics: